### v1

# Concept-Based Interpretable Reinforcement Learning with Limited to No Human Labels

Abstract
Recent advances in reinforcement learning (RL) have predominantly leveraged neural network-based policies for decision-making, yet these models often lack interpretability, posing challenges for stakeholder comprehension and trust. Concept bottleneck models offer an interpretable alternative by integrating human-understandable concepts into neural networks. However, a significant limitation in prior work is the assumption that human annotations for these concepts are readily available during training, necessitating continuous real-time input from human annotators. To overcome this limitation, we introduce a novel training scheme that enables RL algorithms to efficiently learn a concept-based policy by only querying humans to label a small set of data, or in the extreme case, without any human labels. Our algorithm, LICORICE, involves three main contributions: interleaving concept learning and RL training, using a concept ensembles to actively select informative data points for labeling, and decorrelating the concept data with a simple strategy. We show how LICORICE reduces manual labeling efforts to to 500 or fewer concept labels in three environments. Finally, we present an initial study to explore how we can use powerful vision-language models to infer concepts from raw visual inputs without explicit labels at minimal cost to performance.
1 Introduction
In reinforcement learning (RL), agents are tasked with learning a policy, a rule that makes sequential, reactive decisions in complex environments. In recent RL work, agents typically represent the policy as a neural network, as such representations tend to lead to high performance [17]. However, this choice can come at a cost: such policies are challenging for stakeholders to interpret — particularly when the inputs to the network are also complex, such as high-dimensional sensor data. This opacity can become a significant hurdle, especially in applications where understanding the rationale behind decisions is critical, such as healthcare [32] or finance [15]. In such applications, decisions can have significant consequences, so it is essential for stakeholders to fully grasp the reasoning behind actions in order to confidently adopt or collaborate on a policy.
To address interpretability concerns in the supervised learning setting, recent works have integrated human-understandable concepts into the decision-making process through concept bottleneck models [13, 9]. These models incorporate a bottleneck layer whose units correspond to interpretable concepts, ensuring that the final decisions depend on these concepts instead of just on opaque raw inputs. By training the model both to have high accuracy on the target task and to accurately match experts’ concept labels, these models learn a high-level concept-based representation that is simultaneously meaningful to humans and useful for machine learning tasks. As an example, a concept-based explanation for a bird classification task might include a unit that encodes the bird’s wing color.
More recently, these techniques have been applied to RL by incorporating a concept bottleneck in the policy [11, 33], such that the actions taken by the agent are a function of the human-understandable concepts. However, past work assumes that human-provided concept annotations are accessible in the training loop of RL, which presents a significant challenge. In novel domains, we may only have access to data that is not in an interpretable representation, such as RGB or multispectral satellite imagery, while the RL agent needs to know the concepts present in every state and action it encounters during training — even though RL training sets can often be measured in millions or billions of state-action pairs. As a result, the human labelers would need to be actively involved in the RL training process, providing concept labels for unreasonable numbers of observations. This level of involvement is not only detrimental to the human labeler but also risks introducing potential biases and errors in the model training process due to fatigue [16].
In this work, we aim to address this challenge of requiring frequent human interventions during the training of interpretable policies using RL. We propose LICORICE (Label-efficient Interpretable COncept-based ReInforCEment learning), a novel training paradigm consisting of three main contributions. First, LICORICE interleaves concept learning and RL training: it alternately freezes the network layers corresponding to either the concept learning part or the decision-making part. We believe this scheme improves our ability to learn from limited labeled data by reducing interference between the learning tasks. Additionally, concept learning uses data that is more on-policy, which means it is more relevant and useful for learning accurate concepts that directly impact the decision-making process. Second, LICORICE utilizes concept ensembles to actively select the most informative data points for labeling. By focusing on samples that are predicted to provide the most valuable information for model improvement, this technique substantially reduces the number of labels needed to achieve both high performance and high concept accuracy. Third, LICORICE includes a strategy to decorrelate the concept data collected under the current policy. By generating a diverse set of training data, this approach ensures the data remains unbiased and representative of various scenarios. We demonstrate how these changes yield both higher concept accuracy and higher reward while requiring fewer queries on three environments with image input, including an image input version of CartPole and two Minigrid environments.
Given these results, we ask whether we can further reduce the reliance on manual concept labeling by leveraging the potential of vision-language models (VLMs). This capability is important in scenarios where manual concept annotation is impractical. We present an initial exploration of whether VLMs can further reduce the concept annotation burden. We find that VLMs can indeed serve as concept annotators for some but not all of the above environments. In these cases, the resulting policies can achieve $83$ to $93$% of the optimal performance.
2 Preliminaries
2.1 Reinforcement Learning
In RL, an agent learns to make decisions by interacting with an environment [27]. The environment is commonly modeled as a Markov decision process [21], consisting of the following components: a set of states $\mathcal{S}$, a set of actions $\mathcal{A}$, a state transition function $T:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ that indicates the probability of transitioning from one state to another given an action, a reward function $R:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ that assigns a reward for each state-action-state transition, and a discount factor $\gamma\in[0,1]$ that determines the present value of future rewards. The agent learns a policy $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$, which maps states to actions with the aim of maximizing the expected cumulative discounted reward. We evaluate a policy via its value function, which is defined as $V^{\pi}(s)=\mathbb{E}_{\pi}[\sum_{k=0}^{\infty}\gamma^{k}r_{t+k+1}\mid s_{0}=s],\forall s\in\mathcal{S}$. The ultimate aim in RL is to determine the optimal policy, $\pi^{*}$. To do so, the agent iteratively refines its policy based on feedback from the environment.
2.2 Concept Bottleneck Models
Concept-based explanations have emerged as a common paradigm in explainable AI [20]. They explain a model’s decision-making process through human-interpretable attributes and abstractions. Concept bottleneck models [13] are an example of concept-based explanations: they learn a mapping from samples $x\in X$ to labels $y\in Y$ through two functions, $g$ and $f$. The concept encoder function $g:X\rightarrow C$ maps samples from the input space to an intermediate space of human-interpretable concepts. The label predictor function $f:C\rightarrow Y$ maps samples from the concept space to a downstream task space, such as labels for supervised classification. As a result, the prediction $\hat{y}=f(g(x))$ relies on the input $x$ entirely through the bottleneck $\hat{c}=g(x)$.
Training these models requires a dataset of $X\times C\times Y$, in which each sample consists of input features $x$, a ground truth concept vector $c\in\{0,1\}^{k}$, and a task label $y$. The functions $f$ and $g$ are parameterized by neural networks. These models are trained independently, jointly, or sequentially. The independent training method trains $\hat{f}$ and $\hat{g}$ independently. Critically, $\hat{f}$ is trained using the true $c$, but at test time it takes $\hat{g}(x)$ as input. If directly applied to the RL setting, this approach would require continuous access to ground-truth concepts for training $\hat{f}$, which we would like to avoid. In joint training, the model minimizes the weighted sum of the concept losses with the task performance loss. This paradigm can be problematic, as the losses can interfere with each other, and it may require careful tuning to find the right balance between the two losses. In sequential training, the model first learns $\hat{g}$. It then uses the concept predictions $\hat{g}(x)$ to learn $\hat{f}(\hat{g}(x))$. As we will later show, this setup is problematic for RL, as concepts may only emerge after the policy is sufficiently performant.
3 LICORICE
We now describe LICORICE, our algorithm for interactively querying for concept labels under a limited concept label budget during RL training. In Section 3.1, we describe the architecture of the concept bottleneck policy. We then detail our iterative training process (Section 3.2) and explain our active learning approach for choosing informative training points (Section 3.3). Finally, in Section 3.4, we detail how we use VLMs as a substitute for human concept labeling.
3.1 Concept Bottleneck Policies
We insert a concept bottleneck layer into the policy network, such that $\pi$ maps from states $\mathcal{S}$ to concepts $C$ to actions $\mathcal{A}$, $\pi=f(g(s))$. These concepts serve as intermediaries in the policy network, which subsequently maps the concept vector to a distribution over possible actions. This setup allows the policy to base its decisions on understandable and meaningful features. We describe additional modifications necessary to accommodate the concept bottleneck in Section A.2. As a result, we can use any RL algorithm as long as we include an additional loss function for concept prediction.
Loss Function
We now describe how we use two loss functions for training in our iterative training scheme, described in more detail in Section 3.2. Because we employ an iterative training scheme, we disentangle the two loss functions — the concept prediction loss and the standard RL loss — to prevent interference between them. In this way, the concept prediction loss $L^{C}$ only affects $g$, the part of the model responsible for predicting concepts, and the standard RL loss $L^{\text{RL}}$ only affects $f$. For training $g$, we employ two types of loss functions depending on the nature of the concepts (MSE for continous concepts; cross-entropy loss for categorical concepts). If the problem requires mixed-type concepts, we could discretize the continuous attributes, converting them into categorical forms suitable for classification. This approach ensures that our method accommodates a diverse range of concept types.
3.2 Training Process
We now present LICORICE, our novel algorithm for optimizing concept-based models for RL within a fixed concept query budget. Our iterative training approach extends the sequential bottleneck training paradigm by incorporating multiple phases. Intuitively, this strategy can be advantageous for obtaining accurate on-policy concept estimates. To collect candidate training data for $g$, we employ our current policy for rollouts. To minimize data point correlation and collect a diverse range of data, we use a decorrelation strategy. To select more informative points to query for concept labels, we use a concept ensemble to calculate disagreement.
Algorithm 1 LICORICE (Label-efficient Interpretable COncept-based ReInforCEment learning)
1:Input: Total budget $B$, number of iterations $M$, sample acceptance threshold $p$, ratio $\tau$ for active learning, batch size for querying $b$, number of concept models $N$ to ensemble
2:Initialize: training set $\mathcal{D}_{\text{train}}\leftarrow\emptyset$, and validation set $\mathcal{D}_{\text{val}}\leftarrow\emptyset$
3:for $m=1$ to $M$ do
4:     Allocate budget for iteration $m$: $B_{m}\leftarrow\frac{B}{M}$
5:     while $|\mathcal{U}_{m}|<\tau\cdot B_{m}$ do
6:         Execute policy $\pi_{m}$ to collect unlabeled data $\mathcal{U}_{m}$ using acceptance rate $p$
7:     end while
8:     Initialize iteration-specific datasets for collecting labeled data: $\mathcal{D}^{\prime}_{\text{train}}\leftarrow\emptyset$, $\mathcal{D}^{\prime}_{\text{val}}\leftarrow\emptyset$
9:     while $B_{m}>0$ do
10:         Train $N$ concept models $\{\tilde{g}_{i}\}^{N}_{i=1}$ on $\mathcal{D}_{\text{train}}\cup\mathcal{D}^{\prime}_{\text{train}}$, using $\mathcal{D}_{\text{val}}\cup\mathcal{D}^{\prime}_{\text{val}}$ for early stopping
11:         Calculate acquisition function value $\alpha(s)$ for all $s\in\mathcal{U}_{m}\setminus(\mathcal{D}^{\prime}_{\text{train}}\cup\mathcal{D}^{\prime}_{\text{val}})$
12:         Choose $b_{m}=\min(b,B_{m})$ unlabeled points from $\mathcal{U}_{m}$ according to $\arg\max_{s}\alpha(s)$
13:         Query for concept labels to obtain new dataset $\mathcal{D}_{m}\leftarrow\{(s,c)\}^{b_{m}}$
14:         Split $\mathcal{D}_{m}$ into train and validation splits and add to $\mathcal{D}^{\prime}_{\text{train}},\mathcal{D}^{\prime}_{\text{val}}$
15:         Decrement budget: $B_{m}\leftarrow B_{m}-b_{m}$
16:     end while
17:     Aggregate datasets: $\mathcal{D}_{\text{train}}\leftarrow\mathcal{D}_{\text{train}}\cup\mathcal{D}^{\prime}_{\text{train}}$, $\mathcal{D}_{\text{val}}\leftarrow\mathcal{D}_{\text{val}}\cup\mathcal{D}^{\prime}_{\text{validation}}$
18:     Continue training the concept network $g$ on $\mathcal{D}_{\text{train}}$, using $\mathcal{D}_{\text{val}}$ for early stopping
19:     Freeze $g$ and train $f$ using standard RL training to obtain $\pi_{m+1}$
20:end for
More specifically, LICORICE (Algorithm 1) proceeds in three main stages within each iteration $m\in M$: data collection, concept ensemble training for data selection, and concept bottleneck policy training. First, in the data collection stage (lines 4 to 8), we collect a dataset of unlabeled concept data $\mathcal{U}_{m}$ by rolling out our current policy $\pi_{m}$. Crucially, during this step, we decorrelate the data to ensure the data points are diverse (line 6). Specifically, we set an acceptance probability $p$ for adding the data point to the dataset, as they are generated by the temporally-correlated RL policy. This prepares a diverse dataset for the next stage, with the aim of promoting the training of a more performant concept ensemble for data selection and, eventually, a more accurate concept bottleneck portion of the policy.
The second stage is concept ensemble training for data selection (lines 9 to 16). In this stage, we train the concept ensemble — consisting of $N$ independent concept models — from scratch on the dataset of training points $\mathcal{D}_{\text{train}}$ that has been aggregated over all iterations (line 10). We use this ensemble to calculate the disagreement-based acquisition function, which evaluates whether each candidate in our unlabeled dataset $\mathcal{U}_{m}$ ought to receive a ground-truth concept label (line 11). This function targets samples where prediction disagreement is highest, detailed in Section 3.3. We then query for $B_{m}$ ground-truth concept labels (line 13) to prepare us for the next stage.
We are now prepared for the third stage: concept bottleneck policy training (lines 17 to 19). We aggregate the concept datasets from the second stage with data from previous iterations, and continue training the concept network $g$ on $\mathcal{D}_{\text{train}}$ (line 18). After early stopping dictates that we stop $g$ training, we freeze $g$ and train $f$ using any standard RL algorithm — for example, PPO — to obtain the policy for collecting data in the next iteration (line 19). With that, we are prepared to start the process again from the first stage.
3.3 Active Concept Learning
We propose to leverage an ensemble of concept models to calculate the disagreement-based acquisition function for actively querying for concept labels. To quantify the concept disagreement of a state $U(s)$, we use two different formulations, depending on the concept learning task.
Classification
For concept classification, we use a query-by-committee [24] approach. When the models produce diverse predictions, it indicates that the instance is difficult and more information would be particularly valuable. Conversely, if all models agree, the instance is likely already well understood, and additional labels would be less beneficial. More specifically, we prioritize points with a high proportion of predicted class labels that are not the modal class prediction (also called the variation ratio [3]). This is given by $U(s)=1-\max_{c\in C}\frac{1}{N}\sum_{i=1}^{N}[\tilde{g}_{i}(s)=c]$, where $\tilde{g}_{i}(s)$ is the concept prediction of the $i$-th model on state $s$.
Regression
For concept regression, we observe that we can instead directly use variance as a measure of disagreement. Specifically, the concept disagreement of a state is quantified by the unbiased estimation of variance $U(s)=\sigma^{2}(s)$ of the predictions across the concept models, due to Bessel’s correction, defined as: $U(s)=\sigma^{2}(s)=\frac{1}{N-1}\sum_{i=1}^{N}(\tilde{g}_{i}(s)-\mu(s))^{2}$, where $\mu(s)=\frac{1}{N}\sum_{i=1}^{N}\tilde{g}_{i}(s)$ is the mean prediction of the $N$ concept models.
3.4 Using Vision-Language Models for Concept Labeling
Equipped with our overall algorithm, we now seek to further reduce human annotation burden. To do so, we turn to VLMs due to their remarkable performance in understanding and generating human-like descriptions of visual content [22]. Within the pipeline of LICORICE, we keep all aspects of our algorithm the same but use the VLM as the concept annotator in the training loop. By doing so, we effectively decrease the human annotation effort to zero, albeit with some labeling inaccuracy introduced.
In our experiments, we use GPT-4o [1], a closed-source model that is possibly the most capable vision-language model in the world at the time of writing the paper. During the training loop of LICORICE, we query GPT-4o each time the algorithm requires a concept label (line 13 in Algorithm 1). As we are using a pre-specified concept set, we design prompts with detailed general descriptions of the scene layout and definitions of all concepts, in a similar fashion to giving labeling instructions to real humans. We then prompt the GPT-4o to obtain the generated labels. In environments where continuous concept values are required, we ask GPT-4o to give as accurate estimates as possible; in environments with concepts that are discrete and more intuitive to label, we provide clear instructions of how to read the concept numbers according to the input image. More details about our prompts can be found in Section A.3.
4 Experiments
In our experiments, we investigate the following questions:
  1. RQ 1
    Does LICORICE enable both high concept accuracy and high environment reward?
  2. RQ 2
    To what extent does LICORICE enable a reduction in human labeling effort without substantially impacting performance?
  3. RQ 3
    Can LICORICE be used alongside powerful vision-language models to further alleviate the labeling burden?
4.1 Environments
We evaluate our approach on three environments. For each environment, we define an interpretable concept set, where we define concepts to describe properties of objects of interest.
PixelCartPole
In PixelCartPole-v1 [31], the states are images, and the concepts are the original continuous features in the standard CartPole environment: the cart position, the cart velocity, the pole angle, and the pole angular velocity. To obtain the mapping from images to concepts, we use a fixed window of the four most recent images for the temporal concepts, such as the cart velocity. We also include the last action as input to ensure that the information is sufficient to infer concept values. This domain is deceptively difficult due to the temporal concepts.
DoorKey
In DoorKey-7x7 [5], the agent operates in a 5x5 grid [^1] In both of the Minigrid environments, the actual usable area is smaller than what shows in its name, as the outermost layer is a boundary. to pick up an item to unlock the door, then reach the green goal square. The states are fully observable images. We define 6 groups of concepts with 12 concept variables in total: 1) the agent’s position, 2) the agent’s direction, 3) the key’s position or a fixed position (0, 0) outside the grid if the key has been picked up, 4) the door’s position, 5) whether the door is open, and 6) whether the agent can move along each of four directions, i.e., the corresponding cell is empty. Each group contains one or more concept variables. For example, position groups contain x and y coordinates.
DynamicObstacles
In DynamicObstacles-5x5 [5], the agent operates in a 3x3 grid to avoid two moving obstacles while navigating to the goal. Colliding with the dynamic obstacles yields a large penalty, so the agent must correctly learn concepts to safely reach the goal. The states are fully observable images. We define 5 groups of concepts with 11 concept variables in total: 1) the agent’s position, 2) the agent’s direction, 3) the first obstacle’s position, 4) the second obstacle’s position, and 5) whether the agent can move along each of four directions, i.e., the corresponding cell is empty. Each group contains one or more concept variables. For example, position groups contain x and y coordinates.
4.2 Experiment Details
In each experiment, we run each algorithm 5 times, each with a random seed. More implementation details and hyperparameters are in Section A.2.
Architecture
For the backbone model and algorithm, we use PPO (implementation provided by Stable Baselines 3 [23]). We preserve most of the network architecture, but we add a concept layer in the policy network. Specifically, after passing the image through the feature extractor, which consists of 3 CNN layers, we introduce a linear layer to map to the concept layer and obtain predicted concept values. Each predicted concept is represented as a real value for the regression case and one categorical value derived from a classification head for the classification case. This portion of the network corresponds to $g$. For $f$, we utilize an MLP extractor with 2 fully connected layers, each with 64 neurons and a Tanh activation function. We share the default CNN feature extractor between policy and value functions, as we found this to be beneficial in preliminary experiments.
Concept Representation
We model concept learning for PixelCartPole-v1 as a regression problem (minimizing mean squared error) as the concepts are real-valued, and concept learning for DoorKey-7x7 and DynamicObstacles-5x5 as classification problems as the concepts are categorical. The observation images have resolutions of $(240,160)$ for PixelCartPole-v1, and $(160,160)$ for DoorKey-7x7 and DynamicObstacles-5x5. These dimensions maintain a short side of 160 pixels while preserving the aspect ratio of the original rendered images, as we observe lower resolutions may impact performance.
Performance Metrics
To obtain an upper bound on the reward for each environment, we directly use ground-truth concept labels to learn a policy, leading to a reward of $500$, $0.97$, and $0.91$ for the three environments respectively. In the following sections, we report the reward as a ratio of this upper bound. Percentages (or ratios) make sense here since the minimum reasonable reward is 0 in all environments: in PixelCartPole-v1 and DoorKey-7x7, all rewards are nonnegative; in DynamicObstacles-5x5, a random policy would have negative reward due to collisions, but the agent can ensure nonnegative reward by simply staying in place. We additionally report the concept error (MSE for regression; $1$ accuracy for classification). All the reported numbers are calculated during the testing stage, where we evaluate the models on 100 episodes.
4.3 Results
PixelCartPole-v1 DoorKey-7x7 DynamicObstacles-5x5
$c$ Labels Algorithm $R$uparrow $c$ MSE $\downarrow$ $R$uparrow $c$ Error $\downarrow$ $R$uparrow $c$ Error $\downarrow$
$B$ Sequential-Q 0.232 4720.2324720.232\,4720.232 472 0.102 570.102570.102\,570.102 57 0.462 5660.4625660.462\,5660.462 566 0.462 3270.4623270.462\,3270.462 327 0.788 3480.7883480.788\,3480.788 348 0.014 1360.0141360.014\,1360.014 136
Disagreement-Q 0.363 6240.3636240.363\,6240.363 624 0.097 4570.0974570.097\,4570.097 457 0.759 4640.7594640.759\,4640.759 464 0.392 3940.3923940.392\,3940.392 394 0.992 530.992530.992\,530.992 53 0.000 7790.0007790.000\,7790.000 779
Random-Q 0.337 720.337720.337\,720.337 72 0.065 6130.0656130.065\,6130.065 613 0.886 5670.8865670.886\,5670.886 567 0.269 730.269730.269\,730.269 73 0.981 0960.9810960.981\,0960.981 096 0.022 4280.0224280.022\,4280.022 428
LICORICE $1.0$ 0.024 7970.0247970.024\,7970.024 797 0.990 2970.9902970.990\,2970.990 297 0.047 5040.0475040.047\,5040.047 504 0.990 6070.9906070.990\,6070.990 607 0.000 1930.0001930.000\,1930.000 193
$\infty$ CPM  [33] $1.0$ 0.011 5450.0115450.011\,5450.011 545 1.001 0241.0010241.001\,0241.001 024 $0.0$ 0.969 5770.9695770.969\,5770.969 577 $0.0$
Table 1: Evaluation of the reward $R$ and concept error achieved by all methods in all environments. The reward is reported as the fraction of the reward upper bound. Unless otherwise stated, concept error is 1 - accuracy. Methods in the top section are given a budget of $B=[500,300,300]$, respectively; CPM is given an unlimited budget (in practice, it uses 4M, 4M, 1M concept labels respectively). For each column, the top-performing method with a limited budget is in bold. If there are several methods with overlapping standard deviations, we bold all of them. Full results with standard deviation are in Table 6, Appendix B.
4.3.1 Balancing Concept Performance and Environment Reward
We first validate that LICORICE can achieve high reward in all test environments while accurately identifying concepts. We first compare against baselines with a fixed budget of $B=[500,300,300]$ queries for PixelCartPole-v1, DoorKey-7x7, and DynamicObstacles-5x5, respectively, as depicted in the first section of Table 1. We choose these budgets by starting from $500$ and then decreasing by units of $100$ until we find that LICORICE can no longer achieve $99\%$ of the reward upper bound if we further decrease the budget. We treat the minimum budget under which LICORICE can achieve $99\%$ of the reward upper bound as the maximum budget considered for that environment.
Comparison with Budget-Constrained Baselines
To our knowledge, no previous algorithms exist that seek to minimize the number of train-time labels for concept-based RL training, so we implement our own baselines. In Sequential-Q, the agent spends all of $B$ queries on the first $B$ states it encounters during the initial policy rollout. In Disagreement-Q, the agent similarly spends its budget at the beginning of its learning process; however, it uses active learning with the same acquisition function as LICORICE to strategically choose the training data. In Random-Q, the agent receives $B$ concept labels at random points in the training process using a probability to decide whether to query for a concept for each state. We show the results in the first section of Table 1. In the three environments, LICORICE performs similarly to or better than all baselines in terms of both reward and concept error. The main performance differences can be seen for PixelCartPole-v1 and DoorKey-7x7, where LICORICE achieves $100\%$ and $99\%$ of the optimal reward, respectively, while the second best algorithm achieves $36\%$ and $89\%$, respectively. Random-Q, Disagreement-Q, and LICORICE perform similarly on DynamicObstacles-5x5, indicating that this environment is relatively simple and may not benefit from a more advanced strategy. Its simplicity may be due to the lack of concept distribution shift that is present in the other two environments. For example, in PixelCartPole-v1, the policy must be further refined to improve its estimate of the more on-policy concepts. In the initial training stages, when the policy is more random, the agent needs to estimate a wide range of possible values for the pole angle, as the pole starts upright and frequently falls. As the agent’s policy improves, it requires more precise estimates of the pole angle primarily around the vertical position.
Comparison with Budget-Unconstrained Baseline
To better understand the performance of our approach, we compare it against existing work that is not constrained by concept label budgets. Specifically, we implement Concept Policy Model (CPM) from previous work in multi-agent RL [33] but for the single-agent setting. This approach jointly trains the concept bottleneck and the policy, assuming unlimited access to concept labels. It represents an upper bound on concept accuracy, as the agent receives continuous concept feedback throughout learning. Surprisingly, LICORICE outperforms CPM in PixelCartPole-v1and has similar performance to CPM in the other two environments in this unfair comparison. We emphasize that CPM is given an unlimited budget, and in fact, it uses over 1M concept labels for each environment, whereas LICORICE uses $500$ or fewer. We therefore answer RQ 1 affirmatively: compared to baselines, LICORICE demonstrates both high concept accuracy (low concept error) and high reward on our three test environments, all while using substantially fewer concept labels than existing algorithms.
PixelCartPole-v1 DoorKey-7x7 DynamicObstacles-5x5
$B$ 300 400 500 100 200 300 100 200 300
$R$ 0.276 3440.2763440.276\,3440.276 344 0.766 2880.7662880.766\,2880.766 288 $1.0$ 0.765 3910.7653910.765\,3910.765 391 0.934 5320.9345320.934\,5320.934 532 0.990 2970.9902970.990\,2970.990 297 0.957 4390.9574390.957\,4390.957 439 0.979 8190.9798190.979\,8190.979 819 0.990 6070.9906070.990\,6070.990 607
$c$ Error 0.153 5730.1535730.153\,5730.153 573 0.052 4580.0524580.052\,4580.052 458 0.024 7970.0247970.024\,7970.024 797 0.260 4040.2604040.260\,4040.260 404 0.086 6350.0866350.086\,6350.086 635 0.047 5040.0475040.047\,5040.047 504 0.040 7940.0407940.040\,7940.040 794 0.007 8340.0078340.007\,8340.007 834 0.000 1930.0001930.000\,1930.000 193
Table 2: Performance of LICORICE on all environments for varying maximum budgets. The reward is reported as the fraction of the reward upper bound. For PixelCartPole-v1, $c$ error is MSE; for the other environments, $c$ error is 1 - accuracy. For all environments, the reward increases and the error decreases as we increase the maximum budget. Full results with standard deviation are in Table 7Appendix B.
4.3.2 Budget Allocation Effectiveness
To understand the extent to which LICORICE enables a reduction in human labeling effort without substantially impacting performance, we conduct experiments across three different environments with varying concept labeling budgets. We start with the baseline budget $B$ from our previous experiments and test two additional budgets, decreasing by steps of 100. The only component of our algorithm that we vary is the number of iterations $M$. For each budget and environment, we report a single performance value that maximizes the sum of the relative reward and inverse concept error, both scaled to $[0,1]$. The results of this experiment are shown in Table 2.
Our findings reveal that the impact of budget reductions (human labeling effort) varies across environments. In general, we see an increase in concept error and a decrease in reward for all environments with decreasing budget. However, both DynamicObstacles-5x5 and DoorKey-7x7 show more resilience to budget reductions, while PixelCartPole-v1 exhibits higher sensitivity. Specifically, we find that even with $B=100$ for DynamicObstacles-5x5, LICORICE can still achieve $96\%$ of the optimal reward. With $B=200$, LICORICE can still achieve $93\%$ of the optimal reward on DoorKey-7x7. In contrast, we see large decreases in reward from the budget reductions for PixelCartPole-v1. We therefore provide a more nuanced answer to RQ 2: LICORICE can indeed enable reductions in human labeling effort without substantially impacting performance, but the extent of this reduction is environment-dependent. In simpler environments, human effort can be dramatically reduced while maintaining near-optimal performance. However, in more complex environments like PixelCartPole-v1, the trade-off between reduced human effort and maintained performance becomes more pronounced. This finding aligns with the intuition that different environments have varying levels of complexity and would require different levels of human input to accurately capture all relevant concepts and their relationships.
Figure 1: Training curves of reward and concept error for PixelCartPole-v1. Shaded values correspond to 95% CI, calculated using 1000 bootstrap samples. All curves in Figure 5.
4.3.3 Integration with Vision-Language Models
We now seek to answer the question of whether VLMs can successfully provide concept labels in lieu of a human annotator within our LICORICE framework. For consistency, we use the same hyperparameters for LICORICE as in Section 4.3.2. Because using VLMs incurs costs and users requiring interpretable policies for their environments may still face budget constraints, we operate within the same budget-constrained setting as described in Section 4.3.2. Here, the reduction in the number of “human” labels required translates to cost savings.
We show the results in Table 3. We find that GPT-4o can indeed serve as an annotator for some but not all environments. In DoorKey-7x7 and DynamicObstacles-5x5, LICORICE with GPT-4o labels achieves 83% and 88% of the optimal reward, respectively. The concept error rate is comparable to GPT-4o’s labeling error, both evaluated on the same rollout observations, with sampling applied to the latter due to cost constraints. Additionally, the concept error gap between our trained concept model and GPT-4o almost diminishes with more labels. Taken together, these results suggest that LICORICE can work reasonably well with GPT-4o labels.
In PixelCartPole-v1, however, there are challenges for GPT-4o in providing accurate labels due to the continuous nature of the concepts and the lack of necessary knowledge of physical rules and quantities in this specific environment. As a result, LICORICE with GPT-4o labels achieves generally low reward across varying budgets. Interestingly, the concept error increases with more budget, likely due to the additional labeling noise introduced to the concept network. This suggests that while the VLM can handle labeling tasks to some extent, its performance in PixelCartPole-v1 is limited by the complexity of accurately labeling concepts. We also posit the mean reward drop in DynamicObstacles-5x5 at 300 budget is due to the randomness introduced by GPT-4o; however, we note that the range of standard deviation overlaps with the other two budgets. We therefore answer RQ 3 with cautious optimism: there are indeed cases in which LICORICE could be used alongside VLMs. Generally speaking, a more human-intuitive and less complex environment is more likely to work well with VLMs by enabling clear and detailed labeling instructions in the prompt.
PixelCartPole-v1 DoorKey-7x7 DynamicObstacles-5x5
$B$ 300 400 500 100 200 300 100 200 300
$R$ (LICORICE GPT-4o) 0.064 7120.0647120.064\,7120.064 712 0.066 2160.0662160.066\,2160.066 216 0.064 940.064940.064\,940.064 94 0.711 4030.7114030.711\,4030.711 403 0.735 0890.7350890.735\,0890.735 089 0.830 1130.8301130.830\,1130.830 113 0.268 7110.2687110.268\,7110.268 711 0.926 0950.9260950.926\,0950.926 095 0.875 0150.8750150.875\,0150.875 015
$c$ Error (LICORICE GPT-4o) 0.170 0680.1700680.170\,0680.170 068 0.246 0870.2460870.246\,0870.246 087 0.246 895 7330.2468957330.246\,895\,7330.246 895 733 0.450 3260.4503260.450\,3260.450 326 0.329 3480.3293480.329\,3480.329 348 0.312 2880.3122880.312\,2880.312 288 0.200 7110.2007110.200\,7110.200 711 0.133 8850.1338850.133\,8850.133 885 0.132 220.132220.132\,220.132 22
$c$ Error (GPT-4o) 0.193 750.193750.193\,750.193 75 0.224 6830.2246830.224\,6830.224 683 0.241 3650.2413650.241\,3650.241 365 $0.313$ 0.313 6670.3136670.313\,6670.313 667 $0.304$ 0.159 6360.1596360.159\,6360.159 636 0.123 6360.1236360.123\,6360.123 636 0.127 6360.1276360.127\,6360.127 636
Table 3: Performance of LICORICE with GPT-4o integrated into the loop for all environments across different budgets. We also report the concept labeling error on a random sample of 50 observations from the same rollout set used for LICORICE+GPT-4o, due to API cost constraints. Full results with standard deviation are in Table 8, Appendix B.
PixelCartPole-v1 DoorKey-7x7 DynamicObstacles-5x5
Algorithm $R$uparrow $c$ MSE $\downarrow$ $R$uparrow $c$ Error $\downarrow$ $R$uparrow $c$ Error $\downarrow$
LICORICE-IT 0.533 8960.5338960.533\,8960.533 896 0.079 6220.0796220.079\,6220.079 622 0.992 4120.9924120.992\,4120.992 412 0.089 5350.0895350.089\,5350.089 535 0.996 8890.9968890.996\,8890.996 889 0.001 2020.0012020.001\,2020.001 202
LICORICE-DE $0.9669$ 0.031 1060.0311060.031\,1060.031 106 0.783 1130.7831130.783\,1130.783 113 0.200 0220.2000220.200\,0220.200 022 0.980 5370.9805370.980\,5370.980 537 0.000 1140.0001140.000\,1140.000 114
LICORICE-AC 0.910 7840.9107840.910\,7840.910 784 0.022 3140.0223140.022\,3140.022 314 0.815 7550.8157550.815\,7550.815 755 0.161 8420.1618420.161\,8420.161 842 0.583 0530.5830530.583\,0530.583 053 0.002 2910.0022910.002\,2910.002 291
LICORICE $1.0$ 0.024 7970.0247970.024\,7970.024 797 0.990 2970.9902970.990\,2970.990 297 0.047 5040.0475040.047\,5040.047 504 0.990 6070.9906070.990\,6070.990 607 0.000 1930.0001930.000\,1930.000 193
Table 4: Ablation study results for LICORICE in all environments. All of our components generally help achieve a better reward and lower concept error. Full results with standard deviation are in Table 9, Appendix B.
4.3.4 Ablation Study
Given the previous results, we now conduct ablations to confirm the effectiveness of our three main contributions: iterative training, decorrelation, and active learning. LICORICE-IT corresponds to LICORICE with only one iteration, LICORICE-DE corresponds to LICORICE without decorrelation, and LICORICE-AC corresponds to LICORICE without active learning (instead, it uses the entire unlabeled dataset for querying). We show the learning curves for PixelCartPole-v1 in Figure 1; the full set of curves is in Appendix B.
Table 4 depicts the results of our ablation study on all environments. The bottom row corresponds to the upper bound on performance by LICORICE, with all components included. All of our contributions are critical to achieving both high reward and low concept error. However, the component that most contributes to the reward or concept performance differs depending on the environment. For example, compared with LICORICE, LICORICE-IT exhibits the largest reward gap for PixelCartPole-v1; however, LICORICE-AC yields the largest reward gap for DynamicObstacles-5x5, and LICORICE-DE yields the largest gap for DoorKey-7x7. We suspect that this difference is because the concepts in DynamicObstacles-5x5 are simple enough such that one iteration is sufficient for learning, meaning that the largest gains can be made by using active learning. In contrast, PixelCartPole-v1 requires further refinement of the policy to better estimate on-distribution concept values, so the largest gains can be made by leveraging multiple iterations.
5 Related Work
Interpretable Reinforcement Learning Interpretable RL has gained significant attention in recent years [10]. One prominent approach uses rule-based methods — such as decision trees [26, 29], logic [8], and programs [30, 19] — to represent policies. These works either assume that the state is already interpretable or that the policy is pre-specified. Unlike prior work, our method involves learning the interpretable representation (through concept training) for policy learning.
Concept Learning for Reinforcement Learning Inspired by successes in the supervised setting [6, 25, 34], concept-based explanations have recently been incorporated into RL. One approach [7] learns a joint embedding model between state-action pairs and concept-based explanations to expedite learning via reward shaping. Unlike our work, their policy is not a strict function of the concepts, allowing our techniques to be combined to provide both concept-based explanations and a concept-based interpretable policy. Another example, CPM [33], is a multi-agent RL concept architecture that focuses on trade-offs between interpretability and accuracy. They assume that concept labels are available continuously during training. As we have shown, this approach uses over $1$M concept labels in our test environments, whereas our approach reduces the need for continuous human intervention, requiring only 500 or fewer concept labels to achieve similar or better performance in single-agent environments.
Learning Concepts with Human Feedback Another line of work explores how to best leverage human concept labels but not in the RL setting, and does not focus on reducing the labeling burden. In contrast, our approach aims to reduce the concept labeling burden during training. One work [14] instead has users label additional information about the relevance of certain feature dimensions to the concept label. Another work [4] develops an intervention policy at prediction time to choose which concepts to request a label for with the goal of improving the final prediction. Future work could explore using these techniques alongside our method.
6 Discussion and Conclusion
In this work, we proposed LICORICE, a novel RL algorithm that addresses the critical issue of model interpretability while minimizing the reliance on continuous human annotation. We introduced a training scheme that enables RL algorithms to learn concepts more efficiently from little to no labeled concept data. Our approach interleaves concept learning and RL training, uses an ensemble-based active learning technique to select informative data points for labeling, and uses a simple sampling strategy to better decorrelate the concept data. We demonstrated how this approach reduces manual labeling effort. Finally, we conducted initial experiments to demonstrate how we can leverage powerful VLMs to infer concepts from raw visual inputs without explicit labels in some environments. There are broader societal impacts of our work that must be considered. These include both the impacts of using VLMs in real-world applications, as well as considerations around interpretability more generally. For a more detailed discussion, please refer to Appendix C.
Limitations and Future Work Although VLMs can be successfully used for automatic labeling of some concepts, there are still hallucination issues [2] and other failure cases, such as providing inaccurate counts. We believe that future work that seeks to improve general VLM capabilities and mitigate hallucinations would also help overcome this limitation. Another exciting direction is the work in human-AI complementarity and learning-to-defer algorithms [18] to train an additional classifier for deferring concept labeling to a person when the chance of making an error is high.
Acknowledgments and Disclosure of Funding
This work was supported in part by NSF grant IIS-2046640 (CAREER). Co-author Fang is supported in part by Sloan Research Fellowship. We thank NVIDIA for providing computing resources. We thank Zhicheng Zhang for some helpful discussions, Naveen Raman for pointing us to relevant literature, and Rex Chen and Zhiyu (Zoey) Chen for their constructive comments on a draft of this work.
Appendix A Experimental Result Reproducibility
In this section, we provided detailed descriptions to achieve reproducibility.
A.1 Concepts Definitions
Environment Concept Name Type Value Ranges GPT-4o Error
PixelCartPole-v1 Cart Position Continuous $(-2.4,2.4)$ $0.01$\pm$0.0$$
Cart Velocity Continuous $\mathbb{R}$ $0.31$\pm$0.12$$
Pole Angle Continuous (.2095,.2095).2095.2095(-.2095,.2095)( - .2095 , .2095 ) $0.01$\pm$0.0$$
Pole Angular Velocity Continuous $\mathbb{R}$ $0.63$\pm$0.14$$
DoorKey-7x7 Agent Position x Discrete 5 $0.36$\pm$0.08$$
Agent Position y Discrete 5 $0.41$\pm$0.12$$
Agent Direction Discrete 4 $0.28$\pm$0.06$$
Key Position x Discrete 6 $0.2$\pm$0.03$$
Key Position y Discrete 6 $0.32$\pm$0.11$$
Door Position x Discrete 5 $0.37$\pm$0.07$$
Door Position y Discrete 5 $0.32$\pm$0.04$$
Door Open Discrete 2 $0.38$\pm$0.09$$
Direction Movable Right Discrete 2 $0.3$\pm$0.03$$
Direction Movable Down Discrete 2 $0.19$\pm$0.06$$
Direction Movable Left Discrete 2 $0.33$\pm$0.06$$
Direction Movable Up Discrete 2 $0.2$\pm$0.08$$
DynamicObstacles-5x5 Agent Position x Discrete 3 $0.03$\pm$0.06$$
Agent Position y Discrete 3 $0.04$\pm$0.06$$
Agent Direction Discrete 4 $0.2$\pm$0.05$$
Obstacle 1 Position x Discrete 3 $0.08$\pm$0.02$$
Obstacle 1 Position y Discrete 3 $0.19$\pm$0.04$$
Obstacle 2 Position x Discrete 3 $0.22$\pm$0.09$$
Obstacle 2 Position y Discrete 3 $0.12$\pm$0.02$$
Direction Movable Right Discrete 2 $0.19$\pm$0.06$$
Direction Movable Down Discrete 2 $0.14$\pm$0.05$$
Direction Movable Left Discrete 2 $0.13$\pm$0.02$$
Direction Movable Up Discrete 2 $0.07$\pm$0.08$$
Table 5: Concepts and their possible values for all environments. For discrete concepts, we report the number of categories. We also provide the mean GPT-4o labeling error (MSELoss for continuous values and 1 - accuracy for discrete ones) for each single concept, averaged across 5 seeds and corresponding to the same evaluation protocols as the ones with largest budgets in Table 8. The $\pm$ [value] part shows the standard deviation. The average errors over concepts are $0.24$\pm$0.06$$, $0.3$\pm$0.02$$, and $0.13$\pm$0.03$$ respectively for the three environments.
Table 5 provides more details on the concepts used in each environment, categorizing them by their names, types, and value ranges. For the PixelCartPole-v1 environment, all concepts such as Cart Position, Cart Velocity, Pole Angle, and Pole Angular Velocity are continuous. In contrast, the DoorKey-7x7 environment features discrete concepts like Agent Position (x and y), Key Position (x and y), and Door Open status, each with specific value ranges. Similarly, the DynamicObstacles-5x5 environment lists discrete concepts, including Agent and Obstacle positions, with corresponding value ranges.
Figure 2: Example start configurations of the DoorKey-7x7 environment. This shows that concepts must general enough to apply to many different configurations. As a result, the agent cannot memorize the exact position of the door and key.
We also visualize the start configurations for DoorKey-7x7 in Figure 2 to illustrate the importance of concept definitions. As shown, the concepts must be defined in such a way to allow the agent to generalize to all possible environment configurations.
A.2 LICORICE Details
In this section, we provide additional implementation details for LICORICE. The model architecture has been mostly described in the main text and we state all additional details here. The number of neurons in the concept layer is exactly the number of concepts. For continuous concept values, we directly use a linear layer to map from features to concept values. For discrete concept values, since different concepts have different numbers of categories, we create one linear classification head for each single concept, and to predict the final action, we calculate the class with the largest predicted probability for each concept.
Value-Based Methods
If we were to use a value-based method as the RL backbone, we would need to make the following changes. First, we would need to modify $V(s,a)$ or $Q(s,a)$ to include a concept bottleneck, such that $Q(s,a)=f(g(s))$. Then, we can conduct interleaved training in a similar way to LICORICE.
Feature Extractor If we use the actor-critic paradigm, we propose to share a feature extractor between the policy and value networks, shown in Figure 3. Intuitively, this choice can offer several advantages compared with using image or predicted concepts as input for both networks.
Figure 3: Architecture of our concept-bottleneck actor-critic method.
Sharing a feature extractor enables both networks to benefit from a common, rich representation of the input data, reducing the number of parameters to be trained. More importantly, it balances the updates of the policy and value networks. In experiments, we observed that directly using the raw image as input for both networks complicated policy learning. Conversely, relying solely on predicted concepts for the value network may limit its accuracy in value estimation, particularly if the concepts do not capture all the nuances relevant to the value predictions.
A.3 VLM Details
We detail our prompts for each environment here.
[…]
  • Cart Position, within the range (-2.4, 2.4)
  • Cart Velocity
  • Pole Angle, within the range (-0.2095, 0.2095)
  • Pole Angular Velocity
Additionally, please note that the last action taken was [last action].
Please carefully determine the following values and give concise answers one by one. Make sure to return an estimated value for each parameter, even if the task may look challenging.
Follow the reporting format:
  • Cart Position: estimated_value
  • Cart Velocity: estimated_value
  • Pole Angle: estimated_value
  • Pole Angular Velocity: estimated_value
[…]
Detailed Instructions:
  1. 1.
    Agent Position: Identify and report the coordinates (x, y) of the red triangle (agent). Ensure the accuracy by double-checking the agent’s exact location within the grid.
  2. 2.
    Agent Direction: Specify the direction the red triangle is facing, which is the orientation of the vertex (pointy corner) of the isosceles triangle. Choose from ’right’, ’down’, ’left’, or ’up’. Clarify that this direction is independent of movement options.
  3. 3.
    Key Position: Provide the coordinates (x, y) where the key is located. If the key is absent, report as (0, 0). Verify visually that the key is present or not before reporting.
  4. 4.
    Door Position:
    • Position: Determine and report the coordinates (x, y) of the door.
    • Status: Assess whether the door is open or closed (closed means the door is visible as a whole yellow cell, while open means the door disappears visually). Report as ’true’ for open and ’false’ for closed. Double-check the door’s appearance to confirm if it is open or closed.
  5. 5.
    Direction Movable: Evaluate and report whether the agent can move one cell in each specified direction, namely, the neighboring cell in that direction is active and empty (not key, closed door, or grey inactive cell):
    • Right (x + 1): Check the cell to the right.
    • Down (y + 1): Check the cell below.
    • Left (x - 1): Check the cell to the left.
    • Up (y - 1): Check the cell above.
    Each direction’s feasibility should be reported as ’true’ if clear and within the grid, and ’false’ otherwise.
[…]
  1. 1.
    Agent Position: Identify and report the coordinates (x, y) of the red triangle (agent). Ensure the accuracy by double-checking the agent’s exact location within the grid.
  2. 2.
    Agent Direction: Specify the direction the red triangle is facing, which is the orientation of the vertex (pointy corner) of the isosceles triangle. Choose from ’right’, ’down’, ’left’, or ’up’. Clarify that this direction is independent of movement options.
  3. 3.
    Obstacle Position: Identify and report the coordinates of the two obstacles in ascending order. Compare the coordinates by their x-values first. If the x-values are equal, compare by their y-values.
    1. (a)
      First Obstacle: Provide the coordinates (x, y) of the first blue ball.
    2. (b)
      Second Obstacle: Provide the coordinates (x, y) of the second blue ball.
  4. 4.
    Direction Movable: Evaluate and report whether the agent can move one cell in each specified direction, namely, the neighboring cell in that direction is active and empty (not obstacle or out of bounds):
    • Right (x + 1): Check the cell to the right.
    • Down (y + 1): Check the cell below.
    • Left (x - 1): Check the cell to the left.
    • Up (y - 1): Check the cell above.
    Each direction’s feasibility should be reported as ’true’ if clear and within the grid, and ’false’ otherwise.
[…]
A.4 Experimental Details
For the PPO hyperparameters, we set $4\cdot 10^{6}$ total timesteps for PixelCartPole-v1and DoorKey-7x7, and $10^{6}$ for DynamicObstacles-5x5. Besides that, for all environments, we use $8$ vectorized environments, horizon $T=4096$, $10$ epochs for training, batch size of $512$, learning rate $3\cdot 10^{-4}$, entropy coefficient $0.01$, and value function coefficient $1$. For all other hyperparameters, we use the default values from Stable Baselines 3 [23].
For the concept training, we set $100$ epochs with Adam optimizer with the learning rate linearly decaying from $3\cdot 10^{-4}$ to 00 for each iteration in PixelCartPole-v1. In DoorKey-7x7 and DynamicObstacles-5x5, we use the same optimizer and initial learning rate, yet set $50$ epochs instead and set early stopping with threshold linearly increasing from $10$ to $20$, to incentivize the concept network not to overfit in earlier iterations. The batch size is $32$.
For LICORICE, we set the sample acceptance rate $p=0.05$, the ratio for active learning $\tau=10$, batch size to query labels in the active learning module $b=20$, and the number of ensemble models $N=5$. The number of iterations chosen in our algorithm is $M=4$ for PixelCartPole-v1and $M=2$ for the other two environments.
We use NVIDIA A6000 and NVIDIA RTX 6000 Ada Generation. Each of our programs uses less than $2$GB GPU memory. For PixelCartPole-v1 and DoorKey-7x7, each run takes less than 9 hours to finish. For DynamicObstacles-5x5, each run takes less than 2 hours to finish.
For all experiments, we consistently use 5 seeds [$123$, $456$, $789$, $1011$, $1213$] to train the models. We then evaluate on the environment with seed $42$ with 100 episodes and take the average of the reward across the episodes and the concept error across the observations within each episode.
Appendix B Additional Results
B.1 Balancing Concept Performance and Environment Reward
Algorithm $R$uparrow $c$ Error $\downarrow$
PixelCartPole-v1 Sequential-Q 0.232 472±0.085 242plus-or-minus0.2324720.085242\$0.232\,472$\pm$\$0.085\,242\$0.232 472 ± 0.085 242 0.102 57±0.037 255plus-or-minus0.102570.037255\$0.102\,57$\pm$\$0.037\,255\$0.102 57 ± 0.037 255
Disagreement-Q 0.363 624±0.200 067plus-or-minus0.3636240.200067\$0.363\,624$\pm$\$0.200\,067\$0.363 624 ± 0.200 067 0.097 457±0.036 755plus-or-minus0.0974570.036755\$0.097\,457$\pm$\$0.036\,755\$0.097 457 ± 0.036 755
Random-Q 0.337 72±0.137 783plus-or-minus0.337720.137783\$0.337\,72$\pm$\$0.137\,783\$0.337 72 ± 0.137 783 0.065 613±0.019 76plus-or-minus0.0656130.01976\$0.065\,613$\pm$\$0.019\,76\$0.065 613 ± 0.019 76
LICORICE $1.0$\pm$0.0$$ 0.024 797±0.004 166plus-or-minus0.0247970.004166\$0.024\,797$\pm$\$0.004\,166\$0.024 797 ± 0.004 166
CPM $1.0$\pm$0.0$$ 0.011 545±0.000 782plus-or-minus0.0115450.000782\$0.011\,545$\pm$\$0.000\,782\$0.011 545 ± 0.000 782
DoorKey-7x7 Sequential-Q 0.462 566±0.091 999plus-or-minus0.4625660.091999\$0.462\,566$\pm$\$0.091\,999\$0.462 566 ± 0.091 999 0.462 327±0.045 955plus-or-minus0.4623270.045955\$0.462\,327$\pm$\$0.045\,955\$0.462 327 ± 0.045 955
Disagreement-Q 0.759 464±0.117 38plus-or-minus0.7594640.11738\$0.759\,464$\pm$\$0.117\,38\$0.759 464 ± 0.117 38 0.392 394±0.073 599plus-or-minus0.3923940.073599\$0.392\,394$\pm$\$0.073\,599\$0.392 394 ± 0.073 599
Random-Q 0.886 567±0.033 546plus-or-minus0.8865670.033546\$0.886\,567$\pm$\$0.033\,546\$0.886 567 ± 0.033 546 0.269 73±0.082 918plus-or-minus0.269730.082918\$0.269\,73$\pm$\$0.082\,918\$0.269 73 ± 0.082 918
LICORICE 0.990 297±0.007 106plus-or-minus0.9902970.007106\$0.990\,297$\pm$\$0.007\,106\$0.990 297 ± 0.007 106 0.047 504±0.011 097plus-or-minus0.0475040.011097\$0.047\,504$\pm$\$0.011\,097\$0.047 504 ± 0.011 097
CPM 1.001 024±0.000 06plus-or-minus1.0010240.00006\$1.001\,024$\pm$\$0.000\,06\$1.001 024 ± 0.000 06 $0.0$\pm$0.0$$
DynamicObstacles-5x5 Sequential-Q 0.788 348±0.440 724plus-or-minus0.7883480.440724\$0.788\,348$\pm$\$0.440\,724\$0.788 348 ± 0.440 724 0.014 136±0.009 627plus-or-minus0.0141360.009627\$0.014\,136$\pm$\$0.009\,627\$0.014 136 ± 0.009 627
Disagreement-Q 0.992 53±0.014 194plus-or-minus0.992530.014194\$0.992\,53$\pm$\$0.014\,194\$0.992 53 ± 0.014 194 0.000 779±0.000 458plus-or-minus0.0007790.000458\$0.000\,779$\pm$\$0.000\,458\$0.000 779 ± 0.000 458
Random-Q 0.981 096±0.006 375plus-or-minus0.9810960.006375\$0.981\,096$\pm$\$0.006\,375\$0.981 096 ± 0.006 375 0.022 428±0.013 904plus-or-minus0.0224280.013904\$0.022\,428$\pm$\$0.013\,904\$0.022 428 ± 0.013 904
LICORICE 0.990 607±0.003 563plus-or-minus0.9906070.003563\$0.990\,607$\pm$\$0.003\,563\$0.990 607 ± 0.003 563 0.000 193±0.000 432plus-or-minus0.0001930.000432\$0.000\,193$\pm$\$0.000\,432\$0.000 193 ± 0.000 432
CPM 0.969 577±0.021 855plus-or-minus0.9695770.021855\$0.969\,577$\pm$\$0.021\,855\$0.969 577 ± 0.021 855 $0.0$\pm$0.0$$
Table 6: Evaluation of the reward $R$ and concept error achieved by all methods in all environments. This is a extended table from Table 1. The reward is reported as the fraction of the reward upper bound. For PixelCartPole-v1, the $c$ error is the MSE. For the other two environments, the $c$ error is 1 - accuracy. The first four algorithms are given a budget of $B=[500,300,300]$ for each environment, from top to bottom; CPM is given an unlimited budget (in practice, it uses 4M, 4M, 1M concept labels respectively). The $\pm$ [value] part shows the standard deviation. This shows a more complete version of the results in Table 1.
In Table 6, we present an extension of the results in Table 1, including standard deviation. Our method enjoys low variance across all environments in terms of both concept error and reward.
B.2 Budget Allocation Effectiveness
$B$ $M$ $R$uparrow $c$ Error $\downarrow$
PixelCartPole-v1 300 1 0.276 344±0.104 789plus-or-minus0.2763440.104789\$0.276\,344$\pm$\$0.104\,789\$0.276 344 ± 0.104 789 0.153 573±0.065 982plus-or-minus0.1535730.065982\$0.153\,573$\pm$\$0.065\,982\$0.153 573 ± 0.065 982
400 3 0.766 288±0.224 139plus-or-minus0.7662880.224139\$0.766\,288$\pm$\$0.224\,139\$0.766 288 ± 0.224 139 0.052 458±0.012 336plus-or-minus0.0524580.012336\$0.052\,458$\pm$\$0.012\,336\$0.052 458 ± 0.012 336
500 4 $1.0$\pm$0.0$$ 0.024 797±0.004 166plus-or-minus0.0247970.004166\$0.024\,797$\pm$\$0.004\,166\$0.024 797 ± 0.004 166
DoorKey-7x7 100 1 0.765 391±0.039 967plus-or-minus0.7653910.039967\$0.765\,391$\pm$\$0.039\,967\$0.765 391 ± 0.039 967 0.260 404±0.041 564plus-or-minus0.2604040.041564\$0.260\,404$\pm$\$0.$0 4$ 0.041 564
200 2 0.934 532±0.015 966plus-or-minus0.9345320.015966\$0.934\,532$\pm$\$0.015\,966\$0.934 532 ± 0.015 966 0.086 635±0.009 289plus-or-minus0.0866350.009289\$0.086\,635$\pm$\$0.009\,289\$0.086 635 ± 0.009 289
300 2 0.990 297±0.007 106plus-or-minus0.9902970.007106\$0.990\,297$\pm$\$0.007\,106\$0.990 297 ± 0.007 106 0.047 504±0.011 097plus-or-minus0.0475040.011097\$0.047\,504$\pm$\$0.011\,097\$0.047 504 ± 0.011 097
DynamicObstacles-5x5 100 1 0.957 439±0.020 369plus-or-minus0.9574390.020369\$0.957\,439$\pm$\$0.020\,369\$0.957 439 ± 0.020 369 0.040 794±0.006 774plus-or-minus0.0407940.006774\$0.040\,794$\pm$\$0.006\,774\$0.040 794 ± 0.006 774
200 1 0.979 819±0.014 765plus-or-minus0.9798190.014765\$0.979\,819$\pm$\$0.014\,765\$0.979 819 ± 0.014 765 0.007 834±0.004 337plus-or-minus0.0078340.004337\$0.007\,834$\pm$\$0.004\,337\$0.007 834 ± 0.004 337
300 2 0.990 607±0.003 563plus-or-minus0.9906070.003563\$0.990\,607$\pm$\$0.003\,563\$0.990 607 ± 0.003 563 0.000 193±0.000 432plus-or-minus0.0001930.000432\$0.000\,193$\pm$\$0.000\,432\$0.000 193 ± 0.000 432
Table 7: Performance of LICORICE on all environments for varying budgets. We select $M$ through grid search, optimizing $R$ $c$ error without additional weighting since they are observed to have the same magnitude. The reward is reported as the fraction of the reward upper bound. For PixelCartPole-v1, $c$ error is MSE; for the other environments, $c$ error is 1 - accuracy. The $\pm$ [value] part shows the standard deviation. This shows a more complete version of the results in Table 2.
In Table 7, we present an extension of the results in Table 2, including the standard deviation. As expected, as the budget increases, the standard deviation for both the reward and concept error tends to decrease. The one exception is the reward for PixelCartPole-v1. Interestingly, the standard deviation is highest for $B=400$. We suspect this is because the concept errors may be more critical here, leading to higher variance in the reward performance.
Environment $B$ $M$ $R$uparrow LICORICE+GPT-4o $c$ Error $\downarrow$ GPT-4o $c$ Error
PixelCartPole-v1 300 1 0.064 712±0.017 953plus-or-minus0.0647120.017953\$0.064\,712$\pm$\$0.017\,953\$0.064 712 ± 0.017 953 0.170 068±0.188 407plus-or-minus0.1700680.188407\$0.170\,068$\pm$\$0.188\,407\$0.170 068 ± 0.188 407 0.193 75±0.248 575plus-or-minus0.193750.248575\$0.193\,75$\pm$\$0.248\,575\$0.193 75 ± 0.248 575
400 2 0.066 216±0.021 619plus-or-minus0.0662160.021619\$0.066\,216$\pm$\$0.021\,619\$0.066 216 ± 0.021 619 0.246 087±0.147 974plus-or-minus0.2460870.147974\$0.246\,087$\pm$\$0.147\,974\$0.246 087 ± 0.147 974 0.224 683±0.151 258plus-or-minus0.2246830.151258\$0.224\,683$\pm$\$0.151\,258\$0.224 683 ± 0.151 258
500 2 0.064 94±0.010 829 728plus-or-minus0.064940.010829728\$0.064\,94$\pm$\$0.010\,829\,728\$0.064 94 ± 0.010 829 728 0.246 895 733±0.077 473 22plus-or-minus0.2468957330.07747322\$0.246\,895\,733$\pm$\$0.077\,473\,22\$0.246 895 733 ± 0.077 473 22 0.241 365±0.069 363plus-or-minus0.2413650.069363\$0.241\,365$\pm$\$0.069\,363\$0.241 365 ± 0.069 363
DoorKey-7x7 100 1 0.711 403±0.042 15plus-or-minus0.7114030.04215\$0.711\,403$\pm$\$0.042\,15\$0.711 403 ± 0.042 15 0.450 326±0.036 168plus-or-minus0.4503260.036168\$0.450\,326$\pm$\$0.036\,168\$0.450 326 ± 0.036 168 0.313±0.034 549plus-or-minus0.3130.034549\$0.313$\pm$\$0.034\,549\$0.313 ± 0.034 549
200 2 0.735 089±0.043 674plus-or-minus0.7350890.043674\$0.735\,089$\pm$\$0.043\,674\$0.735 089 ± 0.043 674 0.329 348±0.010 042plus-or-minus0.3293480.010042\$0.329\,348$\pm$\$0.010\,042\$0.329 348 ± 0.010 042 0.313 667±0.031 697plus-or-minus0.3136670.031697\$0.313\,667$\pm$\$0.031\,697\$0.313 667 ± 0.031 697
300 2 0.830 113±0.056 988plus-or-minus0.8301130.056988\$0.830\,113$\pm$\$0.056\,988\$0.830 113 ± 0.056 988 0.312 288±0.014 362plus-or-minus0.3122880.014362\$0.312\,288$\pm$\$0.014\,362\$0.312 288 ± 0.014 362 0.304±0.027 654plus-or-minus0.3040.027654\$0.304$\pm$\$0.027\,654\$0.304 ± 0.027 654
DynamicObstacles-5x5 100 1 0.268 711±0.373 027plus-or-minus0.2687110.373027\$0.268\,711$\pm$\$0.373\,027\$0.268 711 ± 0.373 027 0.200 711±0.0328plus-or-minus0.2007110.0328\$0.200\,711$\pm$\$0.0328\$0.200 711 ± 0.0328 0.159 636±0.029 329plus-or-minus0.1596360.029329\$0.159\,636$\pm$\$0.029\,329\$0.159 636 ± 0.029 329
200 1 0.926 095±0.048 753plus-or-minus0.9260950.048753\$0.926\,095$\pm$\$0.048\,753\$0.926 095 ± 0.048 753 0.133 885±0.029 764plus-or-minus0.1338850.029764\$0.133\,885$\pm$\$0.029\,764\$0.133 885 ± 0.029 764 0.123 636±0.046 069plus-or-minus0.1236360.046069\$0.123\,636$\pm$\$0.046\,069\$0.123 636 ± 0.046 069
300 1 0.875 015±0.088 616plus-or-minus0.8750150.088616\$0.875\,015$\pm$\$0.088\,616\$0.875 015 ± 0.088 616 0.132 22±0.025 588plus-or-minus0.132220.025588\$0.132\,22$\pm$\$0.025\,588\$0.132 22 ± 0.025 588 0.127 636±0.029 385plus-or-minus0.1276360.029385\$0.127\,636$\pm$\$0.029\,385\$0.127 636 ± 0.029 385
Table 8: Performance of LICORICE with GPT-4o integrated into the loop for all environments across different budgets, along with the concept labeling error of GPT-4o as a reference. We select $M$ through grid search, optimizing $R$ $c$ error without additional weighting since they are observed to have the same magnitude. GPT-4o $c$ error is evaluated on a random sample of 50 observations from the same rollout set used for LICORICE+GPT-4o, due to API cost constraints. This shows a more complete version of the results in Table 3.
B.3 Integration with Vision-Language Models
In Table 8, we present an extension of the results in Table 3 in the main paper, including the standard deviation. Interestingly, the standard deviation for the reward obtained by using LICORICE with GPT-4o as the annotator does not always follow the same trend as shown in Table 7 (when we assume access to a more accurate human annotator). Instead, the standard deviation is relatively consistent for PixelCartPole-v1, regardless of the budget. It steadily decreases for DoorKey-7x7, as expected. However, in DynamicObstacles-5x5, we see an increase when $B=300$. We are not sure of the cause of this. Perhaps at this point the algorithm begins overfitting to the errors in the labels from GPT-4o (the concept error rate is the same for $200$ and $300$ labels). Further investigation is required to understand the underlying factors contributing to this anomaly.
Figure 4: Concept policy model trained with concept labels provided by GPT-4o waits until the coast is clear to move to the goal. It appears to make a small mistake in the bottom left, requiring it to wait slightly longer than necessary to navigate to the goal.
GPT-4o Concept Labeling Errors Table 5 list detailed concept errors for concepts in all environments. In PixelCartPole-v1, cart position and pole angle have smaller errors, while velocities require understanding multiple frames and thus are harder to predict accurately. For DoorKey-7x7 and DynamicObstacles-5x5, different concepts have slightly varying concept errors, indicating visual tasks have different difficulties for GPT-4o. Agent direction has as high as around 0.3 prediction error for both DoorKey-7x7 and DynamicObstacles-5x5. Direction movable is also hard for DoorKey-7x7, with a high concept error even if it is a binary concept. We posit it requires the correct understanding of more than one particular object to ensure correctness. The concept accuracies in DoorKey-7x7 are generally higher than DynamicObstacles-5x5, suggesting GPT-4o more struggles with a larger grid.
Example Behavior of Agent with GPT-4o-labeled concepts In Figure 4, we show an example of a GPT-4o-trained RL agent on DynamicObstacles-5x5, in which the agent appears to wait until it is safe to move towards the goal: the green square. The image sequence shows the agent (red triangle) starting from its initial position and moving to the right. It then is cornered by an obstacle (blue circle), then both obstacles. In the bottom left corner frame, it appears to make a mistake by turning to the right, meaning it missed a window to escape. Finally, it moves to the goal when the path is clear in the second-to-last and last frames (bottom right). This behavior highlights that the agent may still learn reasonable behavior even if the concept labels may be incorrect (causing it to make a mistake, as in the bottom left frame).
B.4 Ablation
Algorithm $R$uparrow $c$ Error $\downarrow$
PixelCartPole-v1 LICORICE-IT 0.533 896±0.257 169plus-or-minus0.5338960.257169\$0.533\,896$\pm$\$0.257\,169\$0.533 896 ± 0.257 169 0.079 622±0.026 734plus-or-minus0.0796220.026734\$0.079\,622$\pm$\$0.026\,734\$0.079 622 ± 0.026 734
LICORICE-DE 0.9669±0.051 17plus-or-minus0.96690.05117\$0.9669$\pm$\$0.051\,17\$0.9669 ± 0.051 17 0.031 106±0.010 96plus-or-minus0.0311060.01096\$0.031\,106$\pm$\$0.010\,96\$0.031 106 ± 0.010 96
LICORICE-AC 0.910 784±0.096 086plus-or-minus0.9107840.096086\$0.910\,784$\pm$\$0.096\,086\$0.910 784 ± 0.096 086 0.022 314±0.007 917plus-or-minus0.0223140.007917\$0.022\,314$\pm$\$0.007\,917\$0.022 314 ± 0.007 917
LICORICE $1.0$\pm$0.0$$ 0.024 797±0.004 166plus-or-minus0.0247970.004166\$0.024\,797$\pm$\$0.004\,166\$0.024 797 ± 0.004 166
DoorKey-7x7 LICORICE-IT 0.992 412±0.007 62plus-or-minus0.9924120.00762\$0.992\,412$\pm$\$0.007\,62\$0.992 412 ± 0.007 62 0.089 535±0.015 02plus-or-minus0.0895350.01502\$0.089\,535$\pm$\$0.015\,02\$0.089 535 ± 0.015 02
LICORICE-DE 0.783 113±0.079 936plus-or-minus0.7831130.079936\$0.783\,113$\pm$\$0.079\,936\$0.783 113 ± 0.079 936 0.200 022±0.051 719plus-or-minus0.2000220.051719\$0.200\,022$\pm$\$0.051\,719\$0.200 022 ± 0.051 719
LICORICE-AC 0.815 755±0.096 997plus-or-minus0.8157550.096997\$0.815\,755$\pm$\$0.096\,997\$0.815 755 ± 0.096 997 0.161 842±0.057 998plus-or-minus0.1618420.057998\$0.161\,842$\pm$\$0.057\,998\$0.161 842 ± 0.057 998
LICORICE 0.990 297±0.007 106plus-or-minus0.9902970.007106\$0.990\,297$\pm$\$0.007\,106\$0.990 297 ± 0.007 106 0.047 504±0.011 097plus-or-minus0.0475040.011097\$0.047\,504$\pm$\$0.011\,097\$0.047 504 ± 0.011 097
DynamicObstacles-5x5 LICORICE-IT 0.996 889±0.002 139plus-or-minus0.9968890.002139\$0.996\,889$\pm$\$0.002\,139\$0.996 889 ± 0.002 139 0.001 202±0.000 982plus-or-minus0.0012020.000982\$0.001\,202$\pm$\$0.000\,982\$0.001 202 ± 0.000 982
LICORICE-DE 0.980 537±0.008 092plus-or-minus0.9805370.008092\$0.980\,537$\pm$\$0.008\,092\$0.980 537 ± 0.008 092 0.000 114±0.000 18plus-or-minus0.0001140.00018\$0.000\,114$\pm$\$0.000\,18\$0.000 114 ± 0.000 18
LICORICE-AC 0.583 053±0.532 307plus-or-minus0.5830530.532307\$0.583\,053$\pm$\$0.532\,307\$0.583 053 ± 0.532 307 0.002 291±0.002 759plus-or-minus0.0022910.002759\$0.002\,291$\pm$\$0.002\,759\$0.002 291 ± 0.002 759
LICORICE 0.990 607±0.003 563plus-or-minus0.9906070.003563\$0.990\,607$\pm$\$0.003\,563\$0.990 607 ± 0.003 563 0.000 193±0.000 432plus-or-minus0.0001930.000432\$0.000\,193$\pm$\$0.000\,432\$0.000 193 ± 0.000 432
Table 9: Ablation study results for LICORICE in all environments, where we delete every single component to compare with LICORICE. This shows a more complete version of the results in Table 4.
Figure 5: Learning curves for all ablations. Shaded region shows 95% CI, calculated using 1000 bootstrap samples.
In Table 9, we present an extension of the results in Table 4, including standard deviation. Figure 5 shows all learning curves for ablations in all environments. In PixelCartPole-v1, we clearly see the benefit of iterative strategies on reward: LICORICE, LICORICE-DE and LICORICE-AC consistently increase in reward and converge at high levels, but LICORICE-IT converges at less than $50\%$ of the optimal reward. We also see that LICORICE-IT also achieves higher concept error, indicating that it struggles to learn the larger distribution of concepts induced by a non-optimal policy. In DoorKey-7x7, all algorithms steadily increase in reward. However, we can see a dip at around $10^{6}$ environment steps where we begin the second iteration for LICORICE, LICORICE-AC, and LICORICE-DE. Although LICORICE-IT achieves similar reward to LICORICE, LICORICE achieves lower concept error, which is beneficial for the goal of interpretability. Finally, in DynamicObstacles-5x5, LICORICE-AC lags behind the most in terms of both reward and concept error. This result indicates that the active learning component is most important for this environment.
Figure 6: Performance of all ablations at the final training iteration, with black bars representing the 95% confidence interval, calculated using 1000 bootstrap samples.
We now inspect the performance at the last training iteration in Figure 6. Overall, we find that LICORICE performs better than or equal to the ablations on all environments. It enjoys the best reward performance on PixelCartPole-v1, while performing similarly to LICORICE-IT on DoorKey-7x7 and to LICORICE-IT and LICORICE-DE on DynamicObstacles-5x5. It exhibits the lowest concept error on PixelCartPole-v1 and DoorKey-7x7, and achieves similarly low error to LICORICE-DE and LICORICE-IT on DynamicObstacles-5x5.
In PixelCartPole-v1, LICORICE achieves the highest reward, indicating the benefits of multiple iterations, active learning, and decorrelation, which are absent in LICORICE-IT, LICORICE-AC, and LICORICE-DE, respectively. LICORICE also enjoys the lowest concept error, showcasing its ability to learn and apply concepts accurately. LICORICE-AC and LICORICE-DE perform well in minimizing concept error, while LICORICE-IT shows a relatively higher error rate, underscoring the importance of multiple iterations in some environments.
Appendix C Additional Discussion
C.1 Limitations and Future Work
While our approach has demonstrated promising results, there are several limitations to be addressed in future work. One significant challenge is the difficulty VLMs face with certain types of concepts, especially continuous variables. This limitation can impact the overall performance of concept-based models, especially in domains where continuous data is prevalent. Addressing this issue could involve developing specialized techniques or using existing tools and libraries to better complement VLM capabilities.
Another area for future improvement is the refinement of our active learning and sampling schemes. Our current method employs an disagreement-based acquisition function to select the most informative data points for labeling. While this approach is effective, there is potential for exploring more sophisticated active learning strategies, such as incorporating advanced exploration-exploitation trade-offs or leveraging recent advancements in active learning algorithms [28].
Finally, designing a concept-based representation for RL remains an open challenge. Our work provides a few illustrative examples, but the exact design of these representations can significantly impact performance, often for reasons that are not entirely clear — especially when using VLMs as annotators. Prior work [7] proposed some desiderata for concepts in RL, but future work could refine these principles, especially in the face of VLM annotators. Future work could also include systematically investigating the factors that influence the effectiveness of different concept-based representations in RL. This could involve extensive empirical studies, theoretical analyses, and the development of new design principles that guide the creation of effective concept representations. Understanding these factors better will help in creating more reliable and interpretable RL models, ultimately advancing the field and broadening the applicability of concept-based approaches in various RL tasks.
C.2 Broader Impacts
Interpretability in RL
Incorporating concept learning with RL presents both positive and negative societal impacts. On the positive side, promoting interpretability and transparency in decision-making fosters trust and accountability. However, in cases where it yields incorrect results, stakeholders might be misled into trusting flawed decisions due to the perceived transparency of the model [12]. Unintended misuse could also occur if stakeholders lack the technical expertise to accurately interpret the models, leading to erroneous conclusions and potentially harmful outcomes. To mitigate these risks, an avenue for future work is developing clear guidelines for interpreting these models and tools to scaffold non-experts’ understanding of the model outputs.
Using VLMs for Concept Labeling
On one hand, VLMs have the potential to significantly improve the efficiency and scalability of labeling processes, which can accelerate advancements in various fields. By automating the labeling of large datasets, VLMs can help reduce the time and cost associated with manual labeling. However, there are important ethical and social considerations to address. One major concern is the potential for bias in the concept labels generated by VLMs. If these models are trained on biased or unrepresentative data, they may perpetuate or even amplify existing biases, leading to unfair or discriminatory outcomes. This is particularly problematic in sensitive applications like hiring, lending, or law enforcement, where biased decisions can have significant negative impacts on individuals and communities. Furthermore, there are privacy concerns related to the data used to train VLMs. Large-scale data collection often involves personal information, and improper handling of this data can lead to privacy violations. To mitigate these risks, future work could include developing robust data governance frameworks to protect individuals’ privacy and comply with relevant regulations.

### v2

@@
-Concept-Based Interpretable Reinforcement Learning with Limited to No Human Labels
+LICORICE: Label-Efficient Concept-Based Interpretable Reinforcement Learning
 Abstract
-Recent advances in reinforcement learning (RL) have predominantly leveraged neural network-based policies for decision-making, yet these models often lack interpretability, posing challenges for stakeholder comprehension and trust.
-Concept bottleneck models offer an interpretable alternative by integrating human-understandable concepts into neural networks.
-However, a significant limitation in prior work is the assumption that human annotations for these concepts are readily available during training, necessitating continuous real-time input from human annotators.
-To overcome this limitation, we introduce a novel training scheme that enables RL algorithms to efficiently learn a concept-based policy by only querying humans to label a small set of data, or in the extreme case, without any human labels.
-Our algorithm, LICORICE, involves three main contributions: interleaving concept learning and RL training, using a concept ensembles to actively select informative data points for labeling, and decorrelating the concept data with a simple strategy.
-We show how LICORICE reduces manual labeling efforts to to 500 or fewer concept labels in three environments.
-Finally, we present an initial study to explore how we can use powerful vision-language models to infer concepts from raw visual inputs without explicit labels at minimal cost to performance.
+Recent advances in reinforcement learning (RL) have predominantly leveraged neural network policies for decision-making, yet these models often lack interpretability, posing challenges for stakeholder comprehension and trust.
+Concept bottleneck models offer an interpretable alternative by integrating human-understandable concepts into policies.
+However, prior work assumes that concept annotations are readily available during training.
+For RL, this requirement poses a significant limitation: it necessitates continuous real-time concept annotation, which either places an impractical burden on human annotators or incurs substantial costs in API queries and inference time when employing automated labeling methods.
+To overcome this limitation, we introduce a novel training scheme that enables RL agents to efficiently learn a concept-based policy by only querying annotators to label a small set of data.
+Our algorithm, LICORICE, involves three main contributions: interleaving concept learning and RL training, using an ensemble to actively select informative data points for labeling, and decorrelating the concept data.
+We show how LICORICE reduces human labeling efforts to 500 or fewer concept labels in three environments, and 5000 or fewer in two more complex environments, all at no cost to performance.
+We also explore the use of VLMs as automated concept annotators, finding them effective in some cases but imperfect in others.
+Our work significantly reduces the annotation burden for interpretable RL, making it more practical for real-world applications that necessitate transparency.
+Our code is released.
+[^1] https://github.com/cuizhuyefei/LICORICE
 1 Introduction
-In reinforcement learning (RL), agents are tasked with learning a policy, a rule that makes sequential, reactive decisions in complex environments.
-In recent RL work, agents typically represent the policy as a neural network, as such representations tend to lead to high performance [17].
-However, this choice can come at a cost: such policies are challenging for stakeholders to interpret — particularly when the inputs to the network are also complex, such as high-dimensional sensor data.
-This opacity can become a significant hurdle, especially in applications where understanding the rationale behind decisions is critical, such as healthcare [32] or finance [15].
-In such applications, decisions can have significant consequences, so it is essential for stakeholders to fully grasp the reasoning behind actions in order to confidently adopt or collaborate on a policy.
-To address interpretability concerns in the supervised learning setting, recent works have integrated human-understandable concepts into the decision-making process through concept bottleneck models [13, 9].
-These models incorporate a bottleneck layer whose units correspond to interpretable concepts, ensuring that the final decisions depend on these concepts instead of just on opaque raw inputs.
-By training the model both to have high accuracy on the target task and to accurately match experts’ concept labels, these models learn a high-level concept-based representation that is simultaneously meaningful to humans and useful for machine learning tasks.
+In reinforcement learning (RL), agents learn a policy, which is a strategy for making sequential decisions in complex environments.
+Agents typically represent the policy as a neural network, as this representation tends to yield high performance (Mirhoseini et al., 2021).
+However, this choice can come at a cost: such policies are challenging for stakeholders to interpret — particularly when the network inputs are also complex, such as high-dimensional sensor data.
+This opacity can pose a significant hurdle, especially in applications where understanding the rationale behind decisions is critical, such as healthcare (Yu et al., 2021) or finance (Liu et al., 2022).
+In such applications, decisions can have significant consequences, so it is essential for stakeholders to fully grasp the reasoning behind actions to confidently adopt or collaborate on a policy.
+To address interpretability concerns in supervised learning, recent works have integrated human-understandable concepts into neural networks through concept bottleneck models (Koh et al., 2020; Espinosa Zarlenga et al., 2022).
+These models insert a bottleneck layer whose units correspond to interpretable concepts, ensuring that the final decisions depend on these concepts instead of on opaque raw inputs.
+By training the model both to have high task accuracy and to accurately match experts’ concept labels, these models learn a high-level concept-based representation that is simultaneously meaningful to humans and useful for machine learning tasks.
 As an example, a concept-based explanation for a bird classification task might include a unit that encodes the bird’s wing color.
-More recently, these techniques have been applied to RL by incorporating a concept bottleneck in the policy [11, 33], such that the actions taken by the agent are a function of the human-understandable concepts.
-However, past work assumes that human-provided concept annotations are accessible in the training loop of RL, which presents a significant challenge.
-In novel domains, we may only have access to data that is not in an interpretable representation, such as RGB or multispectral satellite imagery, while the RL agent needs to know the concepts present in every state and action it encounters during training — even though RL training sets can often be measured in millions or billions of state-action pairs.
-As a result, the human labelers would need to be actively involved in the RL training process, providing concept labels for unreasonable numbers of observations.
-This level of involvement is not only detrimental to the human labeler but also risks introducing potential biases and errors in the model training process due to fatigue [16].
-In this work, we aim to address this challenge of requiring frequent human interventions during the training of interpretable policies using RL.
-We propose LICORICE (Label-efficient Interpretable COncept-based ReInforCEment learning), a novel training paradigm consisting of three main contributions.
-First, LICORICE interleaves concept learning and RL training: it alternately freezes the network layers corresponding to either the concept learning part or the decision-making part.
-We believe this scheme improves our ability to learn from limited labeled data by reducing interference between the learning tasks.
-Additionally, concept learning uses data that is more on-policy, which means it is more relevant and useful for learning accurate concepts that directly impact the decision-making process.
-Second, LICORICE utilizes concept ensembles to actively select the most informative data points for labeling.
-By focusing on samples that are predicted to provide the most valuable information for model improvement, this technique substantially reduces the number of labels needed to achieve both high performance and high concept accuracy.
-Third, LICORICE includes a strategy to decorrelate the concept data collected under the current policy.
-By generating a diverse set of training data, this approach ensures the data remains unbiased and representative of various scenarios.
-We demonstrate how these changes yield both higher concept accuracy and higher reward while requiring fewer queries on three environments with image input, including an image input version of CartPole and two Minigrid environments.
-Given these results, we ask whether we can further reduce the reliance on manual concept labeling by leveraging the potential of vision-language models (VLMs).
-This capability is important in scenarios where manual concept annotation is impractical.
-We present an initial exploration of whether VLMs can further reduce the concept annotation burden.
-We find that VLMs can indeed serve as concept annotators for some but not all of the above environments.
-In these cases, the resulting policies can achieve $83$ to $93$% of the optimal performance.
+Concepts also offer a path to compositional generalization by enabling subtask decomposition into meaningful, reusable components (Mao et al., 2022; Wang et al., 2023).
+More recently, these techniques have been applied to RL by incorporating a concept bottleneck in the policy (Grupen et al., 2022; Zabounidis et al., 2023), so that the chosen actions are a function of the human-understandable concepts.
+However, a significant challenge emerges in this method’s practical application: past work assumes that concept annotations are readily available during RL training.
+To learn a mapping from states to concepts, an RL agent requires concept information for every state encountered during training, which is often millions or billions of state-action pairs.
+Because many domains do not automatically provide these concept labels (e.g., autonomous driving), we require a means of obtaining them.
+Relying on human labelers is impractical, risking errors due to fatigue (Marshall & Shipman, 2013).
+Using vision language models (VLMs) for automated concept extraction (Oikarinen et al., 2023), while alleviating the human bottleneck, introduces significant API or inference costs that can be prohibitive.
+Figure 1: LICORICE overview.
+In concept-based RL, the policy first maps from states to the concepts in a bottleneck layer with $g$, and then maps from concepts to (distributions over) actions with $f$.
+During training, LICORICE addresses concept label efficiency concerns with three key components: i) iterative training, ii) data decorrelation, and iii) active learning.
+In this work, we address this challenge with LICORICE (Label-efficient Interpretable COncept-based ReInforCEment learning), a novel training paradigm designed to minimize the number of concept annotation queries while maintaining high task performance.
+Figure 1 illustrates our algorithm.
+LICORICE tackles three key challenges.
+First, it addresses the problem of concept learning on off-policy or outdated data.
+If concepts are learned from data collected by a random policy, the concept distribution may not reflect the data from an optimal policy.
+To ensure that concept learning occurs on more recent and on-policy data, LICORICE interleaves concept learning and RL training through iterative training: it alternately freezes the network layers corresponding to either the concept learning part or the decision-making part.
+Second, LICORICE addresses the problem of limited training data diversity that occurs when an agent interacts with the environment, thereby generating sequences of highly similar, temporally correlated data points.
+To tackle this issue, LICORICE implements a data decorrelation strategy to produce a more diverse set of training samples.
+Third, LICORICE addresses the inefficient use of annotation effort, where labeled samples may provide redundant information.
+To resolve this problem, we employ disagreement-based active learning using a concept ensemble to select the most informative data points for labeling.
+To evaluate the effectiveness of LICORICE, we conduct experiments in two scenarios—perfect human annotation and VLM annotation—on five environments with image input: an image-based version of CartPole, two Minigrid environments, and two Atari environments.
+First, with assumption of perfect human annotation, we show that LICORICE yields both higher concept accuracy and higher reward while requiring fewer annotation queries compared to baseline methods.
+Second, we find that VLMs can indeed serve as concept annotators for some, but not all, of the environments.
+Our contributions are summarized as follows:
+  • To the best of our knowledge, we are the first to investigate the problem of a limited concept annotation budget for interpretable RL.
+We introduce LICORICE, a novel training scheme that enables label efficient learning of concept-based RL policies.
+  • We conduct extensive experiments to show the effectiveness of LICORICE across five environments with varying budget constraints.
+  • We study the use of VLMs as automated concept annotators, demonstrating their effectiveness in certain environments while highlighting challenges in others.
 2 Preliminaries
-2.1 Reinforcement Learning
-In RL, an agent learns to make decisions by interacting with an environment [27].
-The environment is commonly modeled as a Markov decision process [21], consisting of the following components: a set of states $\mathcal{S}$, a set of actions $\mathcal{A}$, a state transition function $T:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ that indicates the probability of transitioning from one state to another given an action, a reward function $R:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ that assigns a reward for each state-action-state transition, and a discount factor $\gamma\in[0,1]$ that determines the present value of future rewards.
-The agent learns a policy $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$, which maps states to actions with the aim of maximizing the expected cumulative discounted reward.
-We evaluate a policy via its value function, which is defined as $V^{\pi}(s)=\mathbb{E}_{\pi}[\sum_{k=0}^{\infty}\gamma^{k}r_{t+k+1}\mid s_{0}=s],\forall s\in\mathcal{S}$.
-The ultimate aim in RL is to determine the optimal policy, $\pi^{*}$.
-To do so, the agent iteratively refines its policy based on feedback from the environment.
-2.2 Concept Bottleneck Models
-Concept-based explanations have emerged as a common paradigm in explainable AI [20].
-They explain a model’s decision-making process through human-interpretable attributes and abstractions.
-Concept bottleneck models [13] are an example of concept-based explanations: they learn a mapping from samples $x\in X$ to labels $y\in Y$ through two functions, $g$ and $f$.
-The concept encoder function $g:X\rightarrow C$ maps samples from the input space to an intermediate space of human-interpretable concepts.
-The label predictor function $f:C\rightarrow Y$ maps samples from the concept space to a downstream task space, such as labels for supervised classification.
-As a result, the prediction $\hat{y}=f(g(x))$ relies on the input $x$ entirely through the bottleneck $\hat{c}=g(x)$.
-Training these models requires a dataset of $X\times C\times Y$, in which each sample consists of input features $x$, a ground truth concept vector $c\in\{0,1\}^{k}$, and a task label $y$.
-The functions $f$ and $g$ are parameterized by neural networks.
-These models are trained independently, jointly, or sequentially.
-The independent training method trains $\hat{f}$ and $\hat{g}$ independently.
-Critically, $\hat{f}$ is trained using the true $c$, but at test time it takes $\hat{g}(x)$ as input.
-If directly applied to the RL setting, this approach would require continuous access to ground-truth concepts for training $\hat{f}$, which we would like to avoid.
-In joint training, the model minimizes the weighted sum of the concept losses with the task performance loss.
-This paradigm can be problematic, as the losses can interfere with each other, and it may require careful tuning to find the right balance between the two losses.
-In sequential training, the model first learns $\hat{g}$.
-It then uses the concept predictions $\hat{g}(x)$ to learn $\hat{f}(\hat{g}(x))$.
-As we will later show, this setup is problematic for RL, as concepts may only emerge after the policy is sufficiently performant.
+Reinforcement Learning.
+In RL, an agent learns to make decisions by interacting with an environment (Sutton & Barto, 2018).
+The environment is commonly modeled as a Markov decision process (Puterman, 2014), consisting of the following components: a set of states $\mathcal{S}$, a set of actions $\mathcal{A}$, a state transition function $T:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow[0,1]$ that indicates the probability of transitioning from one state to another given an action, a reward function $R:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ that assigns a reward for each state-action-state transition, and a discount factor $\gamma\in[0,1]$ that determines the present value of future rewards.
+The agent learns a policy $\pi:\mathcal{S}\times\mathcal{A}\rightarrow[0,1]$, which maps states to distributions over actions with the aim of maximizing the expected cumulative discounted reward.
+We evaluate a policy via its value function, which is defined as $V^{\pi}(s)=\mathbb{E}_{\pi}[\sum_{k=0}^{\infty}\gamma^{k}r_{t+k+1}\mid s_{t}=s]$.
+The ultimate aim in RL is to determine the optimal policy, $\pi^{*}$, through iterative refinement based on environmental feedback.
+Concept Policy Models.
+Concept-based explanations have emerged as a common paradigm in explainable AI (Poeta et al., 2023).
+They explain a model’s decision-making process through human-understandable attributes and abstractions.
+In supervised learning, concept bottleneck models (Koh et al., 2020) implement this approach using two key functions: a concept predictor $g:X\rightarrow C$, mapping inputs to interpretable concepts, and an output predictor $f:C\rightarrow Y$, mapping the concept predictions to a downstream task space, such as labels for classification.
+As a result, the prediction takes the form $\hat{y}=f(g(x))$, so that the input $x$ influences the output solely through the bottleneck layer $\hat{c}=g(x)$.
+In RL, since we want the policy to be interpretable, we insert a concept bottleneck layer into the policy network:
+$\pi(s)=f(g(s)),$
+such that $\pi$ maps from states $s\in\mathcal{S}$ to concepts $c\in C$ to actions $a\in\mathcal{A}$ (Grupen et al., 2022; Zabounidis et al., 2023).
+This setup allows the policy to base its decisions on understandable and meaningful features.
+As a result, we can use any RL algorithm that can be modified to include an additional loss function for concept prediction.
+Training Concept Policy Models.
+Training these models requires a dataset of state-concept-action triplets $(s,c,a)\in\mathcal{S}\times C\times\mathcal{A}$.
+The functions $f$ and $g$ are typically implemented as neural networks, with their parameters collectively defined as $\theta$.
+Previous work (Zabounidis et al., 2023) simply combines the concept prediction loss $L^{C}(\theta)$ and RL loss $L^{\text{RL}}(\theta)$:
+$L(\theta)=L^{\text{RL}}(\theta)+\lambda_{C}L^{C}(\theta),$
+where the exact definitions of $L^{\text{RL}}(\theta)$ and $L^{C}(\theta)$ depend on the choice of RL algorithm and concept learning task.
+The objective is to find the optimal parameters $\theta^{*}$ that minimize this combined loss function with the coefficient $\lambda_{C}$.
+However, this approach requires continuous access to ground-truth concepts for training $f$, which may not always be feasible or desirable in practical RL scenarios.
 3 LICORICE
-We now describe LICORICE, our algorithm for interactively querying for concept labels under a limited concept label budget during RL training.
-In Section 3.1, we describe the architecture of the concept bottleneck policy.
-We then detail our iterative training process (Section 3.2) and explain our active learning approach for choosing informative training points (Section 3.3).
-Finally, in Section 3.4, we detail how we use VLMs as a substitute for human concept labeling.
-3.1 Concept Bottleneck Policies
-We insert a concept bottleneck layer into the policy network, such that $\pi$ maps from states $\mathcal{S}$ to concepts $C$ to actions $\mathcal{A}$, $\pi=f(g(s))$.
-These concepts serve as intermediaries in the policy network, which subsequently maps the concept vector to a distribution over possible actions.
-This setup allows the policy to base its decisions on understandable and meaningful features.
-We describe additional modifications necessary to accommodate the concept bottleneck in Section A.2.
-As a result, we can use any RL algorithm as long as we include an additional loss function for concept prediction.
-Loss Function
-We now describe how we use two loss functions for training in our iterative training scheme, described in more detail in Section 3.2.
-Because we employ an iterative training scheme, we disentangle the two loss functions — the concept prediction loss and the standard RL loss — to prevent interference between them.
-In this way, the concept prediction loss $L^{C}$ only affects $g$, the part of the model responsible for predicting concepts, and the standard RL loss $L^{\text{RL}}$ only affects $f$.
-For training $g$, we employ two types of loss functions depending on the nature of the concepts (MSE for continous concepts; cross-entropy loss for categorical concepts).
-If the problem requires mixed-type concepts, we could discretize the continuous attributes, converting them into categorical forms suitable for classification.
-This approach ensures that our method accommodates a diverse range of concept types.
-3.2 Training Process
-We now present LICORICE, our novel algorithm for optimizing concept-based models for RL within a fixed concept query budget.
-Our iterative training approach extends the sequential bottleneck training paradigm by incorporating multiple phases.
-Intuitively, this strategy can be advantageous for obtaining accurate on-policy concept estimates.
-To collect candidate training data for $g$, we employ our current policy for rollouts.
-To minimize data point correlation and collect a diverse range of data, we use a decorrelation strategy.
-To select more informative points to query for concept labels, we use a concept ensemble to calculate disagreement.
+As we have mentioned, the standard way of training concept-based RL assumes continuous access to an oracle to provide concept labels.
+However, this assumption is problematic due to the large annotation cost incurred by human or automated labeling efforts.
+To reduce the number of concept labels required for concept-based RL, we propose LICORICE, a novel algorithm for interpretable RL consisting of three main components: iterative training, data decorrelation, and disagreement-based active learning.
+The full pseudocode is in Algorithm 1.
+Iterative Training.
+As the agent’s policy improves, the distribution of the visited states and their associated concepts changes.
+Consider an MDP where states are indexed.
+With a random (initial) policy, the agent tends to visit small-index states near the initial state, encountering only the concepts relevant to those states.
+However, as the policy improves, the agent explores higher-index states, leading to a shift in both the state and concept visitation distribution.
+If we exhaust all queries at the beginning of training, we risk training the model only on the concepts associated with small-index states from the random policy, potentially missing important concepts that emerge as the agent explores.
+We therefore propose iterative training to enable LICORICE to progressively refine its understanding of concepts as the policy improves.
+Iterative training consists of two parts: concept learning and behavior learning.
+Assume that we have access to an annotated concept dataset $\mathcal{D}_{\text{train}}$ generated by our current policy $\pi_{m}$.
+Then, concept learning focuses on training the concept portion of the network $g$ on this dataset $\mathcal{D}_{\text{train}}$ (line 18).
+Behavior learning involves freezing $g$ and training the behavior portion of the network $f$ with any standard RL algorithm with its associated loss $L^{\text{RL}}$ (line 19).
+Upon completion of behavior learning, we obtain an updated policy $\pi_{m+1}$, which serves to collect more unlabeled concept data to help train $g$ in the next iteration.
 Algorithm 1 LICORICE (Label-efficient Interpretable COncept-based ReInforCEment learning)
@@
 16:     end while
-17:     Aggregate datasets: $\mathcal{D}_{\text{train}}\leftarrow\mathcal{D}_{\text{train}}\cup\mathcal{D}^{\prime}_{\text{train}}$, $\mathcal{D}_{\text{val}}\leftarrow\mathcal{D}_{\text{val}}\cup\mathcal{D}^{\prime}_{\text{validation}}$
+17:     Aggregate datasets: $\mathcal{D}_{\text{train}}\leftarrow\mathcal{D}_{\text{train}}\cup\mathcal{D}^{\prime}_{\text{train}}$, $\mathcal{D}_{\text{val}}\leftarrow\mathcal{D}_{\text{val}}\cup\mathcal{D}^{\prime}_{\text{val}}$
 18:     Continue training the concept network $g$ on $\mathcal{D}_{\text{train}}$, using $\mathcal{D}_{\text{val}}$ for early stopping
-19:     Freeze $g$ and train $f$ using standard RL training to obtain $\pi_{m+1}$
+19:     Freeze $g$ and continue training $f$ using standard RL training to obtain $\pi_{m+1}$
 20:end for
-More specifically, LICORICE (Algorithm 1) proceeds in three main stages within each iteration $m\in M$: data collection, concept ensemble training for data selection, and concept bottleneck policy training.
-First, in the data collection stage (lines 4 to 8), we collect a dataset of unlabeled concept data $\mathcal{U}_{m}$ by rolling out our current policy $\pi_{m}$.
-Crucially, during this step, we decorrelate the data to ensure the data points are diverse (line 6).
-Specifically, we set an acceptance probability $p$ for adding the data point to the dataset, as they are generated by the temporally-correlated RL policy.
-This prepares a diverse dataset for the next stage, with the aim of promoting the training of a more performant concept ensemble for data selection and, eventually, a more accurate concept bottleneck portion of the policy.
-The second stage is concept ensemble training for data selection (lines 9 to 16).
-In this stage, we train the concept ensemble — consisting of $N$ independent concept models — from scratch on the dataset of training points $\mathcal{D}_{\text{train}}$ that has been aggregated over all iterations (line 10).
-We use this ensemble to calculate the disagreement-based acquisition function, which evaluates whether each candidate in our unlabeled dataset $\mathcal{U}_{m}$ ought to receive a ground-truth concept label (line 11).
-This function targets samples where prediction disagreement is highest, detailed in Section 3.3.
-We then query for $B_{m}$ ground-truth concept labels (line 13) to prepare us for the next stage.
-We are now prepared for the third stage: concept bottleneck policy training (lines 17 to 19).
-We aggregate the concept datasets from the second stage with data from previous iterations, and continue training the concept network $g$ on $\mathcal{D}_{\text{train}}$ (line 18).
-After early stopping dictates that we stop $g$ training, we freeze $g$ and train $f$ using any standard RL algorithm — for example, PPO — to obtain the policy for collecting data in the next iteration (line 19).
-With that, we are prepared to start the process again from the first stage.
-3.3 Active Concept Learning
-We propose to leverage an ensemble of concept models to calculate the disagreement-based acquisition function for actively querying for concept labels.
-To quantify the concept disagreement of a state $U(s)$, we use two different formulations, depending on the concept learning task.
-Classification
-For concept classification, we use a query-by-committee [24] approach.
-When the models produce diverse predictions, it indicates that the instance is difficult and more information would be particularly valuable.
-Conversely, if all models agree, the instance is likely already well understood, and additional labels would be less beneficial.
-More specifically, we prioritize points with a high proportion of predicted class labels that are not the modal class prediction (also called the variation ratio [3]).
-This is given by $U(s)=1-\max_{c\in C}\frac{1}{N}\sum_{i=1}^{N}[\tilde{g}_{i}(s)=c]$, where $\tilde{g}_{i}(s)$ is the concept prediction of the $i$-th model on state $s$.
-Regression
-For concept regression, we observe that we can instead directly use variance as a measure of disagreement.
-Specifically, the concept disagreement of a state is quantified by the unbiased estimation of variance $U(s)=\sigma^{2}(s)$ of the predictions across the concept models, due to Bessel’s correction, defined as: $U(s)=\sigma^{2}(s)=\frac{1}{N-1}\sum_{i=1}^{N}(\tilde{g}_{i}(s)-\mu(s))^{2}$, where $\mu(s)=\frac{1}{N}\sum_{i=1}^{N}\tilde{g}_{i}(s)$ is the mean prediction of the $N$ concept models.
-3.4 Using Vision-Language Models for Concept Labeling
-Equipped with our overall algorithm, we now seek to further reduce human annotation burden.
-To do so, we turn to VLMs due to their remarkable performance in understanding and generating human-like descriptions of visual content [22].
-Within the pipeline of LICORICE, we keep all aspects of our algorithm the same but use the VLM as the concept annotator in the training loop.
-By doing so, we effectively decrease the human annotation effort to zero, albeit with some labeling inaccuracy introduced.
-In our experiments, we use GPT-4o [1], a closed-source model that is possibly the most capable vision-language model in the world at the time of writing the paper.
-During the training loop of LICORICE, we query GPT-4o each time the algorithm requires a concept label (line 13 in Algorithm 1).
-As we are using a pre-specified concept set, we design prompts with detailed general descriptions of the scene layout and definitions of all concepts, in a similar fashion to giving labeling instructions to real humans.
-We then prompt the GPT-4o to obtain the generated labels.
-In environments where continuous concept values are required, we ask GPT-4o to give as accurate estimates as possible; in environments with concepts that are discrete and more intuitive to label, we provide clear instructions of how to read the concept numbers according to the input image.
-More details about our prompts can be found in Section A.3.
+Data Decorrelation.
+We do not obtain ground-truth concept labels for all states when rolling out the current policy $\pi_{m}$.
+Instead, rollouts produce a dataset $\mathcal{U}_{m}$ of unlabeled states as candidates for querying for ground-truth concepts from an oracle.
+Simply collecting and labeling all encountered states would give us long chains of nearly-identical samples, undermining the diversity we need for effective learning.
+To resolve this challenge, we introduce data decorrelation with two key components: a budget multiplier $\tau$ that sets $|\mathcal{U}_{m}|=\tau\cdot B_{m}$, and a per-state acceptance probability $p$.
+This random acceptance/rejection mechanism leverages a fundamental property of Markov processes—states become increasingly independent as their temporal distance grows according to the mixing time—allowing us to build more diverse datasets from our policy rollouts.
+Disagreement-Based Active Learning.
+Equipped with $\mathcal{U}_{m}$, we are now prepared to select data points for querying for concept labels.
+The purpose of this stage is to make use of our limited labeling budget $B_{m}$ to collect a labeled dataset $\mathcal{D}_{m}$ for training $g$.
+We propose to train a concept ensemble—consisting of $N$ independent concept models—from scratch on the dataset of training points $\mathcal{D}_{\text{train}}$ that has been aggregated over all iterations (line 10).
+We use this ensemble to calculate the disagreement-based acquisition function $\alpha(\cdot)$, which determines whether each candidate in $\mathcal{U}_{m}$ ought to receive a ground-truth concept label (line 11).
+This function targets samples where prediction disagreement is highest among ensemble members, as these points often represent areas of uncertainty or decision boundaries where additional labeled data would be most informative.
+For $\alpha(\cdot)$, we use two different formulations, depending on the concept learning task.
+For concept classification, we use a query-by-committee approach (Seung et al., 1992), in which we prioritize points with a high proportion of predicted class labels that are not the modal class prediction (also called the variation ratio (Beluch et al., 2018)):
+$\alpha(s)=1-\max_{c\in C}\frac{1}{N}\sum_{i=1}^{N}[\tilde{g}_{i}(s)=c].$
+Here, $\tilde{g}_{i}(s)$ is the concept prediction of the $i$-th model on state $s$.
+For concept regression, we directly use the estimate of variance across the concept models as a measure of disagreement:
+$\alpha(s)=\sigma^{2}(s)=\frac{1}{N-1}\sum_{i=1}^{N}(\tilde{g}_{i}(s)-\mu(s))^{2},$
+where $\mu(s)=\frac{1}{N}\sum_{i=1}^{N}\tilde{g}_{i}(s)$ is the mean prediction of the $N$ concept models.
+After querying for $B_{m}$ ground-truth concept labels (line 13) using $\alpha(\cdot)$, we are prepared to continue training $g$ during the concept learning stage.
+Implementation Details.
+For training $g$ (line 18), we employ two types of loss functions depending on the nature of the concepts: MSE for continuous concepts and cross-entropy loss for categorical concepts.
+If the problem requires mixed-type concepts, we either discretize continuous attributes or simply use a mixed loss.
+To train $f$ (line 19), we freeze $g$ and use PPO to continue training $f$ from the previous iterations, resulting in the final policy $\pi_{M+1}$.
+For complex environments that require many iterations, we observe that RL-related parameters may get stuck in a local region in early iterations, becoming hard to optimize later.
+To mitigate this issue, in the last iteration, we additionally train a new RL network $f^{\prime}$ from scratch when fixing $g$ with $\pi_{M+1}$ as the anchoring policy to get the final $\pi^{\prime}_{M+1}$.
+The anchoring policy ensures $\pi^{\prime}_{M+1}$ has a similar distribution to the previous one and the annotated observations are still highly relevant.
+Specifically, we initialize another policy $\pi_{\theta}$ with the same frozen $g$ parameters and randomly initialized $f$, then train it with the standard PPO loss function and an additional KL-divergence penalty between the current policy and $\pi_{M+1}$:
+$L(\theta)=L^{\text{PPO}}(\theta)+\beta\cdot\text{KL}[\pi_{M+1}(\cdot\mid s),\pi_{\theta}(\cdot\mid s)].$
 4 Experiments
@@
   1. RQ 1
-    Does LICORICE enable both high concept accuracy and high environment reward?
+    Under a limited concept annotation budget, does LICORICE enable both high concept accuracy and high environment reward?
   2. RQ 2
-    To what extent does LICORICE enable a reduction in human labeling effort without substantially impacting performance?
+    How effective are VLMs as automated concept annotators when used with LICORICE?
   3. RQ 3
-    Can LICORICE be used alongside powerful vision-language models to further alleviate the labeling burden?
-4.1 Environments
-We evaluate our approach on three environments. For each environment, we define an interpretable concept set, where we define concepts to describe properties of objects of interest.
-PixelCartPole
-In PixelCartPole-v1 [31], the states are images, and the concepts are the original continuous features in the standard CartPole environment: the cart position, the cart velocity, the pole angle, and the pole angular velocity.
-To obtain the mapping from images to concepts, we use a fixed window of the four most recent images for the temporal concepts, such as the cart velocity.
-We also include the last action as input to ensure that the information is sufficient to infer concept values.
-This domain is deceptively difficult due to the temporal concepts.
-DoorKey
-In DoorKey-7x7 [5], the agent operates in a 5x5 grid [^1] In both of the Minigrid environments, the actual usable area is smaller than what shows in its name, as the outermost layer is a boundary.
-to pick up an item to unlock the door, then reach the green goal square.
-The states are fully observable images.
-We define 6 groups of concepts with 12 concept variables in total: 1) the agent’s position, 2) the agent’s direction, 3) the key’s position or a fixed position (0, 0) outside the grid if the key has been picked up, 4) the door’s position, 5) whether the door is open, and 6) whether the agent can move along each of four directions, i.e., the corresponding cell is empty.
-Each group contains one or more concept variables.
-For example, position groups contain x and y coordinates.
-DynamicObstacles
-In DynamicObstacles-5x5 [5], the agent operates in a 3x3 grid to avoid two moving obstacles while navigating to the goal.
-Colliding with the dynamic obstacles yields a large penalty, so the agent must correctly learn concepts to safely reach the goal.
-The states are fully observable images.
-We define 5 groups of concepts with 11 concept variables in total: 1) the agent’s position, 2) the agent’s direction, 3) the first obstacle’s position, 4) the second obstacle’s position, and 5) whether the agent can move along each of four directions, i.e., the corresponding cell is empty.
-Each group contains one or more concept variables.
-For example, position groups contain x and y coordinates.
-4.2 Experiment Details
-In each experiment, we run each algorithm 5 times, each with a random seed. More implementation details and hyperparameters are in Section A.2.
-Architecture
-For the backbone model and algorithm, we use PPO (implementation provided by Stable Baselines 3 [23]).
-We preserve most of the network architecture, but we add a concept layer in the policy network.
-Specifically, after passing the image through the feature extractor, which consists of 3 CNN layers, we introduce a linear layer to map to the concept layer and obtain predicted concept values.
-Each predicted concept is represented as a real value for the regression case and one categorical value derived from a classification head for the classification case.
-This portion of the network corresponds to $g$.
-For $f$, we utilize an MLP extractor with 2 fully connected layers, each with 64 neurons and a Tanh activation function.
-We share the default CNN feature extractor between policy and value functions, as we found this to be beneficial in preliminary experiments.
-Concept Representation
-We model concept learning for PixelCartPole-v1 as a regression problem (minimizing mean squared error) as the concepts are real-valued, and concept learning for DoorKey-7x7 and DynamicObstacles-5x5 as classification problems as the concepts are categorical.
-The observation images have resolutions of $(240,160)$ for PixelCartPole-v1, and $(160,160)$ for DoorKey-7x7 and DynamicObstacles-5x5.
-These dimensions maintain a short side of 160 pixels while preserving the aspect ratio of the original rendered images, as we observe lower resolutions may impact performance.
-Performance Metrics
-To obtain an upper bound on the reward for each environment, we directly use ground-truth concept labels to learn a policy, leading to a reward of $500$, $0.97$, and $0.91$ for the three environments respectively.
-In the following sections, we report the reward as a ratio of this upper bound.
-Percentages (or ratios) make sense here since the minimum reasonable reward is 0 in all environments: in PixelCartPole-v1 and DoorKey-7x7, all rewards are nonnegative; in DynamicObstacles-5x5, a random policy would have negative reward due to collisions, but the agent can ensure nonnegative reward by simply staying in place.
-We additionally report the concept error (MSE for regression; $1$ accuracy for classification).
-All the reported numbers are calculated during the testing stage, where we evaluate the models on 100 episodes.
+    How label efficient is LICORICE compared with other methods?
+  4. RQ 4
+    Does LICORICE support test-time concept interventions?
+4.1 Experiment Setup
+In each experiment, we run each algorithm 5 times, each with a random seed.
+All algorithms use PPO (Schulman et al., 2017; Raffin et al., 2021) with a concept bottleneck.
+More implementation details and hyperparameters are in Section A.2.
+Environments.
+We investigate these questions across five environments: PixelCartPole (Yang et al., 2021), DoorKey (Chevalier-Boisvert et al., 2023), DynamicObstacles (Chevalier-Boisvert et al., 2023), Boxing (Bellemare et al., 2013), and Pong (Bellemare et al., 2013).
+Each environment includes a distinct challenge and features a set of interpretable concepts describing key objects and properties.
+We summarize the concepts in Table 1, with more details in Section A.1.
+These environments are characterized by their dual representation: a complex image-based input and a symbolic concept-based representation.
+PixelCartPole, DoorKey, and DynamicObstacles are simpler because we can extract noiseless ground-truth concept labels from their source code.
+In contrast, Boxing, as implemented in OCAtari (Delfosse et al., 2023), uses reverse engineering to extract positions of important objects from the game’s RAM state.
+This extraction process introduces a small amount of noise.
+The concepts in Pong are derived from the VIPER paper (Bastani et al., 2018) that also uses reverse engineering.
+Environment # Concepts Concept Type Binarized # Concepts
+PixelCartPole 4 Continuous N/A
+DoorKey 12 Discrete 46
+DynamicObstacles 11 Discrete 30
+Boxing 8 Discrete 1480
+Pong 7 Discrete 1848
+Table 1: Summary of environments, including their associated concepts. Counts for the binarized version of the concepts provided where applicable to illustrate the problem size.
+Baselines.
+To our knowledge, there exist no previous algorithms that seek to minimize the number of concept labels for interpretable RL, so we implement three: Sequential-Q, Disagreement-Q, and Random-Q.
+In Sequential-Q, the agent spends $B$ queries on the first $B$ states encountered during the initial policy rollout.
+In Disagreement-Q, the agent also spends its budget at the beginning of training; however, it uses active learning with the same $\alpha(\cdot)$ as LICORICE to strategically choose the training data for annotation.
+In Random-Q, the agent receives $B$ concept labels at random points using a probability to decide whether to query in each state.
+As a budget-unconstrained baseline, we implement CPM from previous work in multi-agent RL (Zabounidis et al., 2023) for the single-agent setting.
+As mentioned in §2, CPM jointly trains $g$ and $f$, assuming unlimited access to concept labels, so it also represents an upper bound on concept accuracy.
+VLM Details.
+For our VLM experiments, we use GPT-4o (gpt, ).
+During training, we query GPT-4o each time LICORICE requires a concept label.
+As an example, for the concept related to the position of an object, we prompt GPT-4o with instructions to report the coordinates (x y).
+More details about our prompts can be found in Section A.3.
+Performance Metrics.
+We compare our reward to an upper bound calculated by training PPO with ground-truth concept labels (PixelCartPole: $500$, DoorKey: $0.97$, DynamicObstacles: $0.91$, Boxing: $86.05$, Pong: $21$).
+In the first four environments, percentages (or ratios) make sense since the minimum reasonable reward is 0.222In PixelCartPole and DoorKey, all rewards are nonnegative; in DynamicObstacles, an agent can ensure nonnegative reward by simply staying in place; in Boxing, a random policy can get around 0 reward and all of the trained policies have positive reward.
+In Pong, a random policy has -21 reward so we first get the difference between rewards and -21 and then calculate the ratio.
+We additionally report the concept error (MSE for regression; $1\,-\,$accuracy for classification).
+In Boxing, following the practice of OCAtari, we consider a concept prediction correct if it is within 5 pixels of the ground truth label.
+In Pong, we simply use classification error.
+All reported numbers are calculated using 100 evaluation episodes.
 4.2 Results
-PixelCartPole-v1 DoorKey-7x7 DynamicObstacles-5x5
-$c$ Labels Algorithm $R$uparrow $c$ MSE $\downarrow$ $R$uparrow $c$ Error $\downarrow$ $R$uparrow $c$ Error $\downarrow$
-$B$ Sequential-Q 0.232 4720.2324720.232\,4720.232 472 0.102 570.102570.102\,570.102 57 0.462 5660.4625660.462\,5660.462 566 0.462 3270.4623270.462\,3270.462 327 0.788 3480.7883480.788\,3480.788 348 0.014 1360.0141360.014\,1360.014 136
-Disagreement-Q 0.363 6240.3636240.363\,6240.363 624 0.097 4570.0974570.097\,4570.097 457 0.759 4640.7594640.759\,4640.759 464 0.392 3940.3923940.392\,3940.392 394 0.992 530.992530.992\,530.992 53 0.000 7790.0007790.000\,7790.000 779
-Random-Q 0.337 720.337720.337\,720.337 72 0.065 6130.0656130.065\,6130.065 613 0.886 5670.8865670.886\,5670.886 567 0.269 730.269730.269\,730.269 73 0.981 0960.9810960.981\,0960.981 096 0.022 4280.0224280.022\,4280.022 428
-LICORICE $1.0$ 0.024 7970.0247970.024\,7970.024 797 0.990 2970.9902970.990\,2970.990 297 0.047 5040.0475040.047\,5040.047 504 0.990 6070.9906070.990\,6070.990 607 0.000 1930.0001930.000\,1930.000 193
-$\infty$ CPM  [33] $1.0$ 0.011 5450.0115450.011\,5450.011 545 1.001 0241.0010241.001\,0241.001 024 $0.0$ 0.969 5770.9695770.969\,5770.969 577 $0.0$
-Table 1: Evaluation of the reward $R$ and concept error achieved by all methods in all environments.
-The reward is reported as the fraction of the reward upper bound.
-Unless otherwise stated, concept error is 1 - accuracy.
-Methods in the top section are given a budget of $B=[500,300,300]$, respectively; CPM is given an unlimited budget (in practice, it uses 4M, 4M, 1M concept labels respectively).
-For each column, the top-performing method with a limited budget is in bold.
-If there are several methods with overlapping standard deviations, we bold all of them.
-Full results with standard deviation are in Table 6, Appendix B.
-4.3.1 Balancing Concept Performance and Environment Reward
-We first validate that LICORICE can achieve high reward in all test environments while accurately identifying concepts.
-We first compare against baselines with a fixed budget of $B=[500,300,300]$ queries for PixelCartPole-v1, DoorKey-7x7, and DynamicObstacles-5x5, respectively, as depicted in the first section of Table 1.
-We choose these budgets by starting from $500$ and then decreasing by units of $100$ until we find that LICORICE can no longer achieve $99\%$ of the reward upper bound if we further decrease the budget.
-We treat the minimum budget under which LICORICE can achieve $99\%$ of the reward upper bound as the maximum budget considered for that environment.
-Comparison with Budget-Constrained Baselines
-To our knowledge, no previous algorithms exist that seek to minimize the number of train-time labels for concept-based RL training, so we implement our own baselines.
-In Sequential-Q, the agent spends all of $B$ queries on the first $B$ states it encounters during the initial policy rollout.
-In Disagreement-Q, the agent similarly spends its budget at the beginning of its learning process; however, it uses active learning with the same acquisition function as LICORICE to strategically choose the training data.
-In Random-Q, the agent receives $B$ concept labels at random points in the training process using a probability to decide whether to query for a concept for each state.
-We show the results in the first section of Table 1.
-In the three environments, LICORICE performs similarly to or better than all baselines in terms of both reward and concept error.
-The main performance differences can be seen for PixelCartPole-v1 and DoorKey-7x7, where LICORICE achieves $100\%$ and $99\%$ of the optimal reward, respectively, while the second best algorithm achieves $36\%$ and $89\%$, respectively.
-Random-Q, Disagreement-Q, and LICORICE perform similarly on DynamicObstacles-5x5, indicating that this environment is relatively simple and may not benefit from a more advanced strategy.
-Its simplicity may be due to the lack of concept distribution shift that is present in the other two environments.
-For example, in PixelCartPole-v1, the policy must be further refined to improve its estimate of the more on-policy concepts.
-In the initial training stages, when the policy is more random, the agent needs to estimate a wide range of possible values for the pole angle, as the pole starts upright and frequently falls.
-As the agent’s policy improves, it requires more precise estimates of the pole angle primarily around the vertical position.
-Comparison with Budget-Unconstrained Baseline
-To better understand the performance of our approach, we compare it against existing work that is not constrained by concept label budgets.
-Specifically, we implement Concept Policy Model (CPM) from previous work in multi-agent RL [33] but for the single-agent setting.
-This approach jointly trains the concept bottleneck and the policy, assuming unlimited access to concept labels.
-It represents an upper bound on concept accuracy, as the agent receives continuous concept feedback throughout learning.
-Surprisingly, LICORICE outperforms CPM in PixelCartPole-v1and has similar performance to CPM in the other two environments in this unfair comparison.
-We emphasize that CPM is given an unlimited budget, and in fact, it uses over 1M concept labels for each environment, whereas LICORICE uses $500$ or fewer.
-We therefore answer RQ 1 affirmatively: compared to baselines, LICORICE demonstrates both high concept accuracy (low concept error) and high reward on our three test environments, all while using substantially fewer concept labels than existing algorithms.
-PixelCartPole-v1 DoorKey-7x7 DynamicObstacles-5x5
-$B$ 300 400 500 100 200 300 100 200 300
-$R$ 0.276 3440.2763440.276\,3440.276 344 0.766 2880.7662880.766\,2880.766 288 $1.0$ 0.765 3910.7653910.765\,3910.765 391 0.934 5320.9345320.934\,5320.934 532 0.990 2970.9902970.990\,2970.990 297 0.957 4390.9574390.957\,4390.957 439 0.979 8190.9798190.979\,8190.979 819 0.990 6070.9906070.990\,6070.990 607
-$c$ Error 0.153 5730.1535730.153\,5730.153 573 0.052 4580.0524580.052\,4580.052 458 0.024 7970.0247970.024\,7970.024 797 0.260 4040.2604040.260\,4040.260 404 0.086 6350.0866350.086\,6350.086 635 0.047 5040.0475040.047\,5040.047 504 0.040 7940.0407940.040\,7940.040 794 0.007 8340.0078340.007\,8340.007 834 0.000 1930.0001930.000\,1930.000 193
-Table 2: Performance of LICORICE on all environments for varying maximum budgets.
-The reward is reported as the fraction of the reward upper bound.
-For PixelCartPole-v1, $c$ error is MSE; for the other environments, $c$ error is 1 - accuracy.
-For all environments, the reward increases and the error decreases as we increase the maximum budget.
-Full results with standard deviation are in Table 7Appendix B.
-4.3.2 Budget Allocation Effectiveness
-To understand the extent to which LICORICE enables a reduction in human labeling effort without substantially impacting performance, we conduct experiments across three different environments with varying concept labeling budgets.
-We start with the baseline budget $B$ from our previous experiments and test two additional budgets, decreasing by steps of 100.
-The only component of our algorithm that we vary is the number of iterations $M$.
-For each budget and environment, we report a single performance value that maximizes the sum of the relative reward and inverse concept error, both scaled to $[0,1]$.
-The results of this experiment are shown in Table 2.
-Our findings reveal that the impact of budget reductions (human labeling effort) varies across environments.
-In general, we see an increase in concept error and a decrease in reward for all environments with decreasing budget.
-However, both DynamicObstacles-5x5 and DoorKey-7x7 show more resilience to budget reductions, while PixelCartPole-v1 exhibits higher sensitivity.
-Specifically, we find that even with $B=100$ for DynamicObstacles-5x5, LICORICE can still achieve $96\%$ of the optimal reward.
-With $B=200$, LICORICE can still achieve $93\%$ of the optimal reward on DoorKey-7x7.
-In contrast, we see large decreases in reward from the budget reductions for PixelCartPole-v1.
-We therefore provide a more nuanced answer to RQ 2: LICORICE can indeed enable reductions in human labeling effort without substantially impacting performance, but the extent of this reduction is environment-dependent.
-In simpler environments, human effort can be dramatically reduced while maintaining near-optimal performance.
-However, in more complex environments like PixelCartPole-v1, the trade-off between reduced human effort and maintained performance becomes more pronounced.
-This finding aligns with the intuition that different environments have varying levels of complexity and would require different levels of human input to accurately capture all relevant concepts and their relationships.
-Figure 1: Training curves of reward and concept error for PixelCartPole-v1. Shaded values correspond to 95% CI, calculated using 1000 bootstrap samples. All curves in Figure 5.
-4.3.3 Integration with Vision-Language Models
+Figure 2: LICORICE generally achieves higher reward and lower concept error than all other budget-constrained algorithms (Sequential-Q, Disagreement-Q, Random-Q).
+The budgets for each environment are PixelCartpole: $500$, DoorKey: $300$, DynamicObstacles: $300$, Boxing: $3000$, Pong: $5000$.
+CPM has an unlimited budget (in practice, it uses 4M, 4M, 1M, 15M, 10M labels, respectively).
+Despite this extreme budget difference, LICORICE achieves comparable reward and concept error.
+Full results in Table 5, Appendix B.
+RQ 1: Reward and concept accuracy under a limited annotation budget
+LICORICE achieves similarly high reward and concept performance to the state-of-the-art budget-unconstrained baseline.
+We first seek to understand how LICORICE performs compared to the state-of-the-art in concept-based RL, CPM, which is not constrained by concept label budgets.
+Surprisingly, LICORICE and CPM achieve nearly 100% of the maximum reward in all environments in this unfair comparison (Figure 2).
+They also achieve a similar concept error rate, only seeing a clear difference in error for the most complex environment, Pong.
+We emphasize that CPM is given an unlimited budget; in practice, it uses 1M-15M concept labels for all environments, which is around $2000$ to 13000×13000$\times$13000 × the budget used by LICORICE.
+In contrast, LICORICE operates under a strict budget constraint, with $3000$ labels for Boxing and $5000$ for Pong, and at most $500$ concept labels for the simpler environments (PixelCartPole: $500$, DoorKey: $300$, DynamicObstacles: $300$).
+LICORICE generally achieves higher reward and lower concept error than budget-constrained baselines.
+We study all budget-constrained algorithms (LICORICE plus those in §4.1) under the same fixed concept labeling budget as above.
+Figure 2 shows that LICORICE outperforms all baselines in terms of reward and concept error on all but arguably the easiest environment, DynamicObstacles.
+In that environment, LICORICE performs similarly to the baselines.
+The greatest performance differences occur in PixelCartPole and DoorKey, where LICORICE achieves $100\%$ and $99\%$ of the maximum reward, while the second-best algorithm achieves $36\%$ and $89\%$, respectively.
+We therefore answer RQ 1 affirmatively: not only does LICORICE achieve the same high concept accuracy and high reward as the state-of-the-art in concept-based RL, it also performs on-par to budget-constrained baselines in one environment and outperforms them in the other four.
+RQ 2: VLMs as concept annotators
 We now seek to answer the question of whether VLMs can successfully provide concept labels in lieu of a human annotator within our LICORICE framework.
-For consistency, we use the same hyperparameters for LICORICE as in Section 4.3.2.
-Because using VLMs incurs costs and users requiring interpretable policies for their environments may still face budget constraints, we operate within the same budget-constrained setting as described in Section 4.3.2.
-Here, the reduction in the number of “human” labels required translates to cost savings.
-We show the results in Table 3.
-We find that GPT-4o can indeed serve as an annotator for some but not all environments.
-In DoorKey-7x7 and DynamicObstacles-5x5, LICORICE with GPT-4o labels achieves 83% and 88% of the optimal reward, respectively.
-The concept error rate is comparable to GPT-4o’s labeling error, both evaluated on the same rollout observations, with sampling applied to the latter due to cost constraints.
-Additionally, the concept error gap between our trained concept model and GPT-4o almost diminishes with more labels.
-Taken together, these results suggest that LICORICE can work reasonably well with GPT-4o labels.
-In PixelCartPole-v1, however, there are challenges for GPT-4o in providing accurate labels due to the continuous nature of the concepts and the lack of necessary knowledge of physical rules and quantities in this specific environment.
-As a result, LICORICE with GPT-4o labels achieves generally low reward across varying budgets.
-Interestingly, the concept error increases with more budget, likely due to the additional labeling noise introduced to the concept network.
-This suggests that while the VLM can handle labeling tasks to some extent, its performance in PixelCartPole-v1 is limited by the complexity of accurately labeling concepts.
-We also posit the mean reward drop in DynamicObstacles-5x5 at 300 budget is due to the randomness introduced by GPT-4o; however, we note that the range of standard deviation overlaps with the other two budgets.
-We therefore answer RQ 3 with cautious optimism: there are indeed cases in which LICORICE could be used alongside VLMs.
-Generally speaking, a more human-intuitive and less complex environment is more likely to work well with VLMs by enabling clear and detailed labeling instructions in the prompt.
-PixelCartPole-v1 DoorKey-7x7 DynamicObstacles-5x5
-$B$ 300 400 500 100 200 300 100 200 300
-$R$ (LICORICE GPT-4o) 0.064 7120.0647120.064\,7120.064 712 0.066 2160.0662160.066\,2160.066 216 0.064 940.064940.064\,940.064 94 0.711 4030.7114030.711\,4030.711 403 0.735 0890.7350890.735\,0890.735 089 0.830 1130.8301130.830\,1130.830 113 0.268 7110.2687110.268\,7110.268 711 0.926 0950.9260950.926\,0950.926 095 0.875 0150.8750150.875\,0150.875 015
-$c$ Error (LICORICE GPT-4o) 0.170 0680.1700680.170\,0680.170 068 0.246 0870.2460870.246\,0870.246 087 0.246 895 7330.2468957330.246\,895\,7330.246 895 733 0.450 3260.4503260.450\,3260.450 326 0.329 3480.3293480.329\,3480.329 348 0.312 2880.3122880.312\,2880.312 288 0.200 7110.2007110.200\,7110.200 711 0.133 8850.1338850.133\,8850.133 885 0.132 220.132220.132\,220.132 22
-$c$ Error (GPT-4o) 0.193 750.193750.193\,750.193 75 0.224 6830.2246830.224\,6830.224 683 0.241 3650.2413650.241\,3650.241 365 $0.313$ 0.313 6670.3136670.313\,6670.313 667 $0.304$ 0.159 6360.1596360.159\,6360.159 636 0.123 6360.1236360.123\,6360.123 636 0.127 6360.1276360.127\,6360.127 636
-Table 3: Performance of LICORICE with GPT-4o integrated into the loop for all environments across different budgets.
-We also report the concept labeling error on a random sample of 50 observations from the same rollout set used for LICORICE+GPT-4o, due to API cost constraints.
-Full results with standard deviation are in Table 8, Appendix B.
-PixelCartPole-v1 DoorKey-7x7 DynamicObstacles-5x5
-Algorithm $R$uparrow $c$ MSE $\downarrow$ $R$uparrow $c$ Error $\downarrow$ $R$uparrow $c$ Error $\downarrow$
-LICORICE-IT 0.533 8960.5338960.533\,8960.533 896 0.079 6220.0796220.079\,6220.079 622 0.992 4120.9924120.992\,4120.992 412 0.089 5350.0895350.089\,5350.089 535 0.996 8890.9968890.996\,8890.996 889 0.001 2020.0012020.001\,2020.001 202
-LICORICE-DE $0.9669$ 0.031 1060.0311060.031\,1060.031 106 0.783 1130.7831130.783\,1130.783 113 0.200 0220.2000220.200\,0220.200 022 0.980 5370.9805370.980\,5370.980 537 0.000 1140.0001140.000\,1140.000 114
-LICORICE-AC 0.910 7840.9107840.910\,7840.910 784 0.022 3140.0223140.022\,3140.022 314 0.815 7550.8157550.815\,7550.815 755 0.161 8420.1618420.161\,8420.161 842 0.583 0530.5830530.583\,0530.583 053 0.002 2910.0022910.002\,2910.002 291
-LICORICE $1.0$ 0.024 7970.0247970.024\,7970.024 797 0.990 2970.9902970.990\,2970.990 297 0.047 5040.0475040.047\,5040.047 504 0.990 6070.9906070.990\,6070.990 607 0.000 1930.0001930.000\,1930.000 193
-Table 4: Ablation study results for LICORICE in all environments.
-All of our components generally help achieve a better reward and lower concept error.
-Full results with standard deviation are in Table 9, Appendix B.
-4.3.4 Ablation Study
-Given the previous results, we now conduct ablations to confirm the effectiveness of our three main contributions: iterative training, decorrelation, and active learning.
-LICORICE-IT corresponds to LICORICE with only one iteration, LICORICE-DE corresponds to LICORICE without decorrelation, and LICORICE-AC corresponds to LICORICE without active learning (instead, it uses the entire unlabeled dataset for querying).
-We show the learning curves for PixelCartPole-v1 in Figure 1; the full set of curves is in Appendix B.
-Table 4 depicts the results of our ablation study on all environments.
-The bottom row corresponds to the upper bound on performance by LICORICE, with all components included.
-All of our contributions are critical to achieving both high reward and low concept error.
-However, the component that most contributes to the reward or concept performance differs depending on the environment.
-For example, compared with LICORICE, LICORICE-IT exhibits the largest reward gap for PixelCartPole-v1; however, LICORICE-AC yields the largest reward gap for DynamicObstacles-5x5, and LICORICE-DE yields the largest gap for DoorKey-7x7.
-We suspect that this difference is because the concepts in DynamicObstacles-5x5 are simple enough such that one iteration is sufficient for learning, meaning that the largest gains can be made by using active learning.
-In contrast, PixelCartPole-v1 requires further refinement of the policy to better estimate on-distribution concept values, so the largest gains can be made by leveraging multiple iterations.
+Because using VLMs incurs costs, we operate within the same budget-constrained setting as above.
+We present these results in Figure 2.
+$c$ Error
+Environment LICORICE+GPT-4o GPT-4o
+PixelCartPole 0.246 8960.2468960.246\,8960.246 896 0.241 3650.2413650.241\,3650.241 365
+DoorKey 0.312 2880.3122880.312\,2880.312 288 0.304 $0$.3040000.304\,$0$.304 000
+DynamicObstacles 0.132 220.132220.132\,220.132 22 0.127 6360.1276360.127\,6360.127 636
+Boxing 0.983 520.983520.983\,520.983 52 0.815 8330.8158330.815\,8330.815 833
+Pong 0.728 870.728870.728\,870.728 87 0.874 130.874130.874\,130.874 13
+Table 2: Concept error comparison. The concept error of LICORICE+GPT-4o matches that of GPT-4o alone in all environments except Boxing.
+VLMs can serve as concept annotators for some environments.
+In DoorKey and DynamicObstacles, LICORICE with GPT-4o labels achieves 83% and 88% of the maximum reward, respectively.
+To assess the quality of VLM-generated labels, we compare the concept error rate of our trained model against GPT-4o’s labeling error, both evaluated on the same rollout observations, while GPT-4o evaluation is on a random subset of 50 observations, due to API cost constraints.
+Table 2 shows that the concept error of LICORICE is comparable to that of GPT-4o when GPT-4o labeling error is not high, which suggests that LICORICE can effectively utilize these concept labels.
+VLMs, particularly GPT-4o, could serve as concept annotators for certain concepts in certain environments.
+VLMs struggle to provide accurate concepts in more complex environments.
+In PixelCartPole, Boxing, and Pong environments, however, GPT-4o struggles to provide accurate labels.
+In these environments, not only does LICORICE+GPT-4o achieve less than $25\%$ of the maximum reward, it also incurs a large concept error.
+The challenge with PixelCartPole is the continuous nature of the concepts and the lack of necessary knowledge of physical rules and quantities in this specific environment.
+Boxing is particularly challenging due to the large number of possible concept values that each of the 8 discrete concepts can take on (160 or 210 possible values).
+Similarly, concepts in Pong have hundreds of possible values and some physics quantities require multi-frame inference.
+We therefore answer RQ 2 with cautious optimism: there are indeed cases in which LICORICE could be used alongside VLMs, though not all yet.
+Generally speaking, a less complex environment is more likely to work well with VLMs by enabling clear and detailed labeling instructions in the prompt.
+However, VLM capabilities need improvement before they can fully alleviate the human annotation burden, especially in safety-critical settings.
+RQ 3: Investigation of Budget Requirements for Performance
+Given these results, we now investigate the minimum budget required for all environments to achieve 99% of the reward upper bound.
+To do so, we incrementally increase the budget for each environment starting from a reasonably small budget until LICORICE reaches 99% of the reward upper bound.
+In Pong, we further increase the budget until the concept error is below $0.25$.
+In addition to the baselines, we study LICORICE under perfect (human) annotation and noisy VLM labels in Figure 3.
+LICORICE more rapidly achieves high reward compared to baselines.
+Across all environments, LICORICE is the most query-efficient.
+In three of the environments, it consistently achieves higher reward than baselines across all budget levels.
+In DynamicObstacles and Pong, LICORICE outperforms baselines at the smallest query budget, after which some baseline method achieves comparable reward.
+LICORICE is similarly performant with respect to concept error.
+Except for one budget setting for one environment, LICORICE consistently achieves lower concept error than the baselines.
+For LICORICE+GPT-4o, more concept labels is not always better.
+In DoorKey, LICORICE+GPT-4o exhibits predictable behavior in terms of concept error: as the budget increases, the reward increases and the concept error decreases.
+Counterintuitively, for some environments, the concept error and reward fluctuate with more budget, likely due to the additional labeling noise introduced.
+We therefore provide a more nuanced answer to RQ 3: LICORICE is indeed label efficient across varying budgets, but the benefit is annotator-dependent.
+Figure 3: Reward (top) and concept error (bottom) of all algorithms for different budgets.
+LICORICE more efficiently makes use of the varying budgets, achieving higher reward and lower concept error.
+Full results with standard deviation are in Tables 6, 7 and 8, Appendix B.
+RQ 4: Interpretability Analysis: Test-Time Concept Interventions
+A great property of concept-based networks is the ability for people to successfully intervene on the concepts to correct them.
+In RL, this intervention enables the inspection of how different concepts influence the immediate decisions of the agent.
+Figure 4: Concept intervention results: LICORICE enables test-time concept intervention.
+LICORICE enables test-time concept intervention.
+To validate that LICORICE supports test-time concept intervention, we simulate using a noisy concept model $\hat{g}$ with $f$ to study the impact of concept interventions on reward.
+For PixelCartPole, we add Gaussian noise to each predicted concept with a standard deviation of $0.2$.
+For the other three environments with discrete concepts, each concept label is randomly changed with probability $0.2$.
+Following previous work (Koh et al., 2020), we first intervene individually on concepts, setting them to ground truth, and sort the concepts in descending order of reward improvement.
+We then sequentially intervene on the concepts following this ordering, such that any previously intervened concepts remain intervened.
+According to Figure 4, using a noisy concept model significantly degrades reward.
+However, the performance increases with more interventions, meaning LICORICE supports test-time concept intervention, affirmatively answering RQ 4.
+Figure 5: Test-time intervention examples, where intervening on a single concept corrects the action.
+Concept interventions show how the policy decisions change.
+We now show how domain experts could interact with the model.
+Specifically, we simulate a counterfactual question: What if a concept value is incorrectly predicted?
+Figure 5 depicts two examples in DoorKey.
+Just intervening on the concept of the door being open or closed is sufficient to change the agent’s behavior, both highlighting the importance of this specific concept and the ability of a user to intervene on the concepts to interrogate and understand agent behavior.
+Additional experiments
+PixelCartPole DoorKey DynamicObstacles Boxing
+Algorithm $R$uparrow $c$ MSE $\downarrow$ $R$uparrow $c$ Error $\downarrow$ $R$uparrow $c$ Error $\downarrow$ $R$uparrow $c$ Error $\downarrow$
+LICORICE-IT 0.533 8960.5338960.533\,8960.533 896 0.079 6220.0796220.079\,6220.079 622 0.992 4120.9924120.992\,4120.992 412 0.089 5350.0895350.089\,5350.089 535 0.996 8890.9968890.996\,8890.996 889 0.001 2020.0012020.001\,2020.001 202 0.978 1520.9781520.978\,1520.978 152 0.135 6060.1356060.135\,6060.135 606
+LICORICE-DE $0.9669$ 0.031 1060.0311060.031\,1060.031 106 0.783 1130.7831130.783\,1130.783 113 0.200 0220.2000220.200\,0220.200 022 0.980 5370.9805370.980\,5370.980 537 0.000 1140.0001140.000\,1140.000 114 1.001 0451.0010451.001\,0451.001 045 0.044 7890.0447890.044\,7890.044 789
+LICORICE-AC 0.910 7840.9107840.910\,7840.910 784 0.022 3140.0223140.022\,3140.022 314 0.815 7550.8157550.815\,7550.815 755 0.161 8420.1618420.161\,8420.161 842 0.583 0530.5830530.583\,0530.583 053 0.002 2910.0022910.002\,2910.002 291 0.994 3050.9943050.994\,3050.994 305 0.087 9810.0879810.087\,9810.087 981
+LICORICE $1.0$ 0.024 7970.0247970.024\,7970.024 797 0.990 2970.9902970.990\,2970.990 297 0.047 5040.0475040.047\,5040.047 504 0.990 6070.9906070.990\,6070.990 607 0.000 1930.0001930.000\,1930.000 193 1.004 9971.0049971.004\,9971.004 997 0.049 $0$.0490000.049\,$0$.049 000
+Table 3: Ablation study results for LICORICE. All components generally help achieve better reward and lower concept error. Full results with standard deviation are in Table 9, Appendix B.
+Ablation study: all components of LICORICE contribute positively to its performance.
+We now conduct an ablation study of our three main contributions: iterative training, decorrelation, and disagreement-based active learning.
+LICORICE-IT uses only one iteration, LICORICE-DE does not perform data decorrelation, and LICORICE-AC uses the entire unlabeled dataset for querying instead of a subset chosen by active learning.
+All components are critical to achieving both high reward and low concept error (Table 3; learning curves in Appendix B).
+However, the component that most contributes to the performance differs depending on the environment.
+For example, compared with LICORICE, LICORICE-IT exhibits the largest reward gap for PixelCartPole; however, LICORICE-AC yields the largest reward gap for DynamicObstacles, and LICORICE-DE yields the largest gap for DoorKey.
+This difference may occur because the concepts in DynamicObstacles are simple enough that one iteration is sufficient, so the largest gains can be made with active learning.
+In contrast, PixelCartPole requires further policy refinement to better estimate on-distribution concept values, so the largest gains can be made by leveraging multiple iterations.
+LICORICE is robust to various hyperparameter values.
+LICORICE involves a few key hyperparameters, described in §3 and summarized here for convenience.
+In data decorrelation, $\tau$ governs the size of the unlabeled dataset collected by the current policy and $p$ controls the rate of acceptance of data points.
+Active learning involves a concept ensemble with $N$ models.
+We study the effect of the values of these hyperparameters on DynamicObstacles, finding that LICORICE achieves $96-99\%$ of the maximum reward across $2$ values for $N$, $3$ values for $p$, and $3$ values for $\tau$ (Section B.5).
 5 Related Work
-Interpretable Reinforcement Learning Interpretable RL has gained significant attention in recent years [10].
-One prominent approach uses rule-based methods — such as decision trees [26, 29], logic [8], and programs [30, 19] — to represent policies.
+Interpretable RL.
+Interpretable RL has gained significant attention in recent years (Glanois et al., 2024).
+One prominent area uses rule-based methods—such as decision trees (Silva et al., 2020; Topin et al., 2021), logic (Delfosse et al., 2024), and programs (Verma et al., 2018; Penkov & Ramamoorthy, 2019)—to represent policies.
 These works either assume that the state is already interpretable or that the policy is pre-specified.
 Unlike prior work, our method involves learning the interpretable representation (through concept training) for policy learning.
-Concept Learning for Reinforcement Learning Inspired by successes in the supervised setting [6, 25, 34], concept-based explanations have recently been incorporated into RL.
-One approach [7] learns a joint embedding model between state-action pairs and concept-based explanations to expedite learning via reward shaping.
-Unlike our work, their policy is not a strict function of the concepts, allowing our techniques to be combined to provide both concept-based explanations and a concept-based interpretable policy.
-Another example, CPM [33], is a multi-agent RL concept architecture that focuses on trade-offs between interpretability and accuracy.
-They assume that concept labels are available continuously during training.
-As we have shown, this approach uses over $1$M concept labels in our test environments, whereas our approach reduces the need for continuous human intervention, requiring only 500 or fewer concept labels to achieve similar or better performance in single-agent environments.
-Learning Concepts with Human Feedback Another line of work explores how to best leverage human concept labels but not in the RL setting, and does not focus on reducing the labeling burden.
-In contrast, our approach aims to reduce the concept labeling burden during training.
-One work [14] instead has users label additional information about the relevance of certain feature dimensions to the concept label.
-Another work [4] develops an intervention policy at prediction time to choose which concepts to request a label for with the goal of improving the final prediction.
-Future work could explore using these techniques alongside our method.
+Concept Learning for RL.
+Due to successes in the supervised setting (Collins et al., 2023; Sheth & Ebrahimi Kahou, 2023; Zarlenga et al., 2023), concept models have recently been used in RL.
+Das et al.
+(2023) trains a joint embedding model between state-action pairs and concept-based explanations to expedite learning via reward shaping.
+Unlike LICORICE, their policy is not a strict function of the concepts, allowing our techniques to be combined to provide concept-based explanations alongside an interpretable policy.
+Another example, CPM (Zabounidis et al., 2023), is a multi-agent RL approach that assumes that concept labels are continuously available during training.
+As we have shown, this approach uses over $1$M concept labels, whereas LICORICE achieves similar performance while using 2000×2000$\times$2000 × fewer labels.
+Learning Concepts with Human Feedback.
+Prior research has explored leveraging human concept labels, but not for RL and without focusing on reducing the labeling burden.
+For instance, Lage & Doshi-Velez (2020) has users label additional information about the relevance of certain feature dimensions to the concept label to facilitate concept learning for high-dimensional data.
+In a different vein, Chauhan et al.
+(2023) develop a test-time intervention policy for querying for concept labels to improve the final prediction.
+These methods do not directly tackle the issue of reducing the overall labeling burden during training and could be used alongside our method.
+A concurrent work (Shao et al., 2024) also studies sample efficiency but mainly focus on human task demonstrations.
 6 Discussion and Conclusion
-In this work, we proposed LICORICE, a novel RL algorithm that addresses the critical issue of model interpretability while minimizing the reliance on continuous human annotation.
-We introduced a training scheme that enables RL algorithms to learn concepts more efficiently from little to no labeled concept data.
-Our approach interleaves concept learning and RL training, uses an ensemble-based active learning technique to select informative data points for labeling, and uses a simple sampling strategy to better decorrelate the concept data.
+In this work, we proposed LICORICE, a novel RL algorithm that addresses the critical issue of model interpretability while minimizing reliance on continuous human annotation.
+We introduced a training scheme that enables RL algorithms to learn concepts efficiently from little labeled concept data.
 We demonstrated how this approach reduces manual labeling effort.
-Finally, we conducted initial experiments to demonstrate how we can leverage powerful VLMs to infer concepts from raw visual inputs without explicit labels in some environments.
-There are broader societal impacts of our work that must be considered.
-These include both the impacts of using VLMs in real-world applications, as well as considerations around interpretability more generally.
-For a more detailed discussion, please refer to Appendix C.
-Limitations and Future Work Although VLMs can be successfully used for automatic labeling of some concepts, there are still hallucination issues [2] and other failure cases, such as providing inaccurate counts.
-We believe that future work that seeks to improve general VLM capabilities and mitigate hallucinations would also help overcome this limitation.
-Another exciting direction is the work in human-AI complementarity and learning-to-defer algorithms [18] to train an additional classifier for deferring concept labeling to a person when the chance of making an error is high.
-Acknowledgments and Disclosure of Funding
+We conducted initial experiments to demonstrate how we can leverage powerful VLMs to infer concepts from raw visual inputs without explicit labels in some environments.
+There are broader societal impacts of our work that must be considered, including both the impacts of using VLMs in real-world applications and more general considerations around interpretability.
+For a more detailed discussion of limitations and future work, please refer to Appendix C.
+Acknowledgments
 This work was supported in part by NSF grant IIS-2046640 (CAREER).
-Co-author Fang is supported in part by Sloan Research Fellowship.
+Co-author Fang is supported in part by a Sloan Research Fellowship.
 We thank NVIDIA for providing computing resources.
@@
 Appendix A Experimental Result Reproducibility
-In this section, we provided detailed descriptions to achieve reproducibility.
-A.1 Concepts Definitions
+In this section, we provide descriptions of our concept definitions, prompts, and other experimental details toward the goal of reproducibility.
+A.1 Definitions of Concepts
 Environment Concept Name Type Value Ranges GPT-4o Error
-PixelCartPole-v1 Cart Position Continuous $(-2.4,2.4)$ $0.01$\pm$0.0$$
+PixelCartPole Cart Position Continuous $(-2.4,2.4)$ $0.01$\pm$0.0$$
 Cart Velocity Continuous $\mathbb{R}$ $0.31$\pm$0.12$$
@@
 Pole Angular Velocity Continuous $\mathbb{R}$ $0.63$\pm$0.14$$
-DoorKey-7x7 Agent Position x Discrete 5 $0.36$\pm$0.08$$
+DoorKey Agent Position x Discrete 5 $0.36$\pm$0.08$$
 Agent Position y Discrete 5 $0.41$\pm$0.12$$
@@
 Direction Movable Up Discrete 2 $0.2$\pm$0.08$$
-DynamicObstacles-5x5 Agent Position x Discrete 3 $0.03$\pm$0.06$$
+DynamicObstacles Agent Position x Discrete 3 $0.03$\pm$0.06$$
 Agent Position y Discrete 3 $0.04$\pm$0.06$$
@@
 Direction Movable Up Discrete 2 $0.07$\pm$0.08$$
+Boxing Agent Position x at Frame 1 Discrete 160 $0.81$\pm$0.14$$
+Agent Position y at Frame 1 Discrete 210 $0.93$\pm$0.05$$
+Enemy Position x at Frame 1 Discrete 160 $0.78$\pm$0.13$$
+Enemy Position y at Frame 1 Discrete 210 $0.9$\pm$0.07$$
+Agent Position x at Frame 2 Discrete 160 $0.67$\pm$0.06$$
+Agent Position y at Frame 2 Discrete 210 $0.88$\pm$0.04$$
+Enemy Position x at Frame 2 Discrete 160 $0.71$\pm$0.05$$
+Enemy Position y at Frame 2 Discrete 210 $0.83$\pm$0.08$$
+Pong Ball Position x Discrete 176 $0.99$\pm$0.00$$
+Ball Position y Discrete 88 $0.98$\pm$0.00$$
+Ball Velocity x Discrete 176 $0.95$\pm$0.00$$
+Ball Velocity y Discrete 176 $0.80$\pm$0.00$$
+Paddle Velocity Discrete 176 $0.78$\pm$0.05$$
+Paddle Acceleration Discrete 352 $0.79$\pm$0.04$$
+Paddle Jerk Discrete 704 $0.83$\pm$0.04$$
 Table 4: Concepts and their possible values for all environments.
 For discrete concepts, we report the number of categories.
-We also provide the mean GPT-4o labeling error (MSELoss for continuous values and 1 - accuracy for discrete ones) for each single concept, averaged across 5 seeds and corresponding to the same evaluation protocols as the ones with largest budgets in Table 8.
+We also provide the mean GPT-4o labeling error for each single concept, averaged across 5 seeds and corresponding to the same evaluation protocols as the ones with the largest budgets in Table 7.
+We use MSELoss for continuous values and 1 - accuracy for discrete ones, and in Boxing, we regard one prediction correct if it is within distance 5 to the ground truth.
 The $\pm$ [value] part shows the standard deviation.
-The average errors over concepts are $0.24$\pm$0.06$$, $0.3$\pm$0.02$$, and $0.13$\pm$0.03$$ respectively for the three environments.
-Table 5 provides more details on the concepts used in each environment, categorizing them by their names, types, and value ranges.
-For the PixelCartPole-v1 environment, all concepts such as Cart Position, Cart Velocity, Pole Angle, and Pole Angular Velocity are continuous.
-In contrast, the DoorKey-7x7 environment features discrete concepts like Agent Position (x and y), Key Position (x and y), and Door Open status, each with specific value ranges.
-Similarly, the DynamicObstacles-5x5 environment lists discrete concepts, including Agent and Obstacle positions, with corresponding value ranges.
-Figure 2: Example start configurations of the DoorKey-7x7 environment.
-This shows that concepts must general enough to apply to many different configurations.
+The average errors over concepts are $0.24$\pm$0.06$$, $0.3$\pm$0.02$$, $0.13$\pm$0.03$$, $0.82$\pm$0.06$$, and $0.88\pm 0.02$ respectively for the environments.
+Table 4 provides more details regarding the concepts used in each environment, categorizing them by their names, types, and value ranges.
+For the PixelCartPole environment, all concepts such as Cart Position, Cart Velocity, Pole Angle, and Pole Angular Velocity are continuous.
+In contrast, the DoorKey environment features discrete concepts like Agent Position (x and y), Key Position (x and y), and Door Open status, each with specific value ranges.
+Similarly, the DynamicObstacles environment lists discrete concepts, including Agent and Obstacle positions, with corresponding value ranges.
+Precise concept definitions are crucial for enabling correct agent behavior.
+We visualize the start configurations for DoorKey in Figure 6 to illustrate the importance of well-defined concepts in reinforcement learning.
+DoorKey includes a variety of initial states, each with different positions of the agent, key, and door.
+This diversity highlights a challenge in concept-based RL: how to define concepts that are both specific enough to be meaningful and general enough to apply across all possible scenarios.
+An agent trained with poorly defined concepts might perform well in some configurations but fail to generalize to others, leading to suboptimal performance in new or unseen environments.
+As a result, concept engineering is its own separate but important problem.
+Figure 6: Example start configurations of the DoorKey environment, along with the associated concepts for each configuration.
+This shows that concepts must be general enough to apply to many different configurations.
 As a result, the agent cannot memorize the exact position of the door and key.
-We also visualize the start configurations for DoorKey-7x7 in Figure 2 to illustrate the importance of concept definitions.
-As shown, the concepts must be defined in such a way to allow the agent to generalize to all possible environment configurations.
-A.2 LICORICE Details
+A.2 LICORICE Details.
 In this section, we provide additional implementation details for LICORICE.
@@
 For discrete concept values, since different concepts have different numbers of categories, we create one linear classification head for each single concept, and to predict the final action, we calculate the class with the largest predicted probability for each concept.
-Value-Based Methods
+Architecture.
+The network begins with a CNN-based feature extractor $f_{\theta}:\mathcal{X}\to\mathbb{R}^{d}$, comprising three convolutional layers that map input state $x\in\mathcal{X}$ to a $d$-dimensional feature vector $h=f_{\theta}(x)$.
+We then introduce a linear layer $g_{\phi}:\mathbb{R}^{d}\to\mathbb{R}^{k}$ for concept prediction ($g$), mapping extracted features to $k$ concept values.
+For regression tasks, each of the $k$ predicted concept is represented as a real value; for classification tasks, each concept maps to one categorical value derived from a classification head.
+The action prediction component ($f$) is implemented as an MLP with two fully connected layers, each containing 64 neurons with Tanh activation, taking only $c$ as input to produce action logits $a=\text{MLP}(c)$.
+Notably, we share the CNN feature extractor $f_{\theta}$ between policy and value functions, a decision informed by improved performance observed in preliminary experiments.
+Figure 7: Architecture of our concept-bottleneck actor-critic method.
+Value-Based Methods.
 If we were to use a value-based method as the RL backbone, we would need to make the following changes.
@@
 Then, we can conduct interleaved training in a similar way to LICORICE.
-Feature Extractor If we use the actor-critic paradigm, we propose to share a feature extractor between the policy and value networks, shown in Figure 3.
+Feature Extractor.
+If we use the actor-critic paradigm, we propose to share a feature extractor between the policy and value networks, shown in Figure 7.
 Intuitively, this choice can offer several advantages compared with using image or predicted concepts as input for both networks.
-Figure 3: Architecture of our concept-bottleneck actor-critic method.
 Sharing a feature extractor enables both networks to benefit from a common, rich representation of the input data, reducing the number of parameters to be trained.
@@
     Agent Direction: Specify the direction the red triangle is facing, which is the orientation of the vertex (pointy corner) of the isosceles triangle.
-Choose from ’right’, ’down’, ’left’, or ’up’.
+Choose from ‘right’, ‘down’, ‘left’, or ‘up’.
 Clarify that this direction is independent of movement options.
@@
     • Status: Assess whether the door is open or closed (closed means the door is visible as a whole yellow cell, while open means the door disappears visually).
-Report as ’true’ for open and ’false’ for closed.
+Report as ‘true’ for open and ‘false’ for closed.
 Double-check the door’s appearance to confirm if it is open or closed.
@@
     • Up (y - 1): Check the cell above.
-    Each direction’s feasibility should be reported as ’true’ if clear and within the grid, and ’false’ otherwise.
+    Each direction’s feasibility should be reported as ‘true’ if clear and within the grid, and ‘false’ otherwise.
 […]
@@
     Agent Direction: Specify the direction the red triangle is facing, which is the orientation of the vertex (pointy corner) of the isosceles triangle.
-Choose from ’right’, ’down’, ’left’, or ’up’.
+Choose from ‘right’, ‘down’, ‘left’, or ‘up’.
 Clarify that this direction is independent of movement options.
@@
     • Up (y - 1): Check the cell above.
-    Each direction’s feasibility should be reported as ’true’ if clear and within the grid, and ’false’ otherwise.
+    Each direction’s feasibility should be reported as ‘true’ if clear and within the grid, and ‘false’ otherwise.
+[…]
+[…]
+  • The white player’s x and y coordinates
+  • The black player’s x and y coordinates
+Please carefully determine the following values and give concise answers one by one. Make sure to return an estimated value for each one, even if the task may look challenging.
+Follow the reporting format:
+  • frame 1 white x: estimated_value
+  • frame 1 white y: estimated_value
+  • frame 1 black x: estimated_value
+  • frame 1 black y: estimated_value
+  • frame 2 white x: estimated_value
+  • frame 2 white y: estimated_value
+  • frame 2 black x: estimated_value
+  • frame 2 black y: estimated_value
+[…]
+[…]
+Please determine the following values and give concise answers one by one. Make sure to return an estimated value for each one, even if the task may look challenging.
+  • x coordinate of the ball relative to the paddle in the first frame
+  • y coordinate of the ball in the first frame
+  • x velocity of the ball calculated from the first and second frames
+  • y velocity of the ball calculated from the first and second frames
+  • velocity of the paddle calculated from the first and second frames
+  • acceleration of the paddle calculated from the first three frames
+  • jerk of the paddle calculated from the four frames
+Follow the reporting format in your response.
+  • x coordinate of the ball: estimated_value
+  • y coordinate of the ball: estimated_value
+  • x velocity of the ball: estimated_value
+  • y velocity of the ball: estimated_value
+  • velocity of the paddle: estimated_value
+  • acceleration of the paddle: estimated_value
+  • jerk of the paddle: estimated_value
 […]
 A.4 Experimental Details
-For the PPO hyperparameters, we set $4\cdot 10^{6}$ total timesteps for PixelCartPole-v1and DoorKey-7x7, and $10^{6}$ for DynamicObstacles-5x5.
-Besides that, for all environments, we use $8$ vectorized environments, horizon $T=4096$, $10$ epochs for training, batch size of $512$, learning rate $3\cdot 10^{-4}$, entropy coefficient $0.01$, and value function coefficient $1$.
-For all other hyperparameters, we use the default values from Stable Baselines 3 [23].
-For the concept training, we set $100$ epochs with Adam optimizer with the learning rate linearly decaying from $3\cdot 10^{-4}$ to 00 for each iteration in PixelCartPole-v1.
-In DoorKey-7x7 and DynamicObstacles-5x5, we use the same optimizer and initial learning rate, yet set $50$ epochs instead and set early stopping with threshold linearly increasing from $10$ to $20$, to incentivize the concept network not to overfit in earlier iterations.
+Behavior learning hyperparameters.
+For the PPO hyperparameters, we set $4\cdot 10^{6}$ total timesteps for PixelCartPole and DoorKey, $10^{6}$ for DynamicObstacles, $1.5\cdot 10^{7}$ for Boxing, and $10^{7}$ for Pong.
+For PixelCartPole, DoorKey, and DynamicObstacles, we use $8$ vectorized environments, horizon $T=4096$, $10$ epochs for training, batch size of $512$, learning rate $3\cdot 10^{-4}$, entropy coefficient $0.01$, and value function coefficient $0.5$.
+For Boxing and Pong, we use $8$ vectorized environments, horizon $T=1024$, $4$ epochs for training, batch size of $256$, learning rate $3\cdot 10^{-4}$, entropy coefficient $0.01$, and value function coefficient $0.5$.
+For all other hyperparameters, we use the default values from Stable Baselines 3 Raffin et al.
+(2021).
+Concept learning hyperparameters.
+For the concept training, we set $100$ epochs with Adam optimizer with the learning rate linearly decaying from $3\cdot 10^{-4}$ to 00 for each iteration in PixelCartPole, Boxing, and Pong.
+In DoorKey and DynamicObstacles, we use the same optimizer and initial learning rate, yet set $50$ epochs instead and set early stopping with threshold linearly increasing from $10$ to $20$, to incentivize the concept network not to overfit in earlier iterations.
 The batch size is $32$.
-For LICORICE, we set the sample acceptance rate $p=0.05$, the ratio for active learning $\tau=10$, batch size to query labels in the active learning module $b=20$, and the number of ensemble models $N=5$.
-The number of iterations chosen in our algorithm is $M=4$ for PixelCartPole-v1and $M=2$ for the other two environments.
+We model concept learning for PixelCartPole as a regression problem (minimizing mean squared error).
+We model concept learning for DoorKey, DynamicObstacles, Boxing,and Pong as classification problems.
+LICORICE-specific hyperparameters.
+For LICORICE, we set the ratio for active learning $\tau=10$, batch size to query labels in the active learning module $b=20$, and the number of ensemble models $N=5$ for the first three environments.
+For the complex environments Boxing and Pong, we choose the ratio for active learning $\tau=4$, batch size to query labels in the active learning module $b=B_{m}/5$ and number of ensemble models $N=5$ to make a balance between performance and speed.
+The default number of iterations chosen in our algorithm is $M=4$ for PixelCartPole, $M=2$ for DoorKey, DynamicObstacles, and $M=5$ for Boxing and Pong.
+The sample acceptance rate $p=0.02$ for PixelCartPoleand $p=0.1$ for the other four environments.
+The Random-Q baseline uses the same sample acceptance rate $p$ as LICORICE.
+In the complex environments Boxing and Pong, we use both the KL-divergence penalty and PPO loss at the end of the algorithm to improve optimization with $\beta=0.01$, as mentioned in Section 3.
+Computational resources.
 We use NVIDIA A6000 and NVIDIA RTX 6000 Ada Generation.
-Each of our programs uses less than $2$GB GPU memory.
-For PixelCartPole-v1 and DoorKey-7x7, each run takes less than 9 hours to finish.
-For DynamicObstacles-5x5, each run takes less than 2 hours to finish.
-For all experiments, we consistently use 5 seeds [$123$, $456$, $789$, $1011$, $1213$] to train the models.
+Each of our training program uses less than 3GB GPU memory.
+For Boxingand Pong, each run takes less than 20 hours to finish.
+For PixelCartPole and DoorKey, each run takes less than 9 hours to finish.
+For DynamicObstacles, each run takes less than 2 hours to finish.
+Experimental replicability.
+For all experiments, we use 5 seeds [$123$, $456$, $789$, $1011$, $1213$] to train the models.
 We then evaluate on the environment with seed $42$ with 100 episodes and take the average of the reward across the episodes and the concept error across the observations within each episode.
 Appendix B Additional Results
+In this section, we present both the numerical results as tables from the figures in the main paper, in addition to experimental results not reported in the main paper.
 B.1 Balancing Concept Performance and Environment Reward
 Algorithm $R$uparrow $c$ Error $\downarrow$
-PixelCartPole-v1 Sequential-Q 0.232 472±0.085 242plus-or-minus0.2324720.085242\$0.232\,472$\pm$\$0.085\,242\$0.232 472 ± 0.085 242 0.102 57±0.037 255plus-or-minus0.102570.037255\$0.102\,57$\pm$\$0.037\,255\$0.102 57 ± 0.037 255
+PixelCartPole Sequential-Q 0.232 472±0.085 242plus-or-minus0.2324720.085242\$0.232\,472$\pm$\$0.085\,242\$0.232 472 ± 0.085 242 0.102 57±0.037 255plus-or-minus0.102570.037255\$0.102\,57$\pm$\$0.037\,255\$0.102 57 ± 0.037 255
 Disagreement-Q 0.363 624±0.200 067plus-or-minus0.3636240.200067\$0.363\,624$\pm$\$0.200\,067\$0.363 624 ± 0.200 067 0.097 457±0.036 755plus-or-minus0.0974570.036755\$0.097\,457$\pm$\$0.036\,755\$0.097 457 ± 0.036 755
@@
 LICORICE $1.0$\pm$0.0$$ 0.024 797±0.004 166plus-or-minus0.0247970.004166\$0.024\,797$\pm$\$0.004\,166\$0.024 797 ± 0.004 166
+LICORICE+GPT-4o 0.064 94±0.010 829 728plus-or-minus0.064940.010829728\$0.064\,94$\pm$\$0.010\,829\,728\$0.064 94 ± 0.010 829 728 0.246 895 733±0.077 473 22plus-or-minus0.2468957330.07747322\$0.246\,895\,733$\pm$\$0.077\,473\,22\$0.246 895 733 ± 0.077 473 22
 CPM $1.0$\pm$0.0$$ 0.011 545±0.000 782plus-or-minus0.0115450.000782\$0.011\,545$\pm$\$0.000\,782\$0.011 545 ± 0.000 782
-DoorKey-7x7 Sequential-Q 0.462 566±0.091 999plus-or-minus0.4625660.091999\$0.462\,566$\pm$\$0.091\,999\$0.462 566 ± 0.091 999 0.462 327±0.045 955plus-or-minus0.4623270.045955\$0.462\,327$\pm$\$0.045\,955\$0.462 327 ± 0.045 955
+DoorKey Sequential-Q 0.462 566±0.091 999plus-or-minus0.4625660.091999\$0.462\,566$\pm$\$0.091\,999\$0.462 566 ± 0.091 999 0.462 327±0.045 955plus-or-minus0.4623270.045955\$0.462\,327$\pm$\$0.045\,955\$0.462 327 ± 0.045 955
 Disagreement-Q 0.759 464±0.117 38plus-or-minus0.7594640.11738\$0.759\,464$\pm$\$0.117\,38\$0.759 464 ± 0.117 38 0.392 394±0.073 599plus-or-minus0.3923940.073599\$0.392\,394$\pm$\$0.073\,599\$0.392 394 ± 0.073 599
@@
 LICORICE 0.990 297±0.007 106plus-or-minus0.9902970.007106\$0.990\,297$\pm$\$0.007\,106\$0.990 297 ± 0.007 106 0.047 504±0.011 097plus-or-minus0.0475040.011097\$0.047\,504$\pm$\$0.011\,097\$0.047 504 ± 0.011 097
+LICORICE+GPT-4o 0.830 113±0.056 988plus-or-minus0.8301130.056988\$0.830\,113$\pm$\$0.056\,988\$0.830 113 ± 0.056 988 0.312 288±0.014 362plus-or-minus0.3122880.014362\$0.312\,288$\pm$\$0.014\,362\$0.312 288 ± 0.014 362
 CPM 1.001 024±0.000 06plus-or-minus1.0010240.00006\$1.001\,024$\pm$\$0.000\,06\$1.001 024 ± 0.000 06 $0.0$\pm$0.0$$
-DynamicObstacles-5x5 Sequential-Q 0.788 348±0.440 724plus-or-minus0.7883480.440724\$0.788\,348$\pm$\$0.440\,724\$0.788 348 ± 0.440 724 0.014 136±0.009 627plus-or-minus0.0141360.009627\$0.014\,136$\pm$\$0.009\,627\$0.014 136 ± 0.009 627
+DynamicObstacles Sequential-Q 0.788 348±0.440 724plus-or-minus0.7883480.440724\$0.788\,348$\pm$\$0.440\,724\$0.788 348 ± 0.440 724 0.014 136±0.009 627plus-or-minus0.0141360.009627\$0.014\,136$\pm$\$0.009\,627\$0.014 136 ± 0.009 627
 Disagreement-Q 0.992 53±0.014 194plus-or-minus0.992530.014194\$0.992\,53$\pm$\$0.014\,194\$0.992 53 ± 0.014 194 0.000 779±0.000 458plus-or-minus0.0007790.000458\$0.000\,779$\pm$\$0.000\,458\$0.000 779 ± 0.000 458
@@
 LICORICE 0.990 607±0.003 563plus-or-minus0.9906070.003563\$0.990\,607$\pm$\$0.003\,563\$0.990 607 ± 0.003 563 0.000 193±0.000 432plus-or-minus0.0001930.000432\$0.000\,193$\pm$\$0.000\,432\$0.000 193 ± 0.000 432
+LICORICE+GPT-4o 0.875 015±0.088 616plus-or-minus0.8750150.088616\$0.875\,015$\pm$\$0.088\,616\$0.875 015 ± 0.088 616 0.132 22±0.025 588plus-or-minus0.132220.025588\$0.132\,22$\pm$\$0.025\,588\$0.132 22 ± 0.025 588
 CPM 0.969 577±0.021 855plus-or-minus0.9695770.021855\$0.969\,577$\pm$\$0.021\,855\$0.969 577 ± 0.021 855 $0.0$\pm$0.0$$
+Boxing Sequential-Q 0.865 427±0.195 112plus-or-minus0.8654270.195112\$0.865\,427$\pm$\$0.195\,112\$0.865 427 ± 0.195 112 0.318 058±0.260 921plus-or-minus0.3180580.260921\$0.318\,058$\pm$\$0.260\,921\$0.318 058 ± 0.260 921
+Disagreement-Q 0.825 683±0.143 459plus-or-minus0.8256830.143459\$0.825\,683$\pm$\$0.143\,459\$0.825 683 ± 0.143 459 0.201 04±0.212 373plus-or-minus0.201040.212373\$0.201\,04$\pm$\$0.212\,373\$0.201 04 ± 0.212 373
+Random-Q 0.957 467±0.039 969plus-or-minus0.9574670.039969\$0.957\,467$\pm$\$0.039\,969\$0.957 467 ± 0.039 969 0.142 976±0.099 212plus-or-minus0.1429760.099212\$0.142\,976$\pm$\$0.099\,212\$0.142 976 ± 0.099 212
+LICORICE 1.004 997±0.017 713plus-or-minus1.0049970.017713\$1.004\,997$\pm$\$0.017\,713\$1.004 997 ± 0.017 713 0.049±0.010 967plus-or-minus0.0490.010967\$0.049$\pm$\$0.010\,967\$0.049 ± 0.010 967
+LICORICE+GPT-4o 0.221 577±0.009 677plus-or-minus0.2215770.009677\$0.221\,577$\pm$\$0.009\,677\$0.221 577 ± 0.009 677 0.976 101±0.009 001plus-or-minus0.9761010.009001\$0.976\,101$\pm$\$0.009\,001\$0.976 101 ± 0.009 001
+CPM 1.0±0.052 538plus-or-minus1.00.052538\$1.0$\pm$\$0.052\,538\$1.0 ± 0.052 538 0.000 109±0.000 103plus-or-minus0.0001090.000103\$0.000\,109$\pm$\$0.000\,103\$0.000 109 ± 0.000 103
+Pong Sequential-Q 0.98±0.017 924plus-or-minus0.980.017924\$0.98$\pm$\$0.017\,924\$0.98 ± 0.017 924 0.536 369±0.034 322plus-or-minus0.5363690.034322\$0.536\,369$\pm$\$0.034\,322\$0.536 369 ± 0.034 322
+Disagreement-Q 0.942 857±0.092 857plus-or-minus0.9428570.092857\$0.942\,857$\pm$\$0.092\,857\$0.942 857 ± 0.092 857 0.566 152±0.053 357plus-or-minus0.5661520.053357\$0.566\,152$\pm$\$0.053\,357\$0.566 152 ± 0.053 357
+Random-Q 0.942 143±0.045 617plus-or-minus0.9421430.045617\$0.942\,143$\pm$\$0.045\,617\$0.942 143 ± 0.045 617 0.604 967±0.043 498plus-or-minus0.6049670.043498\$0.604\,967$\pm$\$0.043\,498\$0.604 967 ± 0.043 498
+LICORICE 0.996 333±0.018 424plus-or-minus0.9963330.018424\$0.996\,333$\pm$\$0.018\,424\$0.996 333 ± 0.018 424 0.243 221±0.162 857plus-or-minus0.2432210.162857\$0.243\,221$\pm$\$0.162\,857\$0.243 221 ± 0.162 857
+LICORICE+GPT-4o 0.126 19±0.170 042plus-or-minus0.126190.170042\$0.126\,19$\pm$\$0.170\,042\$0.126 19 ± 0.170 042 0.706 473±0.009 016plus-or-minus0.7064730.009016\$0.706\,473$\pm$\$0.009\,016\$0.706 473 ± 0.009 016
+CPM 0.997 222±0.004 811plus-or-minus0.9972220.004811\$0.997\,222$\pm$\$0.004\,811\$0.997 222 ± 0.004 811 0.004 985±0.007 178plus-or-minus0.0049850.007178\$0.004\,985$\pm$\$0.007\,178\$0.004 985 ± 0.007 178
 Table 5: Evaluation of the reward $R$ and concept error achieved by all methods in all environments.
-This is a extended table from Table 1.
+This is a extended table from Figure 2.
 The reward is reported as the fraction of the reward upper bound.
-For PixelCartPole-v1, the $c$ error is the MSE.
-For the other two environments, the $c$ error is 1 - accuracy.
-The first four algorithms are given a budget of $B=[500,300,300]$ for each environment, from top to bottom; CPM is given an unlimited budget (in practice, it uses 4M, 4M, 1M concept labels respectively).
+For PixelCartPole, the $c$ error is the MSE.
+For the other environments, the $c$ error is 1 - accuracy.
+The first five algorithms are given a budget of $B=[500,300,300,3000,5000]$ for each environment, from top to bottom; CPM is given an unlimited budget (in practice, it uses 4M, 4M, 1M, 15M, 10M concept labels respectively).
 The $\pm$ [value] part shows the standard deviation.
-This shows a more complete version of the results in Table 1.
-In Table 6, we present an extension of the results in Table 1, including standard deviation. Our method enjoys low variance across all environments in terms of both concept error and reward.
+For each environment, we bold the highest reward and lowest concept error among the budget-constrained methods.
+In Table 5, we present all of the numerical results from Figure 2, including standard deviation. Our method enjoys low variance across all environments in terms of both concept error and reward.
 B.2 Budget Allocation Effectiveness
-$B$ $M$ $R$uparrow $c$ Error $\downarrow$
-PixelCartPole-v1 300 1 0.276 344±0.104 789plus-or-minus0.2763440.104789\$0.276\,344$\pm$\$0.104\,789\$0.276 344 ± 0.104 789 0.153 573±0.065 982plus-or-minus0.1535730.065982\$0.153\,573$\pm$\$0.065\,982\$0.153 573 ± 0.065 982
+Environment $B$ $M$ $R$uparrow $c$ Error $\downarrow$
+PixelCartPole 300 1 0.276 344±0.104 789plus-or-minus0.2763440.104789\$0.276\,344$\pm$\$0.104\,789\$0.276 344 ± 0.104 789 0.153 573±0.065 982plus-or-minus0.1535730.065982\$0.153\,573$\pm$\$0.065\,982\$0.153 573 ± 0.065 982
 400 3 0.766 288±0.224 139plus-or-minus0.7662880.224139\$0.766\,288$\pm$\$0.224\,139\$0.766 288 ± 0.224 139 0.052 458±0.012 336plus-or-minus0.0524580.012336\$0.052\,458$\pm$\$0.012\,336\$0.052 458 ± 0.012 336
 500 4 $1.0$\pm$0.0$$ 0.024 797±0.004 166plus-or-minus0.0247970.004166\$0.024\,797$\pm$\$0.004\,166\$0.024 797 ± 0.004 166
-DoorKey-7x7 100 1 0.765 391±0.039 967plus-or-minus0.7653910.039967\$0.765\,391$\pm$\$0.039\,967\$0.765 391 ± 0.039 967 0.260 404±0.041 564plus-or-minus0.2604040.041564\$0.260\,404$\pm$\$0.$0 4$ 0.041 564
+DoorKey 100 1 0.765 391±0.039 967plus-or-minus0.7653910.039967\$0.765\,391$\pm$\$0.039\,967\$0.765 391 ± 0.039 967 0.260 404±0.041 564plus-or-minus0.2604040.041564\$0.260\,404$\pm$\$0.$0 4$ 0.041 564
 200 2 0.934 532±0.015 966plus-or-minus0.9345320.015966\$0.934\,532$\pm$\$0.015\,966\$0.934 532 ± 0.015 966 0.086 635±0.009 289plus-or-minus0.0866350.009289\$0.086\,635$\pm$\$0.009\,289\$0.086 635 ± 0.009 289
 300 2 0.990 297±0.007 106plus-or-minus0.9902970.007106\$0.990\,297$\pm$\$0.007\,106\$0.990 297 ± 0.007 106 0.047 504±0.011 097plus-or-minus0.0475040.011097\$0.047\,504$\pm$\$0.011\,097\$0.047 504 ± 0.011 097
-DynamicObstacles-5x5 100 1 0.957 439±0.020 369plus-or-minus0.9574390.020369\$0.957\,439$\pm$\$0.020\,369\$0.957 439 ± 0.020 369 0.040 794±0.006 774plus-or-minus0.0407940.006774\$0.040\,794$\pm$\$0.006\,774\$0.040 794 ± 0.006 774
+DynamicObstacles 100 1 0.957 439±0.020 369plus-or-minus0.9574390.020369\$0.957\,439$\pm$\$0.020\,369\$0.957 439 ± 0.020 369 0.040 794±0.006 774plus-or-minus0.0407940.006774\$0.040\,794$\pm$\$0.006\,774\$0.040 794 ± 0.006 774
 200 1 0.979 819±0.014 765plus-or-minus0.9798190.014765\$0.979\,819$\pm$\$0.014\,765\$0.979 819 ± 0.014 765 0.007 834±0.004 337plus-or-minus0.0078340.004337\$0.007\,834$\pm$\$0.004\,337\$0.007 834 ± 0.004 337
 300 2 0.990 607±0.003 563plus-or-minus0.9906070.003563\$0.990\,607$\pm$\$0.003\,563\$0.990 607 ± 0.003 563 0.000 193±0.000 432plus-or-minus0.0001930.000432\$0.000\,193$\pm$\$0.000\,432\$0.000 193 ± 0.000 432
+Boxing 1000 5 0.942 475±0.050 04plus-or-minus0.9424750.05004\$0.942\,475$\pm$\$0.050\,04\$0.942 475 ± 0.050 04 0.108 584±0.008 43plus-or-minus0.1085840.00843\$0.108\,584$\pm$\$0.008\,43\$0.108 584 ± 0.008 43
+2000 5 0.980 631±0.047 948plus-or-minus0.9806310.047948\$0.980\,631$\pm$\$0.047\,948\$0.980 631 ± 0.047 948 0.098 675±0.059 967plus-or-minus0.0986750.059967\$0.098\,675$\pm$\$0.059\,967\$0.098 675 ± 0.059 967
+3000 5 1.004 997±0.017 713plus-or-minus1.0049970.017713\$1.004\,997$\pm$\$0.017\,713\$1.004 997 ± 0.017 713 0.049±0.010 967plus-or-minus0.0490.010967\$0.049$\pm$\$0.010\,967\$0.049 ± 0.010 967
+Pong 1000 5 0.994 762±0.002 988plus-or-minus0.9947620.002988\$0.994\,762$\pm$\$0.002\,988\$0.994 762 ± 0.002 988 0.647 428±0.041 13plus-or-minus0.6474280.04113\$0.647\,428$\pm$\$0.041\,13\$0.647 428 ± 0.041 13
+3000 5 0.995 238±0.004 836plus-or-minus0.9952380.004836\$0.995\,238$\pm$\$0.004\,836\$0.995 238 ± 0.004 836 0.428 171±0.046 414plus-or-minus0.4281710.046414\$0.428\,171$\pm$\$0.046\,414\$0.428 171 ± 0.046 414
+5000 5 0.996 333±0.018 424plus-or-minus0.9963330.018424\$0.996\,333$\pm$\$0.018\,424\$0.996 333 ± 0.018 424 0.243 221±0.162 857plus-or-minus0.2432210.162857\$0.243\,221$\pm$\$0.162\,857\$0.243 221 ± 0.162 857
 Table 6: Performance of LICORICE on all environments for varying budgets.
-We select $M$ through grid search, optimizing $R$ $c$ error without additional weighting since they are observed to have the same magnitude.
+We select $M$ to optimize $R$ $c$ error without additional weighting since they are observed to have the same magnitude.
 The reward is reported as the fraction of the reward upper bound.
-For PixelCartPole-v1, $c$ error is MSE; for the other environments, $c$ error is 1 - accuracy.
+For PixelCartPole, $c$ error is MSE; for the other environments, $c$ error is 1 - accuracy.
 The $\pm$ [value] part shows the standard deviation.
-This shows a more complete version of the results in Table 2.
-In Table 7, we present an extension of the results in Table 2, including the standard deviation.
+This shows a more complete version of the results in Figure 3.
+In Table 6, we present an extension of the results in Figure 3, including the standard deviation.
 As expected, as the budget increases, the standard deviation for both the reward and concept error tends to decrease.
-The one exception is the reward for PixelCartPole-v1.
+The one exception is the reward for PixelCartPole.
 Interestingly, the standard deviation is highest for $B=400$.
@@
 Environment $B$ $M$ $R$uparrow LICORICE+GPT-4o $c$ Error $\downarrow$ GPT-4o $c$ Error
-PixelCartPole-v1 300 1 0.064 712±0.017 953plus-or-minus0.0647120.017953\$0.064\,712$\pm$\$0.017\,953\$0.064 712 ± 0.017 953 0.170 068±0.188 407plus-or-minus0.1700680.188407\$0.170\,068$\pm$\$0.188\,407\$0.170 068 ± 0.188 407 0.193 75±0.248 575plus-or-minus0.193750.248575\$0.193\,75$\pm$\$0.248\,575\$0.193 75 ± 0.248 575
+PixelCartPole 300 1 0.064 712±0.017 953plus-or-minus0.0647120.017953\$0.064\,712$\pm$\$0.017\,953\$0.064 712 ± 0.017 953 0.170 068±0.188 407plus-or-minus0.1700680.188407\$0.170\,068$\pm$\$0.188\,407\$0.170 068 ± 0.188 407 0.193 75±0.248 575plus-or-minus0.193750.248575\$0.193\,75$\pm$\$0.248\,575\$0.193 75 ± 0.248 575
 400 2 0.066 216±0.021 619plus-or-minus0.0662160.021619\$0.066\,216$\pm$\$0.021\,619\$0.066 216 ± 0.021 619 0.246 087±0.147 974plus-or-minus0.2460870.147974\$0.246\,087$\pm$\$0.147\,974\$0.246 087 ± 0.147 974 0.224 683±0.151 258plus-or-minus0.2246830.151258\$0.224\,683$\pm$\$0.151\,258\$0.224 683 ± 0.151 258
 500 2 0.064 94±0.010 829 728plus-or-minus0.064940.010829728\$0.064\,94$\pm$\$0.010\,829\,728\$0.064 94 ± 0.010 829 728 0.246 895 733±0.077 473 22plus-or-minus0.2468957330.07747322\$0.246\,895\,733$\pm$\$0.077\,473\,22\$0.246 895 733 ± 0.077 473 22 0.241 365±0.069 363plus-or-minus0.2413650.069363\$0.241\,365$\pm$\$0.069\,363\$0.241 365 ± 0.069 363
-DoorKey-7x7 100 1 0.711 403±0.042 15plus-or-minus0.7114030.04215\$0.711\,403$\pm$\$0.042\,15\$0.711 403 ± 0.042 15 0.450 326±0.036 168plus-or-minus0.4503260.036168\$0.450\,326$\pm$\$0.036\,168\$0.450 326 ± 0.036 168 0.313±0.034 549plus-or-minus0.3130.034549\$0.313$\pm$\$0.034\,549\$0.313 ± 0.034 549
+DoorKey 100 1 0.711 403±0.042 15plus-or-minus0.7114030.04215\$0.711\,403$\pm$\$0.042\,15\$0.711 403 ± 0.042 15 0.450 326±0.036 168plus-or-minus0.4503260.036168\$0.450\,326$\pm$\$0.036\,168\$0.450 326 ± 0.036 168 0.313±0.034 549plus-or-minus0.3130.034549\$0.313$\pm$\$0.034\,549\$0.313 ± 0.034 549
 200 2 0.735 089±0.043 674plus-or-minus0.7350890.043674\$0.735\,089$\pm$\$0.043\,674\$0.735 089 ± 0.043 674 0.329 348±0.010 042plus-or-minus0.3293480.010042\$0.329\,348$\pm$\$0.010\,042\$0.329 348 ± 0.010 042 0.313 667±0.031 697plus-or-minus0.3136670.031697\$0.313\,667$\pm$\$0.031\,697\$0.313 667 ± 0.031 697
 300 2 0.830 113±0.056 988plus-or-minus0.8301130.056988\$0.830\,113$\pm$\$0.056\,988\$0.830 113 ± 0.056 988 0.312 288±0.014 362plus-or-minus0.3122880.014362\$0.312\,288$\pm$\$0.014\,362\$0.312 288 ± 0.014 362 0.304±0.027 654plus-or-minus0.3040.027654\$0.304$\pm$\$0.027\,654\$0.304 ± 0.027 654
-DynamicObstacles-5x5 100 1 0.268 711±0.373 027plus-or-minus0.2687110.373027\$0.268\,711$\pm$\$0.373\,027\$0.268 711 ± 0.373 027 0.200 711±0.0328plus-or-minus0.2007110.0328\$0.200\,711$\pm$\$0.0328\$0.200 711 ± 0.0328 0.159 636±0.029 329plus-or-minus0.1596360.029329\$0.159\,636$\pm$\$0.029\,329\$0.159 636 ± 0.029 329
+DynamicObstacles 100 1 0.268 711±0.373 027plus-or-minus0.2687110.373027\$0.268\,711$\pm$\$0.373\,027\$0.268 711 ± 0.373 027 0.200 711±0.0328plus-or-minus0.2007110.0328\$0.200\,711$\pm$\$0.0328\$0.200 711 ± 0.0328 0.159 636±0.029 329plus-or-minus0.1596360.029329\$0.159\,636$\pm$\$0.029\,329\$0.159 636 ± 0.029 329
 200 1 0.926 095±0.048 753plus-or-minus0.9260950.048753\$0.926\,095$\pm$\$0.048\,753\$0.926 095 ± 0.048 753 0.133 885±0.029 764plus-or-minus0.1338850.029764\$0.133\,885$\pm$\$0.029\,764\$0.133 885 ± 0.029 764 0.123 636±0.046 069plus-or-minus0.1236360.046069\$0.123\,636$\pm$\$0.046\,069\$0.123 636 ± 0.046 069
 300 1 0.875 015±0.088 616plus-or-minus0.8750150.088616\$0.875\,015$\pm$\$0.088\,616\$0.875 015 ± 0.088 616 0.132 22±0.025 588plus-or-minus0.132220.025588\$0.132\,22$\pm$\$0.025\,588\$0.132 22 ± 0.025 588 0.127 636±0.029 385plus-or-minus0.1276360.029385\$0.127\,636$\pm$\$0.029\,385\$0.127 636 ± 0.029 385
+Boxing 1000 5 0.342 437±0.088 819plus-or-minus0.3424370.088819\$0.342\,437$\pm$\$0.088\,819\$0.342 437 ± 0.088 819 0.978 089±0.011 367plus-or-minus0.9780890.011367\$0.978\,089$\pm$\$0.011\,367\$0.978 089 ± 0.011 367 0.781 24±0.005 29plus-or-minus0.781240.00529\$0.781\,24$\pm$\$0.005\,29\$0.781 24 ± 0.005 29
+2000 5 0.228 937±0.107 898plus-or-minus0.2289370.107898\$0.228\,937$\pm$\$0.107\,898\$0.228 937 ± 0.107 898 0.984 122±0.006 578plus-or-minus0.9841220.006578\$0.984\,122$\pm$\$0.006\,578\$0.984 122 ± 0.006 578 0.794 52±0.014 46plus-or-minus0.794520.01446\$0.794\,52$\pm$\$0.014\,46\$0.794 52 ± 0.014 46
+3000 5 0.221 577±0.009 677plus-or-minus0.2215770.009677\$0.221\,577$\pm$\$0.009\,677\$0.221 577 ± 0.009 677 0.976 101±0.009 001plus-or-minus0.9761010.009001\$0.976\,101$\pm$\$0.009\,001\$0.976 101 ± 0.009 001 0.790 14±0.016 773plus-or-minus0.790140.016773\$0.790\,14$\pm$\$0.016\,773\$0.790 14 ± 0.016 773
+Pong 1000 5 0.004 167±0.002 525plus-or-minus0.0041670.002525\$0.004\,167$\pm$\$0.002\,525\$0.004 167 ± 0.002 525 0.794 917±0.083 28plus-or-minus0.7949170.08328\$0.794\,917$\pm$\$0.083\,28\$0.794 917 ± 0.083 28 0.883 29±0.005 45plus-or-minus0.883290.00545\$0.883\,29$\pm$\$0.005\,45\$0.883 29 ± 0.005 45
+3000 5 $0.0$\pm$0.0$$ 0.737 935±0.070 374plus-or-minus0.7379350.070374\$0.737\,935$\pm$\$0.070\,374\$0.737 935 ± 0.070 374 0.860 15±0.031 72plus-or-minus0.860150.03172\$0.860\,15$\pm$\$0.031\,72\$0.860 15 ± 0.031 72
+5000 5 0.126 19±0.170 042plus-or-minus0.126190.170042\$0.126\,19$\pm$\$0.170\,042\$0.126 19 ± 0.170 042 0.706 473±0.009 016plus-or-minus0.7064730.009016\$0.706\,473$\pm$\$0.009\,016\$0.706 473 ± 0.009 016 0.874 07±0.021 957plus-or-minus0.874070.021957\$0.874\,07$\pm$\$0.021\,957\$0.874 07 ± 0.021 957
 Table 7: Performance of LICORICE with GPT-4o integrated into the loop for all environments across different budgets, along with the concept labeling error of GPT-4o as a reference.
-We select $M$ through grid search, optimizing $R$ $c$ error without additional weighting since they are observed to have the same magnitude.
+We select $M$ to optimize $R$ $c$ error without additional weighting since they are observed to have the same magnitude.
 GPT-4o $c$ error is evaluated on a random sample of 50 observations from the same rollout set used for LICORICE+GPT-4o, due to API cost constraints.
-This shows a more complete version of the results in Table 3.
+This shows a more complete version of the results in Figure 3.
+Environment $B$ Sequential-Q Disagreement-Q Random-Q
+$R$uparrow $c$ Error $\downarrow$ $R$uparrow $c$ Error $\downarrow$ $R$uparrow $c$ Error $\downarrow$
+PixelCartPole 300 0.172 72±0.029 967plus-or-minus0.172720.029967\$0.172\,72$\pm$\$0.029\,967\$0.172 72 ± 0.029 967 0.113 868±0.032 088plus-or-minus0.1138680.032088\$0.113\,868$\pm$\$0.032\,088\$0.113 868 ± 0.032 088 0.220 04±0.044 409plus-or-minus0.220040.044409\$0.220\,04$\pm$\$0.044\,409\$0.220 04 ± 0.044 409 0.200 878±0.042 151plus-or-minus0.2008780.042151\$0.200\,878$\pm$\$0.042\,151\$0.200 878 ± 0.042 151 0.243 36±0.070 934plus-or-minus0.243360.070934\$0.243\,36$\pm$\$0.070\,934\$0.243 36 ± 0.070 934 0.137 754±0.044 527plus-or-minus0.1377540.044527\$0.137\,754$\pm$\$0.044\,527\$0.137 754 ± 0.044 527
+400 0.226 48±0.051 626plus-or-minus0.226480.051626\$0.226\,48$\pm$\$0.051\,626\$0.226 48 ± 0.051 626 0.104 433±0.028 986plus-or-minus0.1044330.028986\$0.104\,433$\pm$\$0.028\,986\$0.104 433 ± 0.028 986 0.263 64±0.056 862plus-or-minus0.263640.056862\$0.263\,64$\pm$\$0.056\,862\$0.263 64 ± 0.056 862 0.128 294±0.031 032plus-or-minus0.1282940.031032\$0.128\,294$\pm$\$0.031\,032\$0.128 294 ± 0.031 032 0.2826±0.065 494plus-or-minus0.28260.065494\$0.2826$\pm$\$0.065\,494\$0.2826 ± 0.065 494 0.092 855±0.048 178plus-or-minus0.0928550.048178\$0.092\,855$\pm$\$0.048\,178\$0.092 855 ± 0.048 178
+500 0.227 36±0.065 321plus-or-minus0.227360.065321\$0.227\,36$\pm$\$0.065\,321\$0.227 36 ± 0.065 321 0.102 41±0.028 423plus-or-minus0.102410.028423\$0.102\,41$\pm$\$0.028\,423\$0.102 41 ± 0.028 423 0.386 96±0.283 25plus-or-minus0.386960.28325\$0.386\,96$\pm$\$0.283\,25\$0.386 96 ± 0.283 25 0.086 264±0.026 684plus-or-minus0.0862640.026684\$0.086\,264$\pm$\$0.026\,684\$0.086 264 ± 0.026 684 0.352 44±0.155 268plus-or-minus0.352440.155268\$0.352\,44$\pm$\$0.155\,268\$0.352 44 ± 0.155 268 0.067 721±0.021 766plus-or-minus0.0677210.021766\$0.067\,721$\pm$\$0.021\,766\$0.067 721 ± 0.021 766
+DoorKey 100 0.330 94±0.071 19plus-or-minus0.330940.07119\$0.330\,94$\pm$\$0.071\,19\$0.330 94 ± 0.071 19 0.510 42±0.077 026plus-or-minus0.510420.077026\$0.510\,42$\pm$\$0.077\,026\$0.510 42 ± 0.077 026 0.609 493±0.135 071plus-or-minus0.6094930.135071\$0.609\,493$\pm$\$0.135\,071\$0.609 493 ± 0.135 071 0.325 874±0.070 197plus-or-minus0.3258740.070197\$0.325\,874$\pm$\$0.070\,197\$0.325 874 ± 0.070 197 0.701 847±0.071 458plus-or-minus0.7018470.071458\$0.701\,847$\pm$\$0.071\,458\$0.701 847 ± 0.071 458 0.300 172±0.019 691plus-or-minus0.3001720.019691\$0.300\,172$\pm$\$0.019\,691\$0.300 172 ± 0.019 691
+200 0.427 243±0.093 497plus-or-minus0.4272430.093497\$0.427\,243$\pm$\$0.093\,497\$0.427 243 ± 0.093 497 0.493 205±0.027 703plus-or-minus0.4932050.027703\$0.493\,205$\pm$\$0.027\,703\$0.493 205 ± 0.027 703 0.720 069±0.091 289plus-or-minus0.7200690.091289\$0.720\,069$\pm$\$0.091\,289\$0.720 069 ± 0.091 289 0.275 265±0.073 504plus-or-minus0.2752650.073504\$0.275\,265$\pm$\$0.073\,504\$0.275 265 ± 0.073 504 0.846 412±0.037 362plus-or-minus0.8464120.037362\$0.846\,412$\pm$\$0.037\,362\$0.846 412 ± 0.037 362 0.235 007±0.037 117plus-or-minus0.2350070.037117\$0.235\,007$\pm$\$0.037\,117\$0.235 007 ± 0.037 117
+300 0.509 762±0.110 659plus-or-minus0.5097620.110659\$0.509\,762$\pm$\$0.110\,659\$0.509 762 ± 0.110 659 0.450 234±0.080 155plus-or-minus0.4502340.080155\$0.450\,234$\pm$\$0.080\,155\$0.450 234 ± 0.080 155 0.757 019±0.122 807plus-or-minus0.7570190.122807\$0.757\,019$\pm$\$0.122\,807\$0.757 019 ± 0.122 807 0.411 938±0.055 965plus-or-minus0.4119380.055965\$0.411\,938$\pm$\$0.055\,965\$0.411 938 ± 0.055 965 0.895 01±0.015 27plus-or-minus0.895010.01527\$0.895\,01$\pm$\$0.015\,27\$0.895 01 ± 0.015 27 0.268 422±0.079 03plus-or-minus0.2684220.07903\$0.268\,422$\pm$\$0.079\,03\$0.268 422 ± 0.079 03
+DynamicObstacles 100 0.428 543±0.465 309plus-or-minus0.4285430.465309\$0.428\,543$\pm$\$0.465\,309\$0.428 543 ± 0.465 309 0.044 771±0.037 028plus-or-minus0.0447710.037028\$0.044\,771$\pm$\$0.037\,028\$0.044 771 ± 0.037 028 0.824 501±0.182 298plus-or-minus0.8245010.182298\$0.824\,501$\pm$\$0.182\,298\$0.824 501 ± 0.182 298 0.041 619±0.002 449plus-or-minus0.0416190.002449\$0.041\,619$\pm$\$0.002\,449\$0.041 619 ± 0.002 449 0.335 917±0.460 536plus-or-minus0.3359170.460536\$0.335\,917$\pm$\$0.460\,536\$0.335 917 ± 0.460 536 0.053 889±0.049 941plus-or-minus0.0538890.049941\$0.053\,889$\pm$\$0.049\,941\$0.053 889 ± 0.049 941
+200 0.756 247±0.425 772plus-or-minus0.7562470.425772\$0.756\,247$\pm$\$0.425\,772\$0.756 247 ± 0.425 772 0.029 241±0.032 208plus-or-minus0.0292410.032208\$0.029\,241$\pm$\$0.032\,208\$0.029 241 ± 0.032 208 0.986 525±0.020 203plus-or-minus0.9865250.020203\$0.986\,525$\pm$\$0.020\,203\$0.986 525 ± 0.020 203 0.006 841±0.003 298plus-or-minus0.0068410.003298\$0.006\,841$\pm$\$0.003\,298\$0.006 841 ± 0.003 298 0.758 229±0.424 328plus-or-minus0.7582290.424328\$0.758\,229$\pm$\$0.424\,328\$0.758 229 ± 0.424 328 0.030 011±0.019 832plus-or-minus0.0300110.019832\$0.030\,011$\pm$\$0.019\,832\$0.030 011 ± 0.019 832
+300 0.779 729±0.435 969plus-or-minus0.7797290.435969\$0.779\,729$\pm$\$0.435\,969\$0.779 729 ± 0.435 969 0.012 641±0.009 381plus-or-minus0.0126410.009381\$0.012\,641$\pm$\$0.009\,381\$0.012 641 ± 0.009 381 0.992 074±0.014 955plus-or-minus0.9920740.014955\$0.992\,074$\pm$\$0.014\,955\$0.992 074 ± 0.014 955 0.000 522±0.000 476plus-or-minus0.0005220.000476\$0.000\,522$\pm$\$0.000\,476\$0.000 522 ± 0.000 476 0.969 484±0.014 053plus-or-minus0.9694840.014053\$0.969\,484$\pm$\$0.014\,053\$0.969 484 ± 0.014 053 0.020 619±0.015 441plus-or-minus0.0206190.015441\$0.020\,619$\pm$\$0.015\,441\$0.020 619 ± 0.015 441
+Boxing 1000 0.841 488±0.196 579plus-or-minus0.8414880.196579\$0.841\,488$\pm$\$0.196\,579\$0.841 488 ± 0.196 579 0.346 016±0.1168plus-or-minus0.3460160.1168\$0.346\,016$\pm$\$0.1168\$0.346 016 ± 0.1168 0.931 203±0.053 877plus-or-minus0.9312030.053877\$0.931\,203$\pm$\$0.053\,877\$0.931 203 ± 0.053 877 0.253 524±0.113 155plus-or-minus0.2535240.113155\$0.253\,524$\pm$\$0.113\,155\$0.253 524 ± 0.113 155 0.826 612±0.236 548plus-or-minus0.8266120.236548\$0.826\,612$\pm$\$0.236\,548\$0.826 612 ± 0.236 548 0.252 284±0.046 222plus-or-minus0.2522840.046222\$0.252\,284$\pm$\$0.046\,222\$0.252 284 ± 0.046 222
+2000 0.883 905±0.124 857plus-or-minus0.8839050.124857\$0.883\,905$\pm$\$0.124\,857\$0.883 905 ± 0.124 857 0.188 792±0.106 205plus-or-minus0.1887920.106205\$0.188\,792$\pm$\$0.106\,205\$0.188 792 ± 0.106 205 0.933 527±0.057 27plus-or-minus0.9335270.05727\$0.933\,527$\pm$\$0.057\,27\$0.933 527 ± 0.057 27 0.136 189±0.037 203plus-or-minus0.1361890.037203\$0.136\,189$\pm$\$0.037\,203\$0.136 189 ± 0.037 203 0.832 423±0.109 487plus-or-minus0.8324230.109487\$0.832\,423$\pm$\$0.109\,487\$0.832 423 ± 0.109 487 0.300 118±0.182 217plus-or-minus0.3001180.182217\$0.300\,118$\pm$\$0.182\,217\$0.300 118 ± 0.182 217
+3000 0.865 427±0.195 112plus-or-minus0.8654270.195112\$0.865\,427$\pm$\$0.195\,112\$0.865 427 ± 0.195 112 0.318 058±0.260 921plus-or-minus0.3180580.260921\$0.318\,058$\pm$\$0.260\,921\$0.318 058 ± 0.260 921 0.825 683±0.143 459plus-or-minus0.8256830.143459\$0.825\,683$\pm$\$0.143\,459\$0.825 683 ± 0.143 459 0.201 04±0.212 373plus-or-minus0.201040.212373\$0.201\,04$\pm$\$0.212\,373\$0.201 04 ± 0.212 373 0.957 467±0.039 969plus-or-minus0.9574670.039969\$0.957\,467$\pm$\$0.039\,969\$0.957 467 ± 0.039 969 0.142 976±0.099 212plus-or-minus0.1429760.099212\$0.142\,976$\pm$\$0.099\,212\$0.142 976 ± 0.099 212
+Pong 1000 0.911 429±0.115 539plus-or-minus0.9114290.115539\$0.911\,429$\pm$\$0.115\,539\$0.911 429 ± 0.115 539 0.737 485±0.070 762plus-or-minus0.7374850.070762\$0.737\,485$\pm$\$0.070\,762\$0.737 485 ± 0.070 762 0.836 429±0.205 476plus-or-minus0.8364290.205476\$0.836\,429$\pm$\$0.205\,476\$0.836 429 ± 0.205 476 0.683 161±0.063 495plus-or-minus0.6831610.063495\$0.683\,161$\pm$\$0.063\,495\$0.683 161 ± 0.063 495 0.744 762±0.206 179plus-or-minus0.7447620.206179\$0.744\,762$\pm$\$0.206\,179\$0.744 762 ± 0.206 179 0.734 207±0.051 311plus-or-minus0.7342070.051311\$0.734\,207$\pm$\$0.051\,311\$0.734 207 ± 0.051 311
+3000 0.982 381±0.015 622plus-or-minus0.9823810.015622\$0.982\,381$\pm$\$0.015\,622\$0.982 381 ± 0.015 622 0.576 657±0.081 791plus-or-minus0.5766570.081791\$0.576\,657$\pm$\$0.081\,791\$0.576 657 ± 0.081 791 0.982 857±0.005 746plus-or-minus0.9828570.005746\$0.982\,857$\pm$\$0.005\,746\$0.982 857 ± 0.005 746 0.635 608±0.034 63plus-or-minus0.6356080.03463\$0.635\,608$\pm$\$0.034\,63\$0.635 608 ± 0.034 63 0.962 143±0.041 933plus-or-minus0.9621430.041933\$0.962\,143$\pm$\$0.041\,933\$0.962 143 ± 0.041 933 0.599 326±0.037 327plus-or-minus0.5993260.037327\$0.599\,326$\pm$\$0.037\,327\$0.599 326 ± 0.037 327
+5000 0.98±0.017 924plus-or-minus0.980.017924\$0.98$\pm$\$0.017\,924\$0.98 ± 0.017 924 0.536 369±0.034 322plus-or-minus0.5363690.034322\$0.536\,369$\pm$\$0.034\,322\$0.536 369 ± 0.034 322 0.942 857±0.092 857plus-or-minus0.9428570.092857\$0.942\,857$\pm$\$0.092\,857\$0.942 857 ± 0.092 857 0.566 152±0.053 357plus-or-minus0.5661520.053357\$0.566\,152$\pm$\$0.053\,357\$0.566 152 ± 0.053 357 0.942 143±0.045 617plus-or-minus0.9421430.045617\$0.942\,143$\pm$\$0.045\,617\$0.942 143 ± 0.045 617 0.604 967±0.043 498plus-or-minus0.6049670.043498\$0.604\,967$\pm$\$0.043\,498\$0.604 967 ± 0.043 498
+Table 8: Performance of Sequential-Q, Disagreement-Q, and Random-Q for all environments across different budgets.
+The reward is reported as the fraction of the reward upper bound.
+For PixelCartPole, $c$ error is MSE; for the other environments, $c$ error is 1 - accuracy.
+The $\pm$ [value] part shows the standard deviation.
+This shows a more complete version of the results in Figure 3.
 B.3 Integration with Vision-Language Models
-In Table 8, we present an extension of the results in Table 3 in the main paper, including the standard deviation.
+In Table 7, we present an extension of the results in Figure 3 in the main paper, including the standard deviation.
 Interestingly, the standard deviation for the reward obtained by using LICORICE with GPT-4o as the annotator does not always follow the same trend as shown in Table 6 (when we assume access to a more accurate human annotator).
-Instead, the standard deviation is relatively consistent for PixelCartPole-v1, regardless of the budget.
-It steadily decreases for DoorKey-7x7, as expected.
-However, in DynamicObstacles-5x5, we see an increase when $B=300$.
+Instead, the standard deviation is relatively consistent for PixelCartPole, regardless of the budget.
+It steadily decreases for DoorKey, as expected.
+However, in DynamicObstacles, we see an increase when $B=300$.
 We are not sure of the cause of this.
@@
 Further investigation is required to understand the underlying factors contributing to this anomaly.
-Figure 4: Concept policy model trained with concept labels provided by GPT-4o waits until the coast is clear to move to the goal.
+Figure 8: LICORICE+GPT-4o waits until the coast is clear to move to the goal.
 It appears to make a small mistake in the bottom left, requiring it to wait slightly longer than necessary to navigate to the goal.
-GPT-4o Concept Labeling Errors Table 5 list detailed concept errors for concepts in all environments.
-In PixelCartPole-v1, cart position and pole angle have smaller errors, while velocities require understanding multiple frames and thus are harder to predict accurately.
-For DoorKey-7x7 and DynamicObstacles-5x5, different concepts have slightly varying concept errors, indicating visual tasks have different difficulties for GPT-4o.
-Agent direction has as high as around 0.3 prediction error for both DoorKey-7x7 and DynamicObstacles-5x5.
-Direction movable is also hard for DoorKey-7x7, with a high concept error even if it is a binary concept.
+Different tasks and concepts have various difficulties for GPT-4o.
+Table 4 list detailed concept errors for concepts in all environments.
+In PixelCartPole, cart position and pole angle have smaller errors, while velocities require understanding multiple frames and thus are harder to predict accurately.
+For DoorKey and DynamicObstacles, different concepts have slightly varying concept errors, indicating visual tasks have different difficulties for GPT-4o.
+Agent direction has as high as around 0.3 prediction error for both DoorKey and DynamicObstacles.
+Direction movable is also hard for DoorKey, with a high concept error even if it is a binary concept.
 We posit it requires the correct understanding of more than one particular object to ensure correctness.
-The concept accuracies in DoorKey-7x7 are generally higher than DynamicObstacles-5x5, suggesting GPT-4o more struggles with a larger grid.
-Example Behavior of Agent with GPT-4o-labeled concepts In Figure 4, we show an example of a GPT-4o-trained RL agent on DynamicObstacles-5x5, in which the agent appears to wait until it is safe to move towards the goal: the green square.
+The concept accuracies in DoorKey are generally higher than DynamicObstacles, suggesting GPT-4o more struggles with a larger grid.
+A GPT-4o-trained RL agent recovers from a mistake.
+In Figure 8, we show an example of a GPT-4o-trained RL agent on DynamicObstacles, in which the agent appears to wait until it is safe to move towards the goal: the green square.
 The image sequence shows the agent (red triangle) starting from its initial position and moving to the right.
@@
 Algorithm $R$uparrow $c$ Error $\downarrow$
-PixelCartPole-v1 LICORICE-IT 0.533 896±0.257 169plus-or-minus0.5338960.257169\$0.533\,896$\pm$\$0.257\,169\$0.533 896 ± 0.257 169 0.079 622±0.026 734plus-or-minus0.0796220.026734\$0.079\,622$\pm$\$0.026\,734\$0.079 622 ± 0.026 734
+PixelCartPole LICORICE-IT 0.533 896±0.257 169plus-or-minus0.5338960.257169\$0.533\,896$\pm$\$0.257\,169\$0.533 896 ± 0.257 169 0.079 622±0.026 734plus-or-minus0.0796220.026734\$0.079\,622$\pm$\$0.026\,734\$0.079 622 ± 0.026 734
 LICORICE-DE 0.9669±0.051 17plus-or-minus0.96690.05117\$0.9669$\pm$\$0.051\,17\$0.9669 ± 0.051 17 0.031 106±0.010 96plus-or-minus0.0311060.01096\$0.031\,106$\pm$\$0.010\,96\$0.031 106 ± 0.010 96
@@
 LICORICE $1.0$\pm$0.0$$ 0.024 797±0.004 166plus-or-minus0.0247970.004166\$0.024\,797$\pm$\$0.004\,166\$0.024 797 ± 0.004 166
-DoorKey-7x7 LICORICE-IT 0.992 412±0.007 62plus-or-minus0.9924120.00762\$0.992\,412$\pm$\$0.007\,62\$0.992 412 ± 0.007 62 0.089 535±0.015 02plus-or-minus0.0895350.01502\$0.089\,535$\pm$\$0.015\,02\$0.089 535 ± 0.015 02
+DoorKey LICORICE-IT 0.992 412±0.007 62plus-or-minus0.9924120.00762\$0.992\,412$\pm$\$0.007\,62\$0.992 412 ± 0.007 62 0.089 535±0.015 02plus-or-minus0.0895350.01502\$0.089\,535$\pm$\$0.015\,02\$0.089 535 ± 0.015 02
 LICORICE-DE 0.783 113±0.079 936plus-or-minus0.7831130.079936\$0.783\,113$\pm$\$0.079\,936\$0.783 113 ± 0.079 936 0.200 022±0.051 719plus-or-minus0.2000220.051719\$0.200\,022$\pm$\$0.051\,719\$0.200 022 ± 0.051 719
@@
 LICORICE 0.990 297±0.007 106plus-or-minus0.9902970.007106\$0.990\,297$\pm$\$0.007\,106\$0.990 297 ± 0.007 106 0.047 504±0.011 097plus-or-minus0.0475040.011097\$0.047\,504$\pm$\$0.011\,097\$0.047 504 ± 0.011 097
-DynamicObstacles-5x5 LICORICE-IT 0.996 889±0.002 139plus-or-minus0.9968890.002139\$0.996\,889$\pm$\$0.002\,139\$0.996 889 ± 0.002 139 0.001 202±0.000 982plus-or-minus0.0012020.000982\$0.001\,202$\pm$\$0.000\,982\$0.001 202 ± 0.000 982
+DynamicObstacles LICORICE-IT 0.996 889±0.002 139plus-or-minus0.9968890.002139\$0.996\,889$\pm$\$0.002\,139\$0.996 889 ± 0.002 139 0.001 202±0.000 982plus-or-minus0.0012020.000982\$0.001\,202$\pm$\$0.000\,982\$0.001 202 ± 0.000 982
 LICORICE-DE 0.980 537±0.008 092plus-or-minus0.9805370.008092\$0.980\,537$\pm$\$0.008\,092\$0.980 537 ± 0.008 092 0.000 114±0.000 18plus-or-minus0.0001140.00018\$0.000\,114$\pm$\$0.000\,18\$0.000 114 ± 0.000 18
@@
 LICORICE 0.990 607±0.003 563plus-or-minus0.9906070.003563\$0.990\,607$\pm$\$0.003\,563\$0.990 607 ± 0.003 563 0.000 193±0.000 432plus-or-minus0.0001930.000432\$0.000\,193$\pm$\$0.000\,432\$0.000 193 ± 0.000 432
-Table 9: Ablation study results for LICORICE in all environments, where we delete every single component to compare with LICORICE. This shows a more complete version of the results in Table 4.
-Figure 5: Learning curves for all ablations. Shaded region shows 95% CI, calculated using 1000 bootstrap samples.
+Boxing LICORICE-IT 0.978 152±0.036 465plus-or-minus0.9781520.036465\$0.978\,152$\pm$\$0.036\,465\$0.978 152 ± 0.036 465 0.135 606±0.069 144plus-or-minus0.1356060.069144\$0.135\,606$\pm$\$0.069\,144\$0.135 606 ± 0.069 144
+LICORICE-DE 1.001 046±0.042 38plus-or-minus1.0010460.04238\$1.001\,046$\pm$\$0.042\,38\$1.001 046 ± 0.042 38 0.044 79±0.023 005plus-or-minus0.044790.023005\$0.044\,79$\pm$\$0.023\,005\$0.044 79 ± 0.023 005
+LICORICE-AC 0.994 306±0.009 63plus-or-minus0.9943060.00963\$0.994\,306$\pm$\$0.009\,63\$0.994 306 ± 0.009 63 0.087 981±0.063 943plus-or-minus0.0879810.063943\$0.087\,981$\pm$\$0.063\,943\$0.087 981 ± 0.063 943
+LICORICE 1.004 997±0.017 713plus-or-minus1.0049970.017713\$1.004\,997$\pm$\$0.017\,713\$1.004 997 ± 0.017 713 0.049±0.010 967plus-or-minus0.0490.010967\$0.049$\pm$\$0.010\,967\$0.049 ± 0.010 967
+Table 9: Ablation study results for LICORICE in the first four environments, where we study the effect of each component to compare with LICORICE.
+This shows a more complete version of the results in Table 3.
+Figure 9: Learning curves for ablations in the first three environments. Shaded region shows 95% CI, calculated using 1000 bootstrap samples.
 In Table 9, we present an extension of the results in Table 3, including standard deviation.
 Figure 9 shows all learning curves for ablations in all environments.
-In PixelCartPole-v1, we clearly see the benefit of iterative strategies on reward: LICORICE, LICORICE-DE and LICORICE-AC consistently increase in reward and converge at high levels, but LICORICE-IT converges at less than $50\%$ of the optimal reward.
+In PixelCartPole, we clearly see the benefit of iterative strategies on reward: LICORICE, LICORICE-DE and LICORICE-AC consistently increase in reward and converge at high levels, but LICORICE-IT converges at less than $50\%$ of the optimal reward.
 We also see that LICORICE-IT also achieves higher concept error, indicating that it struggles to learn the larger distribution of concepts induced by a non-optimal policy.
-In DoorKey-7x7, all algorithms steadily increase in reward.
+In DoorKey, all algorithms steadily increase in reward.
 However, we can see a dip at around $10^{6}$ environment steps where we begin the second iteration for LICORICE, LICORICE-AC, and LICORICE-DE.
 Although LICORICE-IT achieves similar reward to LICORICE, LICORICE achieves lower concept error, which is beneficial for the goal of interpretability.
-Finally, in DynamicObstacles-5x5, LICORICE-AC lags behind the most in terms of both reward and concept error.
+Finally, in DynamicObstacles, LICORICE-AC lags behind the most in terms of both reward and concept error.
 This result indicates that the active learning component is most important for this environment.
-Figure 6: Performance of all ablations at the final training iteration, with black bars representing the 95% confidence interval, calculated using 1000 bootstrap samples.
-We now inspect the performance at the last training iteration in Figure 6.
-Overall, we find that LICORICE performs better than or equal to the ablations on all environments.
-It enjoys the best reward performance on PixelCartPole-v1, while performing similarly to LICORICE-IT on DoorKey-7x7 and to LICORICE-IT and LICORICE-DE on DynamicObstacles-5x5.
-It exhibits the lowest concept error on PixelCartPole-v1 and DoorKey-7x7, and achieves similarly low error to LICORICE-DE and LICORICE-IT on DynamicObstacles-5x5.
-In PixelCartPole-v1, LICORICE achieves the highest reward, indicating the benefits of multiple iterations, active learning, and decorrelation, which are absent in LICORICE-IT, LICORICE-AC, and LICORICE-DE, respectively.
-LICORICE also enjoys the lowest concept error, showcasing its ability to learn and apply concepts accurately.
-LICORICE-AC and LICORICE-DE perform well in minimizing concept error, while LICORICE-IT shows a relatively higher error rate, underscoring the importance of multiple iterations in some environments.
+Impact on data diversity.
+We hypothesize that data decorrelation enables better performance by increasing the diversity of the training data.
+As a result, we investigate the role of data decorrelation on state diversity by calculating the average distance to the $k$ nearest samples in the concept label space for all labeled samples.
+The intuition is that a larger average distance to nearest neighbors indicates samples are more spread out in the space, suggesting higher diversity in the training data.
+Let $D=\{(s_{1},c_{1}),\dots,(s_{b},c_{b})\}$ be the labeled dataset and $\text{Neighbor}(c_{i},k)$ be the index set of $k$ nearest neighbors with $k=10$ for all environments, then we calculate diversity as:
+$d=\frac{1}{bk}\sum_{i=1}^{b}\sum_{j\in\text{Neighbor}(c_{i},k)}||c_{i}-c_{j}||.$
+We present $d$ along with the reward in Table 10 for LICORICE and LICORICE-DE.
+As hypothesized, decorrelation increases diversity in 3 of the 4 environments.
+LICORICE achieves higher reward than LICORICE-DE.
+In DynamicObstacles, the state diversity and reward is similar for both LICORICE and LICORICE-DE.
+We believe that, in this environment, decorrelation is less critical because neighboring states are inherently more dissimilar due to the dynamic nature of the obstacles.
+LICORICE LICORICE-DE
+Environment $d$ $R$ $d$ $R$
+PixelCartPole 0.12 $\pm$ 0.00 $1.0$\pm$0.0$$ 0.10 $\pm$ 0.01 0.9669±0.051 17plus-or-minus0.96690.05117\$0.9669$\pm$\$0.051\,17\$0.9669 ± 0.051 17
+DoorKey 1.12 $\pm$ 0.18 0.990 297±0.007 106plus-or-minus0.9902970.007106\$0.990\,297$\pm$\$0.007\,106\$0.990 297 ± 0.007 106 0.25 $\pm$ 0.00 0.783 113±0.079 936plus-or-minus0.7831130.079936\$0.783\,113$\pm$\$0.079\,936\$0.783 113 ± 0.079 936
+DynamicObstacles 0.93 $\pm$ 0.04 0.990 607±0.003 563plus-or-minus0.9906070.003563\$0.990\,607$\pm$\$0.003\,563\$0.990 607 ± 0.003 563 0.99 $\pm$ 0.09 0.980 537±0.008 092plus-or-minus0.9805370.008092\$0.980\,537$\pm$\$0.008\,092\$0.980 537 ± 0.008 092
+Boxing 7.20 $\pm$ 0.20 0.968 879±0.036 656plus-or-minus0.9688790.036656\$0.968\,879$\pm$\$0.036\,656\$0.968 879 ± 0.036 656 6.53 $\pm$ 0.08 0.970 273±0.009 912plus-or-minus0.9702730.009912\$0.970\,273$\pm$\$0.009\,912\$0.970 273 ± 0.009 912
+Table 10: Training data diversity ($d$) and reward ($R$) with and without data decorrelation.
+Higher $d$ is better (more diverse).
+For each environment, the highest $d$ is bolded.
+When the standard deviations overlap, we bold both.
+B.5 Hyperparameter Sensitivity
+Figure 10: Hyperparameter sensitivity results for DynamicObstacles.
+Bars represent standard deviation.
+Here, bars are grouped by the same $p$ values (acceptance probability).
+Bars of the same color share the same $\tau$ value (budget multiplier), and bars with the same pattern share the same value for $N$ (committee size).
+Generally, $p=0.1$ tend to perform better overall with greater stability.
+Having a larger committee ($N=5$ vs.
+$N=3$) also tends to help with performance.
+We now seek to understand how sensitive LICORICE is to the choice of hyperparameters.
+Figure 10 shows the results of this experiment for DynamicObstacles.
+Interestingly, we find that the main difference between the settings is the variation.
+Overall, LICORICE is relatively robust to the choice of hyperparameter settings in this environment, with the final average reward ranging between 96% and 99.5% of the optimal.
+The concept error for all settings was $00.00\pm 0$ 0.00, so we do not plot it here.
+B.6 Choice of Acquisition Strategy
+Budget 100 200 300
+R c Error R c Error R c Error
+Disagreement 0.96 0.05 0.97 0.01 1.00 0.00
+Entropy 0.95 0.07 0.98 0.02 0.99 0.00
+Table 11: Comparison of disagreement- and entropy-based active learning strategies for DynamicObstacles under different labeling budgets. There is essentially no difference between the two strategies.
+We now investigate the impact of the choice of acquisition strategy on both reward and concept accuracy.
+We implemented an entropy-based disagreement measurement for the committee and studying its effects on a MiniGrid environment.
+Specifically, we implement the following vote entropy acquisition function: $U(s_{i})=-\frac{1}{K}\sum_{k=1}^{K}\sum_{j=1}^{P_{k}}\frac{V_{ijk}}{N}\log\frac{V_{ijk}}{N}$, where $s_{i}$ is the input state, $K$ is the number of concepts we learn, $P_{k}$ is the number of classes for the $k$-th concept, $V_{ijk}$ is the number of votes for the $j$-th class of the $k$-th concept for state $s_{i}$, and $N$ is the number of committee members.
+Table 11 shows these results on DynamicObstacles.
+There is essentially no difference between the two strategies in terms of both reward and concept error.
 Appendix C Additional Discussion
@@
 This limitation can impact the overall performance of concept-based models, especially in domains where continuous data is prevalent.
+Furthermore, although VLMs can be successfully used for automatic labeling of some concepts, there are still hallucination issues (Achiam et al., 2023) and other failure cases, such as providing inaccurate counts.
+We believe that future work that seeks to improve general VLM capabilities and mitigate hallucinations would also help overcome this limitation.
 Addressing this issue could involve developing specialized techniques or using existing tools and libraries to better complement VLM capabilities.
@@
 Our current method employs an disagreement-based acquisition function to select the most informative data points for labeling.
-While this approach is effective, there is potential for exploring more sophisticated active learning strategies, such as incorporating advanced exploration-exploitation trade-offs or leveraging recent advancements in active learning algorithms [28].
+While this approach is effective, there is potential for exploring more sophisticated active learning strategies, such as incorporating advanced exploration-exploitation trade-offs or leveraging recent advancements in active learning algorithms (Tharwat & Schenck, 2023).
+Furthermore, scaling challenges may arise from environmental complexity, such as when only a subset of given concepts are relevant or when learning certain concepts is prerequisite to gathering training data for others.
+Future research could explore using attention mechanisms (Vaswani et al., 2017) or sparse coding (Olshausen & Field, 1996) to identify relevant concepts.
+Another exciting direction is the work in human-AI complementarity and learning-to-defer algorithms (Mozannar et al., 2023; Verma et al., 2023) to train an additional classifier for deferring labeling to a person when the chance of an error is high.
+In our work, we assume a known set of concepts for the environment; future work could investigate the use of human-VLM teams to determine a reasonable concept set for the environment.
 Finally, designing a concept-based representation for RL remains an open challenge.
 Our work provides a few illustrative examples, but the exact design of these representations can significantly impact performance, often for reasons that are not entirely clear — especially when using VLMs as annotators.
-Prior work [7] proposed some desiderata for concepts in RL, but future work could refine these principles, especially in the face of VLM annotators.
+Prior work (Das et al., 2023) proposed some desiderata for concepts in RL, but future work could refine these principles, especially in the face of VLM annotators.
 Future work could also include systematically investigating the factors that influence the effectiveness of different concept-based representations in RL.
@@
 C.2 Broader Impacts
-Interpretability in RL
+Interpretability in RL.
 Incorporating concept learning with RL presents both positive and negative societal impacts.
 On the positive side, promoting interpretability and transparency in decision-making fosters trust and accountability.
-However, in cases where it yields incorrect results, stakeholders might be misled into trusting flawed decisions due to the perceived transparency of the model [12].
+However, in cases where it yields incorrect results, stakeholders might be misled into trusting flawed decisions due to the perceived transparency of the model (Kaur et al., 2020).
 Unintended misuse could also occur if stakeholders lack the technical expertise to accurately interpret the models, leading to erroneous conclusions and potentially harmful outcomes.
 To mitigate these risks, an avenue for future work is developing clear guidelines for interpreting these models and tools to scaffold non-experts’ understanding of the model outputs.
-Using VLMs for Concept Labeling
+Using VLMs for concept labeling.
 On one hand, VLMs have the potential to significantly improve the efficiency and scalability of labeling processes, which can accelerate advancements in various fields.